### v1

# Non-commutative rings with infinitely many maximal subrings

Abstract.
We study rings with infinitely (only finitely) many maximal subrings. We prove that if $M$ is a maximal left/right ideal of a ring $T$ which is not an ideal of $T$, and $R$ is the idealizer of $M$, then $T$ has at least $|R/M|+1$ maximal left/right ideals which are not an ideal of $T$; in particular $T$ has at least $|R/M|+1$ distinct maximal subrings. Moreover, if $T$ is a $K$-algebra over an infinite field $K$, then either $T$ has infinitely many maximal subrings or $T$ is a quasi duo ring with certain algebraic properties similar to commutative rings. We prove that for a simple ring $R$, the ring $R\times R$ has only finitely many maximal subrings if and only if $R$ is finite. Also we study rings which are integral over their centers and have only finitely many maximal subrings. We prove that if $T$ is integral over its center and $T$ has more than $2^{\aleph_{0}}$ maximal (left/right) ideals, then $T$ has infinitely many maximal subrings. In particular, we see that if a $J$-semisimple ring $T$ is integral over its center and has only finitely many maximal subrings, then $T$ embeds in $S\times\prod_{i\in I}E_{i}$, where each $E_{i}$ is an absolutely algebraic field and $S$ is a finite semisimple ring. We see that if $T$ is a left Noetherian algebraic $K$-algebra over an infinite field $K$ and $T$ has only finitely many maximal subrings, then $T$ is a countable left Artinian ring which is integral over $\mathbb{Z}_{p}$, where $p=Char(K)$. We exactly determine when $T=\prod_{i\in I}\mathbb{M}_{n_{i}}(E_{i})$, where each $E_{i}$ is a field and $n_{i}\in\mathbb{N}$, has only finitely many maximal subrings. We see that if $R$ is an infinite Artinian ring, then $\mathbb{M}_{n}(R)$, $n>1$, and $R\times R$ have infinitely many maximal subrings.
Key words and phrases:
maximal subring, non-commutative ring, idealizer, left maximal ideal, integrality, direct product
2010 Mathematics Subject Classification:
16U80;16U60;16U70;16N99;16P20;16P40
Department of Mathematics, Faculty of Mathematical Sciences and Computer,
Shahid Chamran University of Ahvaz, Ahvaz-Iran
a-azarang@scu.ac.ir
ORCID ID: orcid.org/0000-0001-9598-2411
1. Introduction
Let $T$ be an associative ring with identity $1\neq 0$ and $RgMax(T)$ be the set of all maximal subrings (which contain $1_{T}$) of $T$, then we study when $RgMax(T)$ is finite/infinite. In [2], the author characterized fields and affine rings which have only finitely many maximal subrings; and in [10], the authors studied commutative rings with infinitely/only finitely many maximal subrings. We present some facts from them (which are needed in this paper) as follows:
  1. (1)
    A field $E$ has only finitely many maximal subrings if and only if $E$ has a subfield $F$ without maximal subrings with $[E:F]$ is finite; if and only if every descending chain $\cdots\subset R_{2}\subset R_{1}\subset R_{0}=E$, where each $R_{i}$ is a maximal subring of $R_{i-1}$, $i\geq 1$, is finite; moreover $F$ is unique, all such chains have a same length, say $m$, and $R_{m}=F$, see [2, Theorems 3.2 and 3.6].
  2. (2)
    Let $R=F[\alpha_{1},\ldots,\alpha_{n}]$ be an affine integral domain over a field $F$. Then $R$ has only finitely many maximal subrings if and only if $F$ has only finitely many maximal subrings and each $\alpha_{i}$ is algebraic over $F$, see [2, Theorem 4.1].
  3. (3)
    For each ring $R$, either $R$ is a Hilbert ring or $RgMax(R)$ is infinite, see [10, Corollary 1.9].
  4. (4)
    If $R$ is an Artinian ring which is either uncountable or of zero characteristic and or is not integral over its prime subring, then $RgMax(R)$ is infinite, see [10, Theorem2.5].
  5. (5)
    If $R$ is a Noetherian ring with $|R|>2^{\aleph_{0}}$, then $|RgMax(R)|\geq 2^{\aleph_{0}}$, see [10, Theorem 2.7].
  6. (6)
    If $\{R_{i}\ |\ i\in I\}$ is a family of rings and $R=\prod_{i\in I}R_{i}$, then in [10, Theorem 3.9], the authors exactly determined when $R$ has only finitely many maximal subrings. Moreover, if $I$ is infinite then $|RgMax(R)|\geq 2^{|I|}$, see [9, Remark 3.18].
  7. (7)
    If $K$ is a field, then $K\times K$ has finitely many maximal subrings if and only if $K$ is a finite field, see [10, Corollary 3.5].
  8. (8)
    If $R$ is a semilocal ring, then either $R$ has infinitely many maximal subrings or $R$ is integral over $\mathbb{Z}_{n}$, for some $n>1$, see [10, Corollary 2.2].
  9. (9)
    If $R$ is a ring with $|Max(R)|>2^{\aleph_{0}}$ (resp. $|R/J(R)|>2^{2^{\aleph_{0}}}$), then $|RgMax(R)|\geq|Max(R)|$ (resp. $|RgMax(R)|\geq 2^{\aleph_{0}}$), see [10, Proposition2.8].
  10. (10)
    If $R$ is a ring, then $R\times R$ has only finitely many maximal subrings if and only if $R$ is a semilocal ring with only finitely many maximal subrings and $R/M$ is finite for each $M\in Max(R)$, see [10, Corollary 3.10].
Finally in [10], the authors generalized $(1)$ for a finite direct products of fields too. For non-commutative rings, the existence of maximal subring first studied in [1], we will mentioned some of the result of [1] in the next section. It is interesting to know that each ring $R$ can be considered as a maximal subring of a larger ring $T$, see [1, Theorem 3.7]. Recently, in [5] and [6] the existence and the structure of maximal subrings in division rings and the existence of maximal subrings in certain non-commutative rings investigated, respectively. In [20] and [7], the authors systematically studied maximal commutative subrings for general noncommutative rings and commutative maximal subrings, respectively. We refer the interested reader to [24], for classification of maximal subrings of finite semisimple rings and their applications to the calculation of the covering number of a finite semisimple ring; and also to [3, 4] for more results about the conductor ideals and certain ideals of the extension $R\subseteq T$, where $R$ is a maximal subring of $T$.
A brief outline of this paper is as follows. In Section $2$, we give some preliminaries results that needed in this paper, specially about the number of maximal left/right ideals of a ring $T$ which are not ideals of $T$. In fact, we prove that if $T$ is a ring and $M$ is a maximal left/right ideal of $T$ which is not an ideal of $T$ and $R$ is the idealizer of $M$ in $T$, then $T$ has at least $|R/M|+1$ maximal left/right ideals $N$ such that $N$ is similar to $M$ (i.e., $T/N\cong T/M$ as left/right $T$-module) and therefore $T$ has at least $|R/M|+1$ maximal left/right ideals which are not ideals of $T$. Moreover, if $T$ is a $K$-algebra over an infinite field $K$, then either $T$ is a quasi duo ring or $T$ has infinitely many maximal left/right ideals which are not ideals of $T$ (in fact, either $|Max_{l}(T)\setminus Max(T)|\geq|K|$ or $|Max_{r}(T)\setminus Max(T)|\geq|K|$). Finally we present some needed results from [27] about the (number of) maximal left/right ideals of $\mathbb{M}_{n}(R)$, where $R$ is a ring and $n\in\mathbb{N}$, and also from [1] about maximal subrings of non-commutative rings.
In Section $3$, we investigate on the number of maximal subrings in general non-commutative rings. First we see that if $D$ is a division ring with finitely many maximal subrings, then the center of $D$ is algebraically closed in $D$. For a simple ring $S$, we exactly determine all maximal subrings of $S\times S$. We see that if $S$ is a simple ring which is not a division ring, then $S$ has only finitely many maximal subrings if and only if $S$ is a finite ring. Consequently, if $S$ is a simple then $S\times S$ has only finitely many maximal subrings if and only if $S$ is a finite ring. Next, we see that if a ring $T$ has only finitely many maximal subrings, then $Max_{l}(T)\setminus Max(T)$, $Max_{r}(T)\setminus Max(T)$, $Max(T)\setminus Max_{l}(T)$, $Prm_{l}(T)\setminus Max(T)$ and $Prm_{l}(T)\setminus Max_{l}(T)$ (and analogously for right sets) are all finite, moreover if $M$ is a maximal left/right ideal of $T$ which is not an ideal of $T$, then $\mathbb{I}(M)/M$ is a finite field and if $N$ is a maximal ideal of $T$ which is not a maximal left ideal of $T$, then $T/N\cong\mathbb{M}_{n}(F)$, for some finite field $F$ and $n>1$. Also we see that if $T$ is a ring with only finitely many maximal subrings, then $J(T)=\bigcap_{M\in Max(T)}M$. In particular, if we assume that $T$ is a $K$-algebra over an infinite field $K$, then either $T$ has infinitely many maximal subrings or $T$ is a quasi duo ring. Finally in Section $3$, we prove that if $T$ is an infinite ring with nonzero characteristic, then either $T$ has infinitely many maximal subrings or $\{M\in Max(T)\ |\ |T/M|<\aleph_{0}\}$ is a finite set.
In Section $4$, we studied rings which are integral over their centers and have only finitely many maximal subrings. In particular, algebraic $K$-algebra over an infinite field $K$. We prove that if $T$ is a ring which is integral over its center and $T$ has more than $2^{\aleph_{0}}$ maximal (left/right) ideals, then $T$ has infinitely many maximal subrings. We observe that, if $T$ is an algebraic $K$-algebra over an infinite field $K$, then either $T$ has infinitely many maximal subring or $T/J(T)$ is a commutative von Neumann regular ring which is integral over $\mathbb{Z}_{p}$, where $p=Char(K)$. We see that if $T$ is integral over its center, then either $T$ has infinitely many maximal subrings or $Prm_{l}(T)=Max(T)=Prm_{r}(T)$, and for each maximal ideal $M$ of $T$, either $T/M$ is an absolutely algebraic field (with only finitely many maximal subrings) or $T/M$ is a finite simple ring; in particular, if in addition $J(T)=0$, then $T$ embeds in $S\times\prod_{i\in I}E_{i}$, where $S$ is a finite semisimple ring and each $E_{i}$ is an absolutely algebraic field.
Finally in Section $5$, we exactly determine when $T=\prod_{i\in I}\mathbb{M}_{n_{i}}(E_{i})$, where each $E_{i}$ is a field and $n_{i}\in\mathbb{N}$, has only finitely many maximal subrings. We see that if $R$ is an infinite Artinian ring, then $\mathbb{M}_{n}(R)$, $n>1$, and $R\times R$ have infinitely many maximal subrings.
All rings in this paper are unital with $1\neq 0$. All subrings, modules and homomorphisms are also unital. If $R\subsetneq T$ is a ring extension and there exists no other subring between $R$ and $T$, then $R$ is called a maximal subring of $T$, or the extension $R\subseteq T$ is called a minimal ring extension, see [14, 15, 16, 18, 25]. If $T$ is a ring, $I$ is an ideal of $T$ and $M$ is a left (resp. right) $T$-module, then $Min_{T}(I)$, $Max_{r}(T)$, $Max_{l}(T)$, $Max(T)$ and $l.ann_{T}(M)$ (resp. $r.ann_{T}(M)$), denote the set of all minimal prime ideals of $I$ in $T$, the set of all maximal right ideals of $T$, the set of all maximal left ideals of $T$, the set of all maximal ideals of $T$ and the left annihilator of $M$ in $T$ (resp. the right annihilator of $M$ in $T$), respectively. We use $Min(T)$ for $Min_{T}(0)$. The set of all left (resp. right) primitive ideals of $T$ is denoted by $Prm_{l}(T)$ (resp. $Prm_{r}(T)$). The characteristic of a ring $T$ is denoted by $Char(T)$. If $T$ is a ring, then $dim(T)$ denotes the classical Krull dimension of $T$ (i.e., the supremum of the lengths of all chains of prime ideals of $T$). Also we use $Aut(T)$ for the set of all ring automorphism of a ring $T$. $v.dim_{K}(T)$ denotes the vector dimension of a $K$-algebra $T$ over a field $K$. $J(T)$ denotes the Jacobson radical of a ring $T$. For a ring $T$, $Nil_{*}(T)$, $Nil^{*}(T)$, $N(T)$ and $U(T)$ are the lower nilradical, the upper nilradical, the set of all nilpotent elements and the units group of $T$, respectively. If $T$ is a ring, the opposite ring of $T$ is denoted by $T^{op}$. A prime ideal $P$ of a ring $T$ is called strongly (resp. completely) prime if $T/P$ has no nonzero nil ideal (resp. $T/P$ is a domain). If $T$ is a ring, then $\mathbb{M}_{n}(T)$ denote the ring of all $n\times n$ square matrices over $T$. If $M$ is a left (resp. right) module over a ring $T$, then the ring of all $T$-module homomorphisms of $M$ is denoted by $End({}_{T}M)$ (resp. $End(M_{T})$). If $A$ is a right ideal of a ring $T$, then $\mathbb{I}(A)$ denotes the idealizer of $A$ in $T$, that is the largest subring of $T$ which $A$ is a two-sided ideal in it, i.e., $\mathbb{I}(A)=\{t\in T\ |\ tA\subseteq A\}$, the similar notation is used when $A$ is a left ideal of $T$. If $A$ is a right ideal of a ring $T$, then $\mathbb{E}_{T}(A):=\mathbb{I}(A)$ is called the eigenring of $A$. If $X$ is a subset of a ring $T$, then $C_{T}(X)$ is the centralizer of $X$ in $T$, in particular, $C(T)=C_{T}(T)$ is the center of $T$. For a ring $T$, the subring $\{n.1_{T}\ |\ n\in\mathbb{Z}\}$ is called the prime subring of $T$. It is clear that $T$ and $T^{op}$ have a same prime subring (moreover, $C(T)=C(T^{op})$) and therefore have a same characteristic. A ring $T$ is called left (resp. right) quasi duo if each maximal left (resp. right) ideal of $T$ is a two-sided ideal of $T$, see [22]. $T$ is called quasi duo, if $T$ is left and right quasi duo ring. A ring $T$ is called left (resp. right) duo if each left (resp. right) ideal of $T$ is two-sided. Similarly duo rings are defined. If $R$ is a subring of a ring $T$ and $t\in T$, then we say that $t$ is left (resp. right) integral over $R$, if $t$ is a root of a left (resp. right) monic polynomial of degree $n\geq 1$ over $R$. $R$ is called left (resp. right) integrally closed in $T$, if every left (resp. right) integral element of $T$ over $R$, belongs to $R$. $R$ is called integrally closed in $T$, whenever $R$ is left and right integrally closed in $T$. If $T$ is a ring and $R$ is a subring of the center of $T$, then $T$ is called an $R$-algebra. It is clear that, if $T$ is an $R$-algebra, then the definitions of right and left integral elements and also left or right integrally closed for the extension $R\subseteq T$ are coincided. If $F\subseteq E$ is a field extension, then $tr.deg(E/F)$ denotes the transcendence degree of $E$ over $F$. For other notations and definitions we refer the reader to [19, 21, 26].
2. Preliminaries
Let $T$ be a ring and $I$ and $J$ be left ideals of $T$. Then $I$ and $J$ are called similar left ideals if $T/I\cong T/J$ as left $T$-module. If $a\in T$ and define $(I:a):=\{x\in T\ |\ xa\in I\}$, then it is not hard to see that $I$ and $J$ are similar if and only if there exists $c\in T$ such that $I+Tc=T$ and $J=(I:c)$, see [13, Proposition 1.3.6]. Since for a left ideal $I$ of a ring $T$, we have the ring isomorphism $\mathbb{I}(I)/I\cong(End({}_{T}(T/I)))^{op}$, see [23, P. 14, 1.11], we immediately conclude that if $I$ and $J$ are similar left ideals of a ring $T$, then $\mathbb{I}(I)/I\cong\mathbb{I}(J)/J$ as rings, i.e., up to ring isomorphism, the ring $\mathbb{I}(I)/I$ is unique for similar left ideals; in particular $|\mathbb{I}(I)/I|=|\mathbb{I}(J)/J|$. In other words the cardinality of the ring $\mathbb{I}(I)/I$ is invariant up to similarity. Moreover, if $M$ is a maximal left ideal of a ring $T$, then we have the following facts too:
  1. (1)
    a maximal left ideal $N$ of $T$ is similar to $M$ if and only if $N=(M:c)$ for some $c\in T\setminus M$.
  2. (2)
    if $N$ is a (maximal) left ideal of $T$ which is similar to $M$, then $N$ is an ideal of $T$ if and only if $N=M$ if and only if $M$ is an ideal of $T$ .
To see the above facts, note that $(1)$ is an immediate consequences of [13, Proposition 1.3.6]. For $(2)$, let $M$ and $N$ be similar maximal left ideals of $T$ and assume that $N$ is an ideal of $T$. Since $T/N\cong T/M$ as left $T$-module we immediately conclude that $N=l.ann_{T}(T/N)=l.ann_{T}(T/M)\subseteq M$ and therefore $M=N$ (note that $N$ is a maximal left ideal of $T$, for $T/N\cong T/M$ as left $T$-module). Hence $(2)$ holds. Consequently, if $M$ is a maximal left ideal of $T$ which is not an ideal of $T$, and $[M]$ be the set of all (maximal) left ideals of $T$ which are similar to $M$, we deduce that $M]\subseteq Max_{l}(T)\setminus Max(T)$, by $(2)$, i.e., each (maximal) left ideals of $T$ which is similar to $M$ is not an ideal of $T$. Clearly, the relation similarity is an equivalence relation on the set of all left ideals of $T$ and in particular on $Max_{l}(T)$. By $(1)$ and $(2)$, it is obvious that if $M$ is a maximal left ideal of $T$, then $|[M]|=1$ (i.e., $M]=\{M\}$) if and only if $M$ is an ideal of $T$ (thus $M\in Max(T)\cap Max_{r}(T)$).
It is clear that whenever $I$ is a left ideal of a ring $T$ and $c\in T$, then $(I:c)=T$ if and only if $c\in I$. Moreover, $I\subseteq(I:c)$ if and only if $c\in\mathbb{I}(I)$. In particular, if $M$ is a maximal left ideal of a ring $T$, from these facts we immediately conclude that $(M:c)=M$ if and only if $c\in\mathbb{I}(M)\setminus M$. Hence $M]=\{M\}\cup\{(M:c)\ |\ c\in T\setminus\mathbb{I}(M)\}$. In the following main result we find a lower bound for $|[M]|$. Before present it, let us remind that from the previous observation for each $N\in[M]$, we have $|\mathbb{I}(M)/M|=|\mathbb{I}(N)/N|$. Now the following is in order.
Theorem 2.1.
Let $T$ be a ring and $M\in Max_{l}(T)$ is not an ideal of $T$ (i.e., $M$ is not a maximal ideal of $T$). Then $|[M]|\geq|\mathbb{I}(M)/M|+1$. In particular, $|Max_{l}(T)\setminus Max(T)|\geq|\mathbb{I}(M)/M|+1$.
Proof.
Let $R=\mathbb{I}(M)$ and $D=R/M$. Since $M$ is not an ideal of $T$ we conclude that $R$ is a proper subring of $T$. Take a fixed element $x\in T\setminus R$. We claim that the function $c+M\longmapsto(M:x+c)$ is a one-one function from $D$ to $M]\setminus\{M\}$. Since $x\in T\setminus R$ and $c\in R$ we deduce that $x+c\in T\setminus R$ and therefore $(M:x+c)\in[M]\setminus\{M\}$, by the previous observation. Hence it remain to show that the function is one-one. Hence assume that $a,b\in R$ and $(M:x+a)=(M:x+b)$, we must show that $a+M=b+M$, i.e., $a-b\in M$. Let $t\in(M:x+a)=(M:x+b)$, then we conclude that $tx+ta,tx+tb\in M$. Therefore $t(a-b)\in M$, i.e., $(M:x+a)\subseteq(M:a-b)$. Now if $a-b\notin M$, then we conclude that $a-b\in R\setminus M$ and therefore (M:ab)=M(M:a-b)=M. Hence $(M:x+a)\subseteq M$ and therefore by maximality of $(M:x+a)$ we obtain that $(M:x+a)=M$ which is absurd, for $x+a\in T\setminus R$. Thus $a-b\in M$ and therefore we are done. ∎
As an immediate consequence of the previous theorem is that if $T$ is a ring and $Max_{l}(T)\setminus Max(T)\neq\emptyset$, then $|Max_{l}(T)\setminus Max(T)|\geq 3$, i.e., if a ring $T$ has a maximal left ideal which is not an ideal of $T$, then $T$ has at least three maximal left ideals which are not ideals of $T$.
Corollary 2.2.
Let $T$ be a $K$-algebra over a field $K$. Then either $T$ is a left quasi duo ring or $T$ has at least $|K|+1$ maximal left ideals which are not ideals of $T$ (i.e., $|Max_{l}(T)\setminus Max(T)|\geq|K|+1$).
Proof.
Assume that $T$ is not a left quasi duo ring and $M$ be a maximal left ideal of $T$ which is not an ideal of $T$. Since $K$ is a field which is contained in the center of $T$, we conclude that $K\subseteq R:=\mathbb{I}(M)$ and clearly $K\cap M=0$. Thus $K$ embeds in $R/M$ and hence we are done by Theorem 2.1. ∎
Hence we have the following immediate fact.
Corollary 2.3.
Let $T$ be an algebra over an infinite field $K$. Then either $T$ is a left quasi duo ring or $Max_{l}(T)\setminus Max(T)$ is infinite (in fact, $|Max_{l}(T)\setminus Max(T)|\geq|K|$).
Thus, if $T$ is a ring and $Max_{l}(T)\setminus Max(T)$ is finite, then for each $M\in Max_{l}(T)\setminus Max(T)$, we deduce that $\mathbb{I}(M)/M$ is a finite field. Moreover, if in addition $T$ is a $K$-algebra over a field $K$ and $Max_{l}(T)\setminus Max(T)$ is finite, then $K$ is a finite field (note that $\mathbb{I}(M)/M$ contains a copy of $K$, for each maximal left ideal $M$ of $T$).
By Theorem 2.1, note that if $Max_{l}(T)\setminus Max(T)=\bigcup_{i\in I}[M_{i}]$ be the partition of $Max_{l}(T)\setminus Max(T)$ in to similarity classes, then $|Max_{l}(T)\setminus Max(T)|\geq|I|+\sum_{i\in I}|\mathbb{I}(M)/M|$.
It is clear that if $K$ is a field and $n>1$, then by the above results we have $|Max_{l}(\mathbb{M}_{n}(K))|\geq|K|+1$ and $|Max_{r}(\mathbb{M}_{n}(K))|\geq|K|+1$ (note, $\mathbb{M}_{n}(K)$ is a simple $K$-algebra). In particular, if $K$ is an infinite field then $\mathbb{M}_{n}(K)$ has at least $|K|$-many maximal left/right ideals (also see Remark 2.5). Next we want to use the notation and results of [27], to show that if $R$ is an infinite division ring and $n>1$, then $T=\mathbb{M}_{n}(R)$ has infinitely many maximal left/right ideals. To see this, note that if $M$ is a maximal left ideal of a ring $R$ and $T=\mathbb{M}_{n}(R)$, then for each $u\in R^{n}\setminus M^{n}$, the set $D(M,u):=\{X\in T\ |\ Xu\in M^{n}\}$ is a maximal left ideal of $T$ (in fact, $T/D(M,u)\cong(R/M)^{n}$ as left $T$-module). Moreover, $Max_{l}(T)=\{D(M,u)\ |\ M\in Max_{l}(R),\ u\in R^{n}\setminus M^{n}\}$, see [27, Theorem 1.2]. In particular, if $R$ is a division ring, then we deduce that $Max_{l}(\mathbb{M}_{n}(R))=\{D(0,u)\ |\ 0\neq u\in R^{n}\}$.
Hence to show that for an infinite division ring $R$, $Max_{l}(\mathbb{M}_{n}(R))$ is infinite we must prove that there exist infinitely many $0\neq u\in R^{n}$ such that $D(0,u)$ are distinct maximal left ideals. By [27, Proposition 2.3], if $M$ is a maximal left ideal of a ring $R$, and $S=\mathbb{I}(M)$ be the idealizer of $M$ in $R$, and $u,v\in S^{n}$, then $D(M,u)=D(M,v)$ (in the ring $T=\mathbb{M}_{n}(R)$) if and only if there exists $c\in S\setminus M$ such that $v\equiv uc\ mod(M)$ (i.e. $v-uc\in M^{n}$). In particular, if $R$ is a division ring, since the only possible for $M$ is the zero ideal and in this case $S=R$, we deduce that for each $u,v\in R^{n}$, $D(0,u)=D(0,v)$ if and only if there exists $0\neq c\in R$ such that $v=uc$, (in other words $v$ and $u$ are right parallel vectors in $R^{n}$ as right $R$-module). Clearly, if $D$ is an infinite division ring and $n>1$, then $D^{n}$ has infinitely many right non-parallel vectors, thus we deduce the following fact.
Corollary 2.4.
Let $R$ be an infinite division ring and $n>1$, then $T=\mathbb{M}_{n}(R)$ has infinitely many maximal left (resp. right) ideals.
Remark 2.5.
Let $R$ be a ring, $n>1$ and $T=\mathbb{M}_{n}(R)$. Then it is obvious that $Max_{l}(T)\cap Max(T)=\emptyset=Max_{r}(T)\cap Max(T)$. Moreover, the function $M\longmapsto M^{t}$ (where $M^{t}$ is the transpose of $M$, i.e., $M^{t}:=\{m^{t}\ |\ m\in M\}$), is a one-one corresponding from $Max_{l}(T)$ onto $Max_{r}(T)$. In particular, $|Max_{l}(T)|=|Max_{r}(T)|$.
Remark 2.6.
Let $T$ be a ring, then we have the following immediate facts:
  1. (1)
    $Max_{l}(T)\cap Max(T)=Max_{r}(T)\cap Max(T)=Prm_{l}(T)\cap Max_{l}(T)=Prm_{r}(T)\cap Max_{r}(T)=Prm_{l}(T)\cap Max_{r}(T)=Prm_{r}(T)\cap Max_{l}(T)$.
  2. (2)
    $Max(T)\setminus Max_{l}(T)=Max(T)\setminus Max_{r}(T)$.
Next let us remind the reader some immediate fact from maximal subrings and also some needed result from [1]. It is easy to see that a proper subring $S$ of a ring $T$ is maximal if and only if for each $x\in T\setminus S$, we have $S[x]=T$. Moreover, if $R$ is a proper subring of $T$ and there exists $\alpha\in T$ such that $R[\alpha]=T$, then by a natural use of Zorn’s Lemma, $T$ has a maximal subring $S$ with $R\subseteq S$ and $\alpha\notin S$. In particular, if $R\subsetneq T$ is a ring extension which is finite as ring (module) over $R$, then $T$ has a maximal subring which contains $R$. Now let $R$ be a ring and $\Delta:=\{(r,r)\ |\ r\in R\}$, be the diagonal subring of $T:=R\times R$. Then it is not hard too see that $\Delta[(1,0)]=T$. Therefore $T$ has a maximal subring which contains $\Delta$. In fact, subrings of $T$ which contains $\Delta$ has the form $\Delta(I):=\Delta+I\times I$, where $I$ is an ideal of $R$. In particular, $\Delta(I)$ is a maximal subring of $T$ if and only if $I$ is a maximal ideal of $R$, see [1, Theorem 2.7]. Therefore a ring $R$ is simple if and only if $\Delta$ is a maximal subring of $R\times R$. In the following we give some results from [1].
Theorem 2.7.
  1. (1)
    [1, Theorem 4.1]. Let $T$ be a ring and $A$ be a maximal right/left ideal of $T$ which is not an ideal of $T$. Then the idealizer of $A$ is a maximal subring of $T$. In particular, either a ring has a maximal subring or is a quasi duo ring.
  2. (2)
    [1, Proposition 4.2]. Let $R$ be a maximal subring of a ring $T$ which contains a maximal one-sided ideal $A$ of $T$ which is not an ideal of $T$. Then $R$ is the idealizer of $A$ in $T$.
  3. (3)
    [1, Theorem 4.4]. Let $T$ be a ring which is not a division ring. If $T$ is left/right primitive ring (in particular, if $T$ is a simple ring), then $T$ has a maximal subring.
  4. (4)
    [1, Theorem 3.1]. For each ring $R$ and $n>1$, the matrix ring $\mathbb{M}_{n}(R)$ has a maximal subring.
Finally, we prove some needed facts as follows. We remind the reader that if $T$ is a ring with $|T|<\mathfrak{a}$, where $\mathfrak{a}$ is an infinite cardinal number, then for each finitely generated left/right $T$-module $M$, we have $|M|<\mathfrak{a}$. To see this, note that since $M$ is finitely generated as $T$-module, we infer that there exists a natural number $n$ such that $M$ is a homomorphic image of $T^{n}$ (as left/right $T$-module). Hence $|M|\leq|T|^{n}<\mathfrak{a}^{n}=\mathfrak{a}$. Now the following is in order.
Lemma 2.8.
Let $T$ be a left Noetherian ring and $I$ be a nilpotent ideal of $T$. If $\mathfrak{a}$ be an infinite cardinal number, then $|T|<\mathfrak{a}$ if and only if $|T/I|<\mathfrak{a}$.
Proof.
It is clear that, if $|T|<\mathfrak{a}$, then $|T/I|<\mathfrak{a}$ for each ideal $I$ of $T$. Conversely, assume that $I$ be a nilpotent ideal of $T$ and $|T/I|<\mathfrak{a}$. Since $I$ is a nilpotent ideal of $T$, we infer that there exists a natural number $n$ such that $I^{m}=0$. Now consider the chain $0=I^{m}\subseteq I^{m-1}\subseteq\cdots\subseteq I\subseteq T$. Since $T$ is a left Noetherian ring, we conclude that for each $i$, $I^{i}$ is a finitely generated left ideal. Hence for each $i$, the left $T$-module $I^{i}/I^{i+1}$, $0\leq i\leq m-1$ is finitely generated. Thus $|I^{i}/I^{i+1}|<\mathfrak{a}$. In particular, $|I^{m-1}|<\mathfrak{a}$. Therefore for each $i$ we deduce that $|I^{i}|<\mathfrak{a}$ and thus $|T|<\mathfrak{a}$. ∎
Note that in the proof of the previous lemma we only use the fact that $I$ is a nilpotent ideal of $T$ such that $I^{i}$ is a finitely generated left ideal of $T$, for each $i$. It is clear that by the condition of the previous lemma we conclude that $T$ is finite (resp. countable) if and only if $T/I$ is finite (resp. countable).
We conclude this section by a result about the structure of maximal subrings of a certain direct product of rings. If $A$ and $B$ be rings, then we say that $A$ and $B$ are homomorphically non-isomorphic rings, whenever for each proper ideals $I$ of $A$ and $J$ of $B$, the rings $A/I$ and $B/J$ are not isomorphic as ring. For example if $A=\mathbb{M}_{n_{1}}(R_{1})\times\cdots\times\mathbb{M}_{n_{k}}(R_{k})$, where $k\geq 1$, each $n_{i}\geq 2$ and $R_{i}$ are arbitrary rings and $B$ is any commutative ring, then $A$ and $B$ are homomorphically non-isomorphic rings (note that clearly, each homomorphic image of $A$ is a non-commutative ring). Now the following is in order.
Lemma 2.9.
Let $A$ and $B$ be homomorphically non-isomorphic rings and $T=A\times B$. Then $R$ is a maximal subring of $T$ if and only if either $R=A_{1}\times B$, where $A_{1}$ is a maximal subring of $A$, or $R=A\times B_{1}$, where $B_{1}$ is a maximal subring of $B$.
Proof.
It is clear that if $R$ is of the forms that is given in the statement of the lemma, then $R$ is a maximal subring of $T$. Conversely, let $R$ be a maximal subrings of $T$. If $0\times B\subseteq R$, then it is not hard to see that $R=A_{1}\times B$, where $A_{1}$ is a maximal subring of $A$. Similarly if $A\times 0\subseteq R$, then $R=A\times B_{1}$, where $B_{1}$ is a maximal subring of $B$. Hence assume that $R$ is a maximal subring of $T$ which does not contain $A\times 0$ and $0\times B$. Let $M=(A\times 0)\cap R=\{(a,0)\ |\ a\in A_{1}\}=A_{1}\times 0$, where $A_{1}$ is a proper subset of $A$ and $N=(0\times B)\cap R=\{(0,b)\ |\ b\in B_{1}\}=0\times B_{1}$, where $B_{1}$ is a proper subset of $B$. Clearly $M$ and $N$ are ideals of $R$ and $M+N=A_{1}\times B_{1}$. Also note that since $A\times 0\nsubseteq R$ and $R$ is a maximal subring of $T$, we have $R+(A\times 0)=T$ and therefore $R/M\cong B$ and similarly $R/N\cong A$. Now we have two cases (note $(1,1)\in R$): $(i)$ $(1,0)\in R$ which equivalent to $(0,1)\in R$. Therefore in this case $M+N=R$ and thus $R=A_{1}\times B_{1}$. Hence $A_{1}$ and $B_{1}$ are proper subrings of $A$ and $B$, respectively. Clearly $R=A_{1}\times B_{1}\subsetneq A\times B_{1}\subsetneq T$, which contradicts to the maximality of $R$. $(ii)$ hence assume that $(1,0)$ and $(0,1)$ are not in $R$. Thus $(1,1)\notin M+N$ and therefore $M+N$ is a proper ideal of $R$ which contains $M$ and $N$. Hence, by $R/M\cong B$ and $R/N\cong A$, we deduce that $A$ and $B$ have epimorphic images that are isomorphic as ring which is impossible. Thus we are done. ∎
3. Rings with Infinitely Many Maximal Subrings
In this section we study the number of maximal subrings for general non-commutative rings, especially for division rings, simple rings and also $K$-algebras over an infinite field $K$. If $E$ is a field with prime subfield $F$, then by [10, Corollary 1.5], whenever $E$ has only finitely many maximal subrings, then $F\cong\mathbb{Z}_{p}$ for some prime number $p$ and $E$ is algebraic over $F$ (for characterization of such fields see [2]). Moreover, if $E$ is not an absolutely algebraic field, then $|RgMax(E)|\geq max\{tr.deg(E/F),|E|\}$, by [10, Corollary 1.5]. For division rings we have the next result.
Proposition 3.1.
Let $D$ be a non-commutative division ring. If $D$ has a maximal subring $R$ which is a division ring, then $D$ has infinitely many maximal subrings. Consequently, if $D$ has only finitely many maximal subrings, then each $x\in D\setminus C_{D}(D)$ is not algebraic over $C_{D}(D)$ (i.e., $C_{D}(D)$ is algebraically closed in $D$) and $C_{D}(R)=C_{D}(D)$ for each maximal subring $R$ of $D$.
Proof.
Assume that $R$ be a maximal subring of $D$ which is a division ring. Then by [17, Theorem 3], there exist infinitely many subrings of the form $xRx^{-1}$, where $x\in D\setminus\{0\}$. It is evident that for each $0\neq x\in D$, $xRx^{-1}$ is a maximal subring of $D$. Thus $D$ has infinitely many maximal subrings. For the next part, assume that $D$ has only finitely many maximal subrings, then by the first part we conclude that each maximal subring of $D$ is not a division ring. Now if $x$ is a non-centeral element of $D$ which is algebraic over the center of $D$, then by [5, Corollary 2.10], $D$ has a maximal subring which is a division ring which is impossible. Hence the center of $D$ is algebraically closed in $D$. For the final part note that if $R$ is a maximal subring of $D$ and $\alpha\in C_{D}(R)$ is not a central element of $D$, then clearly $R=C_{D}(\alpha)$, i.e., $R$ is a division ring which is impossible too. Thus $C_{D}(R)=C_{D}(D)$. ∎
Corollary 3.2.
Let $D$ be an infinite division ring with $C_{D}(D)=F$. Then either $D$ has infinitely many maximal subrings or $D:F]\geq|F|$ (in fact, for each $x\in D\setminus F$ we have $F(x):F]\geq|F|$).
Proof.
Let $x\in D\setminus F$ and $[F(x):F]<|F|$. Then similar to the proof of [21, Theorem 4.20], we deduce that $F(x)$ is algebraic over $F$. Hence $x$ is algebraic over $F$ and thus we are done by the previous result. ∎
Corollary 3.3.
Let $D$ be an existentially complete non-commutative division ring over a field $K$. Then $D$ has infinitely many maximal subrings.
Proof.
By [5, $(1)$ of Proposition 3.5], $D$ has a maximal subring of the form $R=C_{D}(x)$ for some $x\in D$. It is clear that $R$ is a division ring and hence we are done by Proposition 3.1. ∎
For more results first we need to characterize the maximal subrings of $R\times R$ for a simple ring $R$, as follow.
Theorem 3.4.
Let $R$ be a simple ring. Then $S$ is a maximal subring of $T:=R\times R$ if and only if $S$ satisfies exactly one of the following conditions:
  1. (1)
    $S=A\times R$ or $S=R\times A$, where $A$ is a maximal subring of $R$.
  2. (2)
    There exist $\sigma_{1},\sigma_{2}\in Aut(R)$ such that $S=\{(\sigma_{1}(x),\sigma_{2}(x))\ |\ x\in R\}$.
Moreover in case $(2)$, we have $S\cong R$.
Proof.
First assume that $S$ is a subring of $T$ which is of the forms given in $(1)$ or $(2)$. It is clear that if $S$ is of the forms given in $(1)$, then $S$ is a maximal subring of $T$. Hence assume that $S$ is of the form in $(2)$. Define $\sigma:R\longrightarrow S$, by $\sigma(x)=(\sigma_{1}(x),\sigma_{2}(x))$, then one can easily check that $\sigma$ is an isomorphism and therefore $S\cong R$ is a simple ring. To show $S$ is a maximal subring of $T$, we first prove that $S[(1,0)]=T$. Clearly $(0,1)\in S[(1,0)]$, for $(1,1)\in S$. Now let $(a,b)\in T=R\times R$, thus there exist $x,y\in R$ such that $\sigma_{1}(x)=a$ and $\sigma_{2}(y)=b$, for $\sigma_{1},\sigma_{2}\in Aut(R)$. Thus by definition of $S$ we conclude that $(\sigma_{1}(x),\sigma_{2}(x))=(a,\sigma_{2}(x))\in S$ and $(\sigma_{1}(y),\sigma_{2}(y))=(\sigma_{1}(y),b)\in S$. Therefore $(a,b)=(a,\sigma_{2}(x))(1,0)+(\sigma_{1}(y),b)(0,1)\in S[(1,0)]$. Thus $S[(1,0)]=T$. Now we prove that for each $(x,y)\in T\setminus S$, we have $S[(x,y)]=T$, which shows that $S$ is a maximal subring of $T$. Since $\sigma_{2}\in Aut(R)$, there exists $c\in R$ such that $\sigma_{2}(c)=y$. Therefore $(\sigma_{1}(c),y)\in S\subseteq S[(x,y)]$ and therefore $(x-\sigma_{1}(c),0)\in S[(x,y)]$. Since $(x,y)\notin S$, we infer that $x_{0}:=x-\sigma_{1}(c)\neq 0$. Thus $Rx_{0}R=R$, for $R$ is a simple ring. Hence there exist $n\in\mathbb{N}$ and $r_{i},s_{i}\in R$, $1\leq i\leq n$, such that $\sum_{i=1}^{n}r_{i}x_{0}s_{i}=1$. From $\sigma_{1}\in Aut(R)$, we deduce that there exist $a_{i},b_{i}\in R$, $1\leq i\leq n$, such that $\sigma_{1}(a_{i})=r_{i}$ and $\sigma_{1}(b_{i})=s_{i}$. Therefore
$(1,0)=\sum_{i=1}^{n}(\sigma_{1}(a_{i}),\sigma_{2}(a_{i}))(x_{0},0)(\sigma_{1}(b_{i}),\sigma_{2}(b_{i}))\in S[(x,y)]$
Thus $T=S[(1,0)]\subseteq S[(x,y)]$ and therefore $S[(x,y)]=T$. Hence $S$ is a maximal subring of $T$.
Conversely, assume that $S$ be a maximal subring of $T$. Suppose $I=R\times 0$ and $J=0\times R$. If either $I\subseteq S$ or $J\subseteq S$, then $S$ is of the forms which is given in $(1)$. Thus assume that $I$ and $J$ are not contained in $S$. Therefore $I_{1}=I\cap S=\{(x,o)\ |\ x\in S_{1}\}=S_{1}\times 0$ and $J_{1}=J\cap S=\{(0,y)\ |\ y\in S_{2}\}=0\times S_{2}$, where $S_{1}$ and $S_{2}$ are proper subsets of $R$. Clearly $I_{1}+J_{1}=S_{1}\times S_{2}\subseteq S$. Also note that since $S$ is a maximal subring of $T$ which does not contain $I$ and $J$, we deduce that $S+I=T=S+J$ and thus $S/I_{1}\cong R\cong S/J_{1}$. Thus $I_{1}$ and $J_{1}$ are maximal ideals of $S$. Now we have two cases: $(a)$ if $(1,0)\in S$ (which is equivalent to $(0,1)\in S$), then $(1,1)=(1,0)+(0,1)\in I_{1}+J_{1}$, which means $I_{1}+J_{1}=S$ and therefore $S=S_{1}\times S_{2}$. Thus $S_{1}$ and $S_{2}$ are proper subring of $R$. But clearly $S\subsetneq R\times S_{2}\neq T$, which contradicts the maximality of $S$. $(b)$ Hence $(1,0),(1,0)\notin S$. Thus $I_{1}+J_{1}=S_{1}\times S_{2}$ is a proper ideal of $S$ which contains $I_{1}$ and $J_{1}$. But as we see earlier, $I_{1}$ and $J_{1}$ are maximal ideals of $S$. Therefore we immediately conclude that $I_{1}=J_{1}$. Thus $I_{1}=J_{1}=0$ and therefore $S\cong R$. Let $\sigma:R\longrightarrow S$ be a ring isomorphism, and for each $x\in R$, let $\sigma(x)=(\sigma_{1}(x),\sigma_{2}(x))$. Then clearly $\sigma_{1}$ and $\sigma_{2}$ are ring endomorphisms of $R$. Since $R$ is a simple ring we conclude that $\sigma_{1}$ and $\sigma_{2}$ are one-one. Finally note that since $S=\{(\sigma_{1}(x),\sigma_{2}(x))\ |\ x\in R\}$ and $(1,0)\notin S$, we deduce that $S[(1,0)]=T$. This immediately implies that $T=S+S(1,0)$ (note, $(1,0)$ is a central idempotent of $T$). To show that $\sigma_{1}$ and $\sigma_{2}$ are onto, let $x\in R$, then $(x,0)\in S+S(1,0)$. Therefore there exist $a,b\in R$ such that $(x,0)=(\sigma_{1}(a),\sigma_{2}(a))+(\sigma_{1}(b),\sigma_{2}(b))(1,0)$, which immediately shows that $\sigma_{1}(b)=x$ and therefore $\sigma_{1}$ is onto. Similarly, $\sigma_{2}$ is onto and hence we are done. ∎
Next we have the following main results in this paper for simple rings.
Theorem 3.5.
Let $T$ be a simple ring which is not a division ring. Then $T$ has only finitely many maximal subring if and only if $R$ is finite (i.e., $T\cong\mathbb{M}_{n}(F)$, where $F$ is a finite field and $n\geq 2$)
Proof.
First we show that $Max_{l}(T)$ is finite. For otherwise let $M_{1},M_{2},\ldots$ be distinct maximal left ideals of $T$. Then for each $i$, by $(1)$ of Theorem 2.7, $A_{i}=\mathbb{I}_{T}(M_{i})$ is a maximal subring of $T$ which contains $M_{i}$ (note, for each $i$, $M_{i}$ is not an ideal of $T$). Since for $i\neq j$, $M_{i}+M_{j}=T$, we immediately conclude that $A_{i}\neq A_{j}$. Thus $T$ has infinitely many maximal subrings which is absurd. Thus $Max_{l}(T)=\{M_{1},\ldots,M_{n}\}$. Now since $J(T)=0$, we deduce that $M_{1}\cap\cdots\cap M_{n}=0$. Thus $T$ can be embedded in $T/M_{1}\times\cdots\times T/M_{n}$. Hence $T$ is a left Artinian ring. Therefore $T\cong\mathbb{M}_{n}(D)$ for some division ring $D$ and $n\geq 2$ (note, $T$ is not a division ring). If $D$ is an infinite division ring, then $\mathbb{M}_{n}(D)$, where $n\geq 2$, has infinitely many maximal left (right) ideals, by Corollary 2.4. Thus $D$ is finite and therefore $D$ is a finite field. Hence $T$ is a finite ring. ∎
We remind the reader that if $K$ is a field, then $K\times K$ has only finitely many maximal subrings if and only if $K$ is finite, see [10, Corollary 3.5]. Now the following generalization is in order.
Proposition 3.6.
Let $R$ be a simple ring. Then $R\times R$ has finitely many maximal subrings if and only if $R$ is finite.
Proof.
It is clear that if $R$ is finite then $R\times R$ has only finitely many maximal subrings. Conversely, assume that $R\times R$ has only finitely many maximal subrings. Thus $R$ has only finitely many maximal subrings. If $R$ is not a division ring, then we are done by the previous theorem. Hence assume that $R$ is a division ring and $R\times R$ has only finitely many maximal subrings. By the preceding comment we may assume that $R$ is not a field and therefore $C_{R}(R)\subsetneq R$. Now suppose that $R$ is infinite. Then we have two cases: $(a)$ If there exists a non-central element $x$ of $R$ which is algebraic over the center of $R$, then by Proposition 3.1, $R$ has infinitely many maximal subrings. Thus $R\times R$ has infinitely many maximal subring too, which is impossible. $(b)$ Hence assume that each non-central element $x$ of $R$ is not algebraic over the center of $R$. For each $n\geq 1$ and $a\in D$, let $\sigma_{n}(a)=x^{n}ax^{-n}$ be the inner automorphism of $R$ induced from $x^{n}$. If $\sigma_{n}=\sigma_{m}$ for $n>m$, then $\sigma_{1}^{n-m}=1$ and therefore $x^{n-m}$ is a central element of $R$. This immediately implies that $x$ is integral over the center of $R$, which contradict to the assumption in this case. Thus $\sigma_{n}\neq\sigma_{m}$ for each $n\neq m$. Finally note that by Theorem 3.4, for each $n\geq 1$, the set $S_{n}=\{(x,\sigma_{n}(x))\ |\ x\in R\}$ is a maximal subring of $R\times R$ and clearly for $n\neq m$, we have $S_{n}\neq S_{m}$. Therefore $R\times R$ has infinitely many maximal subrings, which is absurd. Thus $R$ is a finite and hence we are done. ∎
By the previous results we have the following immediate results.
Corollary 3.7.
Let $D$ be a division ring. Then $D\times D$ has only finitely many maximal subrings if and only if $D$ is a finite field.
For definition of existentially complete division rings see [12].
Proposition 3.8.
Let $D$ be an infinite countable existentially complete division ring, then $D\times D$ has at least $2^{\aleph_{0}}$ maximal subrings.
Proof.
By [12, Theorem 6.5.11], $D$ has at least $2^{\aleph_{0}}$ automorphisms. Hence by Theorem 3.4 for each $\sigma\in Aut(D)$, $S_{\sigma}:=\{(x,\sigma(x))\ |\ x\in D\}$ is a maximal subring of $D\times D$. It is clear that $S_{\sigma}\neq S_{\sigma^{\prime}}$, whenever $\sigma$ and $\sigma^{\prime}$ are two distinct automorphisms of $D$. Hence $D\times D$ has at least $2^{\aleph_{0}}$ maximal subrings and hence we are done. ∎
Corollary 3.9.
Let $T=R_{1}\times R_{2}$, where $R_{1}$ and $R_{2}$ be two simple ring. Then $T$ has infinitely many maximal subrings if and only if one of the following holds:
  1. (1)
    $R_{1}$ or $R_{2}$ has infinitely many maximal subrings.
  2. (2)
    $R_{1}\cong R_{2}$ are infinite rings.
Proof.
It is clear that if $(1)$ holds then $T$ has infinitely many maximal subrings. Now assume that $(2)$ holds, thus $T\cong R_{1}\times R_{1}$ and therefore by Proposition 3.6, we conclude that $T$ has infinitely many maximal subrings. Conversely, assume that $T$ has infinitely many maximal subrings, we prove that $(1)$ or $(2)$ holds. Suppose that $(2)$ does not hold, i.e., $R_{1}$ and $R_{2}$ are not isomorphic rings. Thus $R_{1}$ and $R_{2}$ are homomorphically non-isomorphic rings. Therefore by Lemma 2.9, we infer that each maximal subrings of $T$ is of the form $S_{1}\times R_{2}$ or $R_{1}\times S_{2}$, where $S_{i}$ is a maximal subring of $R_{i}$. Since $T$ has infinitely many maximal subrings, we immediately conclude that either $R_{1}$ or $R_{2}$ has infinitely many maximal subrings and hence we are done. ∎
For a ring $T$, let us denote the sets $Max_{l}(T)\setminus Max(T)$ by $Max_{l}^{\prime}(T)$, similarly $Max_{r}^{\prime}(T):=Max_{r}(T)\setminus Max(T)$ and $Max^{\prime}(T):=Max(T)\setminus Max_{l}(T)=Max(T)\setminus Max_{r}(T)$. Note that if $P$ is a prime ideal of a ring $T$ and $A$ be a left (resp. right) ideal and $J$ be an ideal of $T$ such that $P=I\cap J$, then either $P=I$ or $P=J$. To see this note that clearly $JI\subseteq I\cap J\subseteq P$ (resp. $IJ\subseteq I\cap J\subseteq P$), therefore $I\subseteq P$ or $J\subseteq P$. Hence $P=I$ or $P=J$.
Proposition 3.10.
Let $T$ be a ring with only finitely many maximal subrings. Then the following hold:
  1. (1)
    $Max_{l}^{\prime}(T)$ and $Max_{r}^{\prime}(T)$ are finite sets and in fact $|Max_{l}^{\prime}(T)|\leq|RgMax(T)|$ and $|Max_{r}^{\prime}(T)|\leq|RgMax(T)|$.
  2. (2)
    for each $M\in Max_{l}^{\prime}(T)\cup Max_{r}^{\prime}(T)$, the ring $\mathbb{I}(M)/M$ is a finite field. In particular, $C(T)/(C(T)\cap M)$ is a finite field.
  3. (3)
    for each $M\in Max_{l}(T)\cap Max(T)=Max_{r}(T)\cap Max(T)$, the center of $T/M$ is algebraically closed in $T/M$.
  4. (4)
    for each $M\in Max^{\prime}(T)$, there exist a natural number $n>1$ and a finite field $F$ such that $T/M\cong\mathbb{M}_{n}(F)$.
  5. (5)
    $Max^{\prime}(T)$ is finite and in fact $|Max^{\prime}(T)|\leq|RgMax(T)|$.
  6. (6)
    $Prm_{l}(T)\setminus Max_{l}(T)$, $Prm_{l}(T)\setminus Max(T)$, $Prm_{r}(T)\setminus Max_{r}(T)$ and $Prm_{r}(T)\setminus Max(T)$ are finite sets.
  7. (7)
    If $P$ a left (resp. right) primitive ideal of $T$, then either $P$ is a maximal ideal of $T$ or $P=\bigcap_{i\in I}M_{i}$, where $M_{i}\in Max_{l}(T)\cap Max(T)$, moreover in this case $I$ is infinite and $T/P$ is a domain (i.e., $P$ is a completely prime ideal of $T$). In particular, $T/P$ is embeddable in a division ring.
  8. (8)
    $J(T)=\bigcap_{M\in Max(T)}M$.
Proof.
$(1)$ for each maximal left ideal $A$ of $T$ which is not an ideal of $T$, by Theorem 2.7, $\mathbb{I}(A)$ is a maximal subring of $T$ which contains $A$. It is clear that whenever $A\neq B$ are in $Max_{l}^{\prime}(T)$, then $A+B=T$. Hence we conclude that $\mathbb{I}(A)\neq\mathbb{I}(B)$. Therefore the function $A\longmapsto\mathbb{I}(A)$, from $Max_{l}^{\prime}(T)$ in to $RgMax(T)$ is one-one and therefore $(1)$ holds. For $Max_{r}^{\prime}(T)$ the proof is similar.
$(2)$ is an immediate consequence of Theorem 2.1 and $(1)$ (note that if $M\in Max_{l}^{\prime}(T)$, then $\mathbb{I}(M)/M\cong(End({}_{T}(T/M)))^{op}$ is a division ring). For the final part of $(2)$ note that $C(T)+M$ is a subring of $T$ which is contained in $\mathbb{I}(M)$ and therefore $(C(T)+M)/M$ is a subring of $\mathbb{I}(M)/M$. Now by the first part of $(2)$, we immediately deduce that $(C(T)+M)/M$ is a finite field and thus $C(T)/(C(T)\cap M)$ is a finite field too.
$(3)$ if $T/M$ is a field (in particular, if $T/M$ is finite), then the conclusion is obvious. Hence assume that $T/M$ is an infinite division ring, then by Theorem 3.1, we deduce that the center of $T/M$ is algebraically closed in $T/M$ (for otherwise, $T/M$ and therefore $T$ have infinitely many maximal subrings, which is absurd).
$(4)$ this is an immediate conclusion of Theorem 3.5 and the fact that $T/M$ has only finitely many maximal subrings (for $T$ has only finitely many maximal subrings).
$(5)$ assume that $Max^{\prime}(T)=\{M_{i}\ |\ i\in I\}$, where $M_{i}\neq M_{j}$ for $i\neq j\in I$. Then by $(4)$, for each $i\in I$ there exist a natural number $n_{i}>1$ and a finite field $F_{i}$ such that $T/M_{i}\cong\mathbb{M}_{n_{i}}(F_{i})$. Therefore by $(4)$ of Theorem 2.7, we infer that for each $i\in I$, $M_{i}$ is contained in a maximal subring $S_{i}$ of $T$. Since for $i\neq j$, we have $M_{i}+M_{j}=T$, we deduce that $S_{i}\neq S_{j}$. Therefore $|I|\leq|RgMax(T)|$ and hence we are done.
$(6)$ let $\{P_{i}\ |\ i\in I\}$ be a family of distinct left primitive ideals of $T$ such that for each $i\in I$, $P_{i}$ is not a maximal (left) ideal of $T$. Since for each $i\in I$, $P_{i}$ is a left primitive ideal of $T$, we deduce that there exists a maximal left ideal $M_{i}$ of $T$ such that $P_{i}=l.ann_{T}(T/M_{i})$. It is clear that whenever $i\neq j$, then $M_{i}\neq M_{j}$, for $P_{i}\neq P_{j}$. Also note that if $M_{i}$ is an ideal of $T$, then $P_{i}=M_{i}$ is a maximal (left) ideal of $T$ which is absurd. Thus for each $i\in I$ we deduce that $M_{i}\in Max_{l}^{\prime}(T)$ and therefore by $(1)$ we conclude that $I$ is finite. Thus the first and the second parts of $(6)$ hold. The proof for the next two parts are similar.
$(7)$ let $P$ be a left primitive ideal of $T$. Hence $P$ is an intersection of a family of maximal left ideals of $T$. Since by $(1)$, $Max_{l}^{\prime}(T)$ is finite, we conclude that there exist $N_{1},\ldots,N_{k}\in Max_{l}^{\prime}(T)$, $k\geq 0$, and a family $\{M_{i}\ |\ i\in I\}\subseteq Max_{l}(T)\cap Max(T)$ such that $P=N_{1}\cap\cdots N_{k}\cap\bigcap_{i\in I}M_{i}$. By the comments preceding the proposition we infer that either $P=N_{1}\cap\cdots N_{k}$ or $P=\bigcap_{i\in I}M_{i}$. If $P=N_{1}\cap\cdots N_{k}$, then the ring $T/P$ embeds in the semisimple left $T$-module $T/N_{1}\times\cdots\times T/N_{k}$, as left $T$-module and therefore $T/P$ is a prime semisimple ring. Thus $T/P$ is a simple ring, i.e., $P$ is a maximal ideal of $T$. Hence assume that $P$ is not a maximal ideal of $T$ and $P=\bigcap_{i\in I}M_{i}$. Since $P$ is not maximal ideal of $T$, but is a prime ideal of $T$, we immediately conclude that $I$ is infinite. Finally note that for each $i\in I$, the ring $T/M_{i}$ is a division ring. Therefore $\prod_{i\in I}T/M_{i}$ is a reduced ring. Since $T/P$ embeds in $\prod_{i\in I}T/M_{i}$, we deduce that $T/P$ is a reduced ring. Thus $T/P$ is a reduced prime ring and hence is a domain. Therefore by [12, Theorem 1.2.4], $T/P$ is embeddable in a division ring.
$(8)$ This is trivial by $(7)$ and the fact that $J(T)$ is the intersection of all left (resp. right) primitive ideals of $T$. ∎
Corollary 3.11.
Let $T$ be an infinite left/right primitive ring which is not embeddable in a division ring, then $T$ has infinitely many maximal subrings.
Proof.
If $T$ is a simple ring then we are done by Theorem 3.5. Hence assume that $T$ is not a simple ring. If $T$ has only finitely many maximal subrings, then by $(7)$ of Proposition 3.10, we conclude that $T$ is embeddable in a division ring which is absurd. Thus $T$ has infinitely many maximal subrings. ∎
Corollary 3.12.
Let $T$ be an infinite prime ring which is not embeddable in a division ring. If $J(T)=0$, then $T$ has infinitely many maximal subrings.
Proof.
Assume that $T$ has only finitely many maximal subrings, then by $(8)$ and $(5)$ of Proposition 3.10, we conclude that $0=J(T)=Q_{1}\cap\cdots\cap Q_{n}\cap N$, where $Q_{i}$ are maximal ideals of $T$ which are not maximal left ideals, $n\geq 0$ and $N=\bigcap_{i\in I}M_{i}$, where $Max(T)\cap Max_{l}(T)=\{M_{i}\ |\ i\in I\}$. Since $T$ is a prime ring, we infer that either there exists $k$ such that $Q_{k}=0$, i.e., $T$ is an infinite simple ring, or $N=0$. If $T$ is an infinite simple ring, then by hypothesise and Theorem 3.5, we deduce that $T$ has infinitely many maximal subrings which is absurd. Thus assume that $N=0$. Therefore $T$ embeds in $\prod_{i\in I}T/M_{i}$, and similar to the proof of $(7)$, we conclude that $T$ is embeddable in a division ring which is absurd. Thus $T$ has infinitely many maximal subrings. ∎
We need the following main lemma for more conclusions.
Lemma 3.13.
Let $T$ be an infinite ring and $\{M_{i}\ |\ i\in I\}\subseteq Max_{l}(T)\cap Max(T)$ such that $\bigcap_{i\in I}M_{i}=0$. Then either $T$ has infinitely many maximal subring or $C(T)$ is integrally closed in $T$. In particular, if $T$ is a (left) quasi duo $J$-semisimple ring, then either $T$ has infinitely many maximal subring or $C(T)$ is integrally closed in $T$.
Proof.
Assume that $T$ has only finitely many maximal subrings. Hence by Theorem 3.1, for each $i\in I$, the center of $T/M_{i}$ is integrally closed in $T/M_{i}$, for otherwise if there exists $i\in I$ such that $T/M_{i}$ has a non-central element which is integral over the center of $T/M_{i}$, then $T/M_{i}$ and therefore $T$ have infinitely many maximal subrings, which is absurd. Now assume that $\alpha$ be a non-central element of $T$ which is integral over the center of $T$. Thus there exists $\beta\in T$ such that $\alpha\beta-\beta\alpha\neq 0$. Since $\bigcap_{i\in I}M_{i}=0$, we infer that there exists $i\in I$ such that $\alpha\beta-\beta\alpha\notin M_{i}$. Therefore $\alpha+M$ is a non-central element of the division ring $T/M_{i}$. Clearly, $\alpha+M_{i}$ is integral over the center of $T/M_{i}$, which is not possible. Thus we conclude that the center of $T$ is integrally closed in $T$. The final part is evident. ∎
Remark 3.14.
Let $T$ be a ring with only finitely many maximal subrings. Then $Max_{l}(T)$ is finite if and only if $T/J(T)$ is finite if and only if $Max_{r}(T)$ is finite. To see this, it suffices to show that if $Max_{l}(T)$ is finite, then $T/J(T)$ is finite. Since $Max_{l}(T)$ is finite, we immediately conclude that $T/J(T)$ is a semisimple ring. By assumption, $T$ and therefore $T/J(T)$ has only finitely many maximal subrings, therefore by Proposition 3.6, we deduce that each simple part of $T/J(T)$ is finite and therefore $T/J(T)$ is a finite ring.
For the next result, by the previous remark, if $T$ is a ring with only finitely many maximal subrings, then we may assume that $Max_{l}(T)$ is infinite and therefore by Proposition 3.10, we conclude that $Max_{l}(T)\cap Max(T)$ is infinite.
Corollary 3.15.
Let $T$ be an infinite ring with only finitely many maximal subrings. Then there exist finitely many distinct maximal ideals $Q_{1},\ldots,Q_{m}$, $m\geq 0$, such that:
  1. (1)
    for each $i$, there exist a natural number $n_{i}>1$ and a finite field $F_{i}$ such that $T/Q_{i}\cong\mathbb{M}_{n_{i}}(F_{i})$.
  2. (2)
    $J(T)=Q_{1}\cap\cdots\cap Q_{m}\cap N$, where $N=\bigcap_{M\in Max(T)\cap Max_{l}(T)}M$, where $N,Q_{1},\ldots,Q_{m}$ are incomparable.
  3. (3)
    $T/J(T)\cong R\times S$, where $R$ is a finite semisimple ring (in fact, $R=\prod_{i=1}^{n}\mathbb{M}_{n_{i}}(F_{i})$) and the center of $S$ is integrally closed in $S$.
Proof.
First note that by $(8)$ of Proposition 3.10, $J(T)=\bigcap_{M\in Max(T)}M$. Since by $(5)$ of Proposition 3.10, $Max^{\prime}(T)$ is finite we deduce that there exists finitely many distinct maximal ideals $Q_{1},\ldots,Q_{m}\in Max^{\prime}(T)$, $m\geq 0$, such that $J(T)=Q_{1}\cap\cdots\cap Q_{m}\cap N$, where $N=\bigcap_{M\in Max(T)\cap Max_{l}(T)}M$ and clearly we may assume that $N$ and $Q_{i}$ are incomparable. Thus $(2)$ holds and note that $(1)$ is obvious by $(4)$ of Proposition 3.10. Since $Q_{i}$ are distinct maximal ideals which are incomparable to $N$, and $J(T)=Q_{1}\cap\cdots\cap Q_{m}\cap N$, we deduce the first part of $(3)$ (where $S=T/N$). The final part of $(3)$ is a consequence of Lemma 3.13. Hence we are done. ∎
Now we have the following main result.
Theorem 3.16.
Let $T$ be a $K$-algebra over an infinite field $K$. Then either $T$ has infinitely many maximal subrings (in fact, $|RgMax(T)|\geq|K|$) or the following hold:
  1. (1)
    $T$ is a quasi duo ring. In particular, $Max(T)=Max_{l}(T)=Max_{r}(T)=Prm_{l}(T)=Prm_{r}(T)$.
  2. (2)
    For each $M\in Max(T)$, the center of $T/M$ is integrally closed in $T/M$.
  3. (3)
    $J(T)=\bigcap_{M\in Max(T)}M$ and $T/J(T)$ is a reduced ring. In particular, $N(T)\subseteq J(T)$. Moreover, the center of $T/J(T)$ is integrally closed in $T/J(T)$.
Proof.
If $T$ has a maximal left ideal $M$, which is not an ideal of $T$, then by Corollary 2.2, we deduce that $|K|\leq|Max_{l}(T)\setminus Max(T)|$. Thus by $(1)$ of Proposition 3.10, we conclude that $|K|\leq|RgMax(T)|$. Hence if $|RgMax(T)|<|K|$ we deduce that each maximal left ideal of $T$ is an ideal of $T$, i.e., $T$ is a left quasi duo ring, and by a similar proof $T$ is a right quasi duo ring. Hence $T$ is a quasi duo ring. Thus $(1)$ holds. $(2)$ is an immediate consequence of Proposition 3.1. For $(3)$, by $(1)$ we infer that $J(T)$ is the intersection of all maximal ideals of $T$ and therefore $T/J(T)$ embeds in the reduce ring $\prod_{M\in Max(T)}T/M$ (note that $T$ is a quasi duo ring and therefore $T/M$ is a division ring for each $M\in Max(T)$). Hence $T/J(T)$ is a reduced ring and thus $N(T)\subseteq J(T)$. The final part of $(3)$ is obvious by Lemma 3.13 (note, $T/J(T)$ is an infinite ring, for it contains a copy of the infinite field $K$). ∎
Corollary 3.17.
Let $T$ be a prime $K$-algebra over an infinite field $K$ with $J(T)=0$, then either $|RgMax(T)|\geq|K|$ or $T$ embeds in a division ring (therefore $T$ is a domain).
Proof.
Assume that $|RgMax(T)|<|K|$, then by the proof of the previous theorem $T$ embeds in the reduced ring $R:=\prod_{M\in Max(T)}T/M$. Thus $T$ is a reduce prime ring. Therefore $T$ is a domain. Since for each maximal ideal $M$ of $T$, $T/M$ is a division ring, we infer that $T$ is a strongly von Neumann regular ring. Hence by [12, Theorem 1.2.4], we conclude that $T$ is embeddable in a division ring (note, in fact $T$ is a division ring, for $T$ is a strongly von Neumann regular domain and therefore $C(T)$ is integrally closed in $T$, by Proposition 3.1). ∎
If $R$ is a commutative ring and $M$ be a maximal ideal of $R$ such that $Char(R/M)=0$, then $R/M$ is a field with infinitely many maximal subrings, see [10, Corollary 1.5]. Therefore $R$ has infinitely many maximal subrings. For non-commutative rings we have the following result.
Corollary 3.18.
Let $T$ be a ring with only finitely many maximal subrings. Then the following hold:
  1. (1)
    For each $M\in Max_{l}^{\prime}(T)\cup Max_{r}^{\prime}(T)$, there exists a prime number $p$ such that $pT\subseteq M$.
  2. (2)
    For each $M\in Max^{\prime}(T)$, there exists a prime number $p$ such that $pT\subseteq M$.
  3. (3)
    If $P$ is a left/right primitive ideal of $T$, then either $P$ is a maximal left and right ideal of $T$ or $Char(T/P)$ is a prime number.
Proof.
$(1)$ Let $M$ be a maximal left/right ideal of $T$ which is not an ideal of $T$. Since $T$ has only finitely many maximal subring, by $(2)$ of Proposition 3.10, we conclude that $\mathbb{I}(M)/M$ is a finite field. Hence $Char(\mathbb{I}(M)/M)=p$, where $p$ is a prime number. Thus $p.1\in M$, for $1\in\mathbb{I}(M)$. Therefore $pT\subseteq M$.
$(2)$ The proof is similar to $(1)$ and using $(4)$ of Proposition 3.10.
$(3)$ Let $P$ be a left primitive ideal of $T$, which is not a left/right maximal ideal of $T$. Hence there exists a maximal left ideal $M$ of $T$ such that $P=l.ann_{T}(T/M)$. If $M$ is an ideal of $T$, then $M=l.ann_{T}(T/M)=P$, which is absurd. Thus $M$ is not an ideal of $T$ and therefore by $(1)$, we deduce that there exists a prime number $p$ such that $p(T/M)=0$. Hence $p\in P$ and thus $Char(T/P)=p$. ∎
By the previous result, if $T$ is a ring with zero characteristic and there exists a maximal left/right ideal $M$ of $T$ which is not an ideal of $T$ and either $M\cap\mathbb{Z}=0$ or $T/M$ is a torsionfree/faithful $\mathbb{Z}$-module, then $T$ has infinitely many maximal subrings.
Proposition 3.19.
Let $T$ be an infinite ring with nonzero characteristic. If $Max_{F}(T):=\{M\in Max(T)\ |\ T/M\ \text{is}\ \text{finite}\}$ is infinite, then $T$ has infinitely many maximal subrings.
Proof.
Since $n$ has only finitely many prime divisors and for each $M\in Max_{F}(T)$, $Char(T/M)$ is a prime divisor of $n$, we conclude that the exists a prime divisor $p$ of $n$ such that $Max_{p,F}(T):=\{M\in Max_{F}(T)\ |\ Char(T/M)=p\}$ is infinite. Also note that for each $M\in Max_{p,F}(T)$, the ring $T/M$ is a finite simple ring, hence we deduce that there exist a natural number $m$ and a finite field $E$ such that $T/M\cong\mathbb{M}_{m}(E)$. Now we have two cases:
$(1)$ there exist distinct $M_{1},M_{2},\ldots\in Max_{p,F}(T)$ such that $T/M_{i}\cong\mathbb{M}_{m_{i}}(E_{i})$, where $m_{i}>1$ and $E_{i}$ is a field. Thus in this case for each $i$, $T$ has a maximal subring $R_{i}$ which contains $M_{i}$, for $\mathbb{M}_{m_{i}}(E_{i})$ has a maximal subring by $(4)$ of Theorem 2.7. Since for $i\neq j$, $M_{i}+M_{j}=T$, we deduce that $R_{i}\neq R_{j}$. Hence $R_{1},R_{2},\ldots$ are distinct maximal subrings of $T$ and hence we are done.
$(2)$ thus assume that there exist infinitely many distinct maximal ideals $M_{1},M_{2},\ldots\in Max_{p,F}(T)$ such that $T/M_{i}=E_{i}$ is a finite field. Clearly by our assumption $\mathbb{Z}_{p}\subseteq E_{i}$, for each $i\geq 1$. Now if for infinitely many $i$, $\mathbb{Z}_{p}\subsetneq E_{i}$, then similar to proof of $(1)$ we obtain that $T$ has infinitely many maximal subrings (note that, $E_{i}$ is a finite field and $\mathbb{Z}_{p}\subsetneq E_{i}$, therefore $E_{i}$ has a maximal subring, see the comments preceding Theorem 2.7, and hence $T$ has a maximal subring $R_{i}$ which contains $M_{i}$). Thus without lose of the generality we may assume that for each $i\geq 1$, we have $\mathbb{Z}_{p}=E_{i}$. Let $P_{i}=M_{2i}$ and $Q_{i}=M_{2i-1}$, where $i\geq 1$. Then clearly for each $i\geq 1$, we have $T/(P_{i}\cap Q_{i})\cong\mathbb{Z}_{p}\times\mathbb{Z}_{p}$. Since $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ has a maximal subring (in fact, $\{(a,a)\ |\ a\in\mathbb{Z}_{p}\}$ is the only proper and therefore maximal subring of $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$, see the comments preceding Theorem 2.7), we conclude that $T$ has a maximal subring $R_{i}$ which contains $P_{i}\cap Q_{i}$. Clearly, whenever $i\neq j$, then $(P_{i}\cap Q_{i})+(P_{j}\cap Q_{j})=T$. Thus we deduce that for $i\neq j$ we have $R_{i}\neq R_{j}$ and therefore $T$ has infinitely many maximal subrings. ∎
It is not hard to see that by the proof of the previous result, in fact there exists a prime divisor $p$ of $n$ such that $|Max_{p,F}(T)|=|Max_{F}(T)|\leq|RgMax(T)|$ for some prime number $p$. Now the next result is in order.
Corollary 3.20.
Let $T$ be a ring. If $|Max_{F}(T)|>\aleph_{0}$, then $T$ has infinitely many maximal subrings. In fact, $|Max_{F}(T)|\leq|RgMax(T)|$.
Proof.
Since $|Max_{F}(T)|>\aleph_{0}$ and the number of prime numbers is countable, we conclude that there exists a prime number $p$ such that $|Max_{p,F}(T)|=|Max_{F}(T)|$. Now by a similar proof of Proposition 3.19, we can see that $|Max_{p,F}(T)|\leq|RgMax(T)|$ and hence we are done. ∎
Finally we conclude this section by the following result.
Proposition 3.21.
Let $T$ be a ring and $S_{1}\ncong S_{2}$ be simple $T$-modules. If any of the following conditions holds, then $T$ has infinitely many maximal subrings.
  1. (1)
    If $E_{1}=End_{T}(S_{1})$ is a field which is not an absolutely algebraic field.
  2. (2)
    If $E_{1}=End_{T}(S_{1})$ and $E_{2}=End_{T}(S_{2})$ are infinite isomorphic ring.
Proof.
First note that since $S_{i}$ is a simple left $T$-module, we infer that there exists maximal left ideals $M_{i}$, $i=1,2$, such that $S_{i}\cong T/M_{i}$, as left $T$-module.
$(1)$ Note that $E_{1}^{op}\cong\mathbb{I}(M_{1})/M_{1}$. Now we have two cases, either $M_{1}$ is an ideal of $T$ or not. If $M_{1}$ is an ideal of $T$, then $T/M_{1}\cong E_{1}^{op}$ as ring and therefore $T/M_{1}$ has infinitely many maximal subrings, by [10, Corollary 1.5]. Hence assume that $M_{1}$ is not an ideal of $T$. Since $E_{1}$ is not an absolutely algebraic field, we conclude that $E_{1}$ is an infinite ring. Therefore $\mathbb{I}(M_{1})/M_{1}$ is an infinite ring. Thus by Theorem 3.1, we conclude that $M_{1}]$ and therefore $Max_{l}(T)\setminus Max(T)$ are infinite and hence we are done by $(1)$ of Proposition 3.10.
$(2)$ We have two cases, if $M_{1}$ or $M_{2}$ is not an ideal of $T$. Then similar to last part of the proof of $(1)$ we infer that $T$ has infinitely many maximal subrings. Hence assume that $M_{1}$ and $M_{2}$ are ideals of $T$. In this case we have the following ring isomorphisms
$T/M_{1}=\mathbb{I}(M_{1})/M_{1}\cong E_{1}^{op}\cong E_{2}^{op}\cong\mathbb{I}(M_{2})/M_{2}=T/M_{2}.$
Thus $T/(M_{1}\cap M_{2})\cong E_{1}^{op}\times E_{1}^{op}$, as ring (note $M_{1}\neq M_{2}$, for $S_{1}\ncong S_{2}$). Now since $E_{1}$ and $E_{2}$ are infinite division rings, by Theorem 3.6, we deduce that $T/(M_{1}\cap M_{2})$ has infinitely many maximal subrings. Therefore $T$ has infinitely many maximal subrings. ∎
4. Rings integral over their centers
In this section we investigate about the number of maximal subrings in a ring $T$ which is integral over its center. First we need some observation and notation. Let $T$ be a ring, then it is clear that $Max_{l}(T)\cap Max(T)=Max_{r}(T)\cap Max(T)=Max_{r}(T)\cap Max_{l}(T)$, we denote this set by $Max_{lr}(T)$ and we define $J^{\prime}(T):=\bigcap_{M\in Max_{lr}(T)}M$. It is clear that $J^{\prime}(T)$ is an ideal of $T$ and $J(T)\subseteq J^{\prime}(T)$. In particular, if $T$ is either left or right quasi duo ring, then $J(T)=J^{\prime}(T)$. Moreover if $Max_{l}(T)\setminus Max(T)=\{M_{1},\ldots,M_{n}\}$ is finite (in particular, by $(5)$ of Proposition 3.10, where $T$ has only finitely many maximal subrings), then $J(T)=J^{\prime}(T)\cap M_{1}\cap\cdots\cap M_{n}$. Note that by Lemma 3.13, if $T$ is an infinite ring with $J^{\prime}(T)=0$, then either $T$ has infinitely many maximal subrings or $C(T)$ is integrally closed in $T$. Now the following is in order.
Proposition 4.1.
Let $T$ be a ring which is integral over its center. Then either $T$ has infinitely many maximal subrings or the following hold:
  1. (1)
    $T/J^{\prime}(T)$ is a reduce Hilbert commutative ring and $|T/J^{\prime}(T)|\leq 2^{2^{\aleph_{0}}}$. In particular, $N(T)\subseteq J^{\prime}(T)$.
  2. (2)
    $|Max_{lr}(T)|\leq 2^{\aleph_{0}}$. For each $M\in Max_{lr}(T)$, $T/M$ is an absolutely algebraic field (thus $|T/M|\leq\aleph_{0}$).
  3. (3)
    for each $M\in Max_{l}^{\prime}(T)\cup Max_{r}^{\prime}(T)$, $J^{\prime}(T)+M=T$.
Proof.
Assume that $T$ has only finitely many maximal subrings. Thus for each $M\in Max_{lr}(T)$, the division ring $T/M$ has only finitely many maximal subrings. Since $T$ is integral over its center, we conclude that for each $M\in Max_{lr}(T)$, the division ring $T/M$ is integral over its center. Thus by Proposition 3.1, we conclude that $T/M$ is a field. Clearly, $T/J^{\prime}(T)$ embeds in $\prod_{M\in Max_{lr}(T)}T/M$, thus $T/J^{\prime}(T)$ is a commutative reduced ring and therefore $N(T)\subseteq J^{\prime}(T)$. Since $T/J^{\prime}(T)$ is a commutative ring with only finitely many maximal subrings, by [10, Corollary 1.9], we deduce that $T/J^{\prime}(T)$ is a Hilbert ring. By [10, Corollary 1.5], for each $M\in Max_{lr}(T)$, the field $T/M$ is an absolutely algebraic field and therefore $|T/M|\leq\aleph_{0}$. Since $T/J^{\prime}(T)$ is a commutative ring, the first part of $(2)$ is an immediate consequence of [10, (1) of Proposition 2.8]. Therefore we deduce that
$|\frac{T}{J^{\prime}(T)}|\leq|\prod_{M\in Max_{lr}(T)}T/M|\leq 2^{2^{\aleph_{0}}}.$
Hence $(1)$ and $(2)$ hold. Finally for $(3)$, let $M$ be a maximal left ideal of $T$ which is not an ideal of $T$. Since $T/J^{\prime}(T)$ is a commutative ring, we conclude that $J^{\prime}(T)\nsubseteq M$ and therefore $M+J^{\prime}(T)=T$ by maximality of $M$. ∎
As we mentioned in the proof of the previous proposition, if $R$ is a commutative ring with $|Max(R)|>2^{\aleph_{0}}$, then $|RgMax(R)|\geq 2^{\aleph_{0}}$, see [10, (1) of Proposition 2.8]. For non-commutative rings we have the following result.
Proposition 4.2.
Let $T$ be a ring which is integral over its center. If one of the following conditions holds, then $T$ has infinitely many maximal subrings.
  1. (1)
    $|Max_{l}(T)|>2^{\aleph_{0}}$.
  2. (2)
    $|Max_{r}(T)|>2^{\aleph_{0}}$.
  3. (3)
    $|Max(T)|>2^{\aleph_{0}}$.
Proof.
We prove $(1)$ and $(3)$, the proof of $(2)$ is similar to $(1)$. For $(1)$, first note that if $Max_{l}(T)\setminus Max(T)$ is infinite, then by $(1)$ of Proposition 3.10, we deduce that $T$ has infinitely many maximal subrings. Hence assume that $Max_{l}(T)\setminus Max(T)$ is finite. Thus $|Max_{l}(T)\cap Max(T)|=|Max_{l}(T)|>2^{\aleph_{0}}$. Now since $Max_{lr}(T)=Max_{l}(T)\cap Max(T)$, we deduce that $T$ has infinitely many maximal subrings by $(2)$ of Proposition 4.1.
$(3)$ If $Max(T)\setminus Max_{l}(T)$ is infinite, then by $(5)$ of Proposition 3.10, we conclude that $T$ has infinitely many maximal subrings. Hence assume that $Max(T)\setminus Max_{l}(T)$ is finite and therefore $|Max(T)\cap Max_{l}(T)|=|Max(T)|>2^{\aleph_{0}}$. By a similar proof of $(1)$ we obtain that $T$ has infinitely many maximal subrings. ∎
As the following main result shows if we assume that a ring $T$ is an algebraic $K$-algebra (and therefore is integral over its center), then either $T$ has infinitely many maximal subrings or $T$ has similar algebraic properties like commutative rings with only finitely many maximal subrings, see [10].
Theorem 4.3.
Let $T$ be an algebraic $K$-algebra over an infinite field $K$. Then either $T$ has infinitely many maximal subrings or the following hold:
  1. (1)
    $T$ is a quasi duo ring. In particular, $J(T)=\bigcap_{M\in Max(T)}M$.
  2. (2)
    For each $M\in Max(T)$, the ring $T/M$ is an absolutely algebraic field which contains a copy of $K$. In particular, $K$ is an absolutely algebraic field and $|K|\leq|T/M|\leq\aleph_{0}$. Moreover, if $D$ is a division ring which is a subring of $T$, then $D$ is an absolutely algebraic field.
  3. (3)
    $|Max(T)|\leq 2^{\aleph_{0}}$ and $|T/J(T)|\leq 2^{2^{\aleph_{0}}}$.
  4. (4)
    $T/J(T)$ is a reduced Hilbert commutative ring. In fact, $T/J(T)$ is a von Neumann regular ring which is integral over its prime subring (i.e., $\mathbb{Z}_{p}$, where $p=Char(K)$ is a prime number). Moreover, $T$ is integral over $\mathbb{Z}_{p}$.
  5. (5)
    For each distinct maximal ideals $M$ and $N$ of $T$, the fields $T/M$ and $T/N$ are not isomorphic.
  6. (6)
    For each proper ideal $I$ of $T$, $N(T/I)=J(T/I)$.
  7. (7)
    Each strongly prime ideal $Q$ of $T$ is a maximal ideal.
  8. (8)
    The contraction of each maximal ideal of $T$ to any subring $R$ of $T$, remains a maximal ideal of $R$. In particular, $J(R)$ is nil.
Proof.
Assume that $T$ has only finitely many maximal subring. Thus by Theorem 3.16, we deduce that $(1)$ holds. The first part of $(2)$ is clear by $(2)$ of Proposition 4.1. For the second part of $(2)$, let $M$ be any maximal ideal of $T$. Clearly, $T/M$ contains a copy of $K$ (and similarly $D$), thus $K$ and $D$ are absolutely algebraic fields. $(3)$ and the first part of $(4)$ are clear by Proposition 4.1. For the second part of $(4)$, note that $T/J(T)$ is a commutative ring which is integral over the field $K$. Hence we deduce that $T/J(T)$ is a zero-dimensional (reduced) ring and therefore is a von Neumann ring. Since $T$ has only finitely many maximal subrings, we infer that $T/J(T)$ has only finitely many maximal subring. Thus by [10, Proposition 2.1], we obtain that $T/J(T)$ is integral over $\mathbb{Z}_{p}$ (note, $K$ is an absolutely algebraic field). Since $T$ is algebraic over $K$, by [21, Theorem 4.20], $J(T)$ is a nil ideal of $T$. Hence if $x\in T$, then there there exist $n\in\mathbb{N}$ and $a_{0},\ldots,a_{n-1}\in\mathbb{Z}_{p}$ such that $(x+J(T))^{n}+a_{n-1}(x+J(T))^{n-1}+\cdots+a_{1}(x+J(T))+a_{0}=0$ in the ring $T/J(T)$. Thus $x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0}\in J(T)$. Now since $J(T)$ is nil, we deduce that there exists a natural number $m$ such that $(x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0})^{m}=0$, i.e., $x$ is integral over $\mathbb{Z}_{p}$. For $(5)$, assume that $T/M\cong T/N$. Thus $T/(M\cap N)\cong T/M\times T/M$. Now note that $T/M$ contains a copy of $K$, and therefore $T/M$ is an infinite field. Hence by Proposition 3.6, $T/M\times T/M$ has infinitely many maximal subrings. Hence we conclude that $T$ has infinitely many maximal subrings, which is absurd. For $(6)$, since $T/I$ contains a copy of $K$ and remains algebraic over $K$ we may assume that $I=0$. Now note that $T/J(T)$ is a reduced ring. Therefore $N(T)\subseteq J(T)$. As we see in the proof of $(4)$, $J(T)$ is nil and thus $J(T)=N(T)$. For $(7)$, note that $T/Q$ has no nonzero nil ideal, hence by $(6)$ we conclude that $J(T/Q)=0$, i.e., $Q$ is an intersection of a family of maximal (left) ideal of $T$. Thus $J(T)\subseteq Q$ and therefore by $(4)$ we immediately deduce that $T/Q$ is a field, i.e., $Q$ is a maximal (left) ideal of $T$. Finally for $(8)$, let $M$ be a maximal ideal of $T$ and $R$ be a subring of $T$. Since $(R+M)/M$ is a subring of the absolutely algebraic field $T/M$, we deduce that $(R+M)/M$ is a field. Therefore $R/(R\cap M)$ is a field. Hence $R\cap M\in Max(R)$. This immediately implies that $J(R)\subseteq J(T)\cap R$ and therefore $J(R)$ is a nil ideal (note, $J(T)$ is nil by $(6)$). ∎
Lemma 4.4.
Let $T$ be a ring which is integral over its center. Assume that $T$ has only finitely many maximal subrings. if $P$ is a prime ideal of $T$ with $J(T/P)=0$, then either $P$ is a maximal ideal of $T$ or $T/P$ is an integral domain. In fact, either $T/P\cong\mathbb{M}_{n}(F)$ for some natural number $n>1$ and a finite field $F$ or $T/P$ is an integral domain.
Proof.
Since $T$ has only finitely many maximal subrings, then by $(1)$ of Proposition 3.10 we deduce that $Max_{l}^{\prime}(T)$ is finite. Hence by $J(T/P)=0$ we infer that there exist $M_{1},\ldots,M_{n}\in Max_{l}^{\prime}(T)$, $n\geq 0$ and $N=\bigcap_{i\in I}M_{i}$, where $M_{i}\in Max(R)\cap Max_{l}(T)$ such that $P=M_{1}\cap\cdots\cap M_{n}\cap N$. Thus by the comments preceding Proposition 3.10, either $P=M_{1}\cap\cdots\cap M_{n}$ or $P=N$, for $P$ is a prime ideal of $T$. If $P=M_{1}\cap\cdots\cap M_{n}$, then $T/P$ is a left $T$-submodule of the semisimple $T$-module $T/M_{1}\times\cdots\times T/M_{n}$. Therefore we conclude that $T/P$ is a simple Artinian ring. Hence $P$ is a maximal ideal of $T$. Thus assume that $P=N$. Since $T$ is integral over its center, we infer that for each $i\in I$, the division ring $T/M_{i}$ is integral over its center. Hence, if for some $i\in I$, $T/M_{i}$ is not a field, then by Theorem 3.1, we obtain that $T/M_{i}$ and therefore $T$ have infinitely many maximal subrings which is impossible. Thus for each $i\in I$, $T/M_{i}$ is a field. Clearly $T/P$ embeds in $\prod_{i\in I}T/M_{i}$ and therefore $T/P$ is a commutative (reduced prime) ring. Hence $T/P$ is an integral domain. For the final par, if $T/P$ is not an integral domain, then since $T/P$ is a simple Artinian ring, we conclude that $T/P\cong\mathbb{M}_{n}(D)$, where $D$ is a division ring and $n$ is a natural number. Now we have two cases: If $n=1$, then $T/P$ is a division ring which is integral over its center, hence by Theorem 3.1, we deduce that $T/P$ is a field, for $T/P$ has only finitely many maximal subrings. Thus $T/P$ is an integral domain which is absurd. Therefore $n>1$. We claim that $D$ is finite. For otherwise, by Theorem 3.5, we deduce that $T/P$ has infinitely many maximal subrings which is absurd. Thus $D$ is a finite field and hence we are done. ∎
Proposition 4.5.
Let $T$ be a ring which is integral over its center. Assume that $T$ has only finitely many maximal subrings. Then the following hold:
  1. (1)
    $Prm_{l}(T)=Max(T)=Prm_{r}(T)$.
  2. (2)
    If $C(R)$ is a zero-dimensional ring, then an ideal $P$ of $T$ is a maximal ideal of $T$ if and only if $P$ is a strongly prime ideal of $T$.
Proof.
$(1)$ Let $P$ be a left primitive ideal of $T$. Clearly, $J(T/P)=0$ and therefore by Lemma 4.4, we conclude that either $P$ is a maximal ideal of $T$ or $T/P$ is a commutative left primitive ring. Since a commutative primitive ring is a field, we infer that $T/P$ is a field and therefore $P$ is a maximal ideal of $T$. Similarly each right primitive ideal of $T$ is a maximal ideal.
$(2)$ It is clear that each maximal ideal $P$ of $T$ is a strongly prime ideal of $T$, for $T/P$ is a simple ring and therefore has no nonzero nil ideals. Conversely, let $P$ be a strongly prime ideal of $T$, i.e., $T/P$ has no nonzero nil ideal. It is not hard to see that $P\cap C(T)$ is a prime ideal of $C(T)$ and therefore by the assumption is a maximal ideal of $C(T)$. Therefore $C(T)/(P\cap C(T))\cong(C(T)+P)/P$ is a field. Since $T$ is integral over its center, we immediately conclude that $T/P$ is integral over the field $E:=(C(T)+P)/P$. Hence by [21, Theorem 4.20], $J(T/P)$ is a nil ideal and therefore $J(T/P)=0$, for $P$ is a strongly prime ideal of $T$. Hence by Lemma 4.4, either $P$ is a maximal ideal of $T$ or $T/P$ is an integral domain. In the second case note that since $T/P$ is integral over the field $E$, we immediately conclude that $P$ is maximal too. ∎
In the next main result in this section, we see that if $T$ is a ring which is integral over its center and has only finitely many maximal subrings, then $T/J(T)$ is embeddable in $R\times S$, where $R$ is a direct product of certain absolutely algebraic fields and $S$ is a finite semisimple ring.
Theorem 4.6.
Let $T$ be a ring which is integral over its center. Assume that $T$ has only finitely many maximal subrings. Then for each maximal ideal $M$ of $T$ either $T/M$ is an absolutely algebraic field with only finitely many maximal subrings or $T/M\cong\mathbb{M}_{n}(F)$ for some finite field $F$ and $n>1$. In particular, if $J(T)=0$, then $T$ embeds in
$\prod_{i\in I}E_{i}\times\prod_{j\in J}\mathbb{M}_{n_{j}}(F_{j})$
where each $E_{i}$ is an absolutely algebraic field with only finitely many maximal subrings, $n_{j}>1$ and each $F_{j}$ is a finite field. Moreover $J$ is finite, $|I|\leq 2^{\aleph_{0}}$ and if $E_{i}$ is infinite for some $i\in I$, then for each $i^{\prime}\in I\setminus\{i\}$, $E_{i}\ncong E_{i^{\prime}}$.
Proof.
Let $M$ be a maximal ideal of $T$, then by assumption $T/M$ has only finitely many maximal subrings and is integral over its center. Hence by Theorem 3.1, we infer that whenever $T/M$ is a division ring, then $T/M$ is a field with only finitely many maximal subrings and therefore $T/M$ is an absolutely algebraic field, by [10, Corollary 1.5]. Hence assume that $T/M$ is not a division ring. Now by Theorem 3.5, whenever $T/M$ is an infinite ring, then $T/M$ has infinitely many maximal subrings which is impossible. Thus $T/M$ is a finite ring. Therefore $T/M\cong\mathbb{M}_{n}(F)$, where $F$ is a finite ring and $n$ is a natural number. Since $T/M$ is not a division ring in this case, we deduce that $n>1$. Also note that since $T$ has only finitely many maximal subrings, then $T$ has only finitely many distinct maximal ideals $M$ for them the ring $T/M$ is not a field, by $(5)$ of Proposition 3.10, say $M_{1},\ldots,M_{k}$ (and assume that $T/M_{i}\cong\mathbb{M}_{n_{i}}(F_{i})$, where $F_{i}$ is a finite field and $n_{i}>1$). For the next part, by $(8)$ of Proposition 3.10, we have $0=J(T)=\bigcap_{M\in Max(T)}M$. Hence we conclude that $0=J(T)=M_{1}\cap\cdots\cap M_{k}\cap\bigcap_{i\in I}N_{i}$, where for each $i\in I$, $T/N_{i}$ is an absolutely algebraic field with only finitely many maximal subrings, by the first part of the theorem. Clearly, we may assume that for $i\neq j$ in $I$, $N_{i}\neq N_{j}$. Thus $T$ embeds in $T/M_{1}\times\cdots\times T/M_{k}\times\prod_{i\in I}T/N_{i}$. Clearly, by Proposition 4.2, we have $|Max(T)|\leq 2^{\aleph_{0}}$ and therefore $|I|\leq 2^{\aleph_{0}}$. Finally we show that if $i\neq j$ in $I$ and $T/N_{i}$ is infinite, then $T/N_{i}\ncong T/N_{j}$. Note that if $T/N_{i}\cong T/N_{j}$, then by Proposition 3.6, we deduce that $T/(N_{i}\cap N_{j})\cong T/N_{i}\times T/N_{i}$ has infinitely many maximal subrings which is impossible. Hence we are done (note that clearly $|J|=k$ is finite). ∎
In the next section we exactly determine when the ring $T=\prod_{i\in I}\mathbb{M}_{n_{i}}(E_{i})$, where each $E_{i}$ is a field and $n_{i}\in\mathbb{N}$, has only finitely many maximal subrings. In the following we obtain some results for Artinian/Noetherian rings. First we need the following remark.
Remark 4.7.
It is not hard to see that if $S$ is a commutative ring with nonzero characteristic, say $n$, $Max(S)=\{M_{1},\ldots,M_{k}\}$ and $J(S)$ is a nil ideal of $S$ (in particular, if $S$ is a commutative Artinian ring with nonzero characteristic $n$), then $S$ is integral over $\mathbb{Z}_{n}$ if and only if each $S/M_{i}$ is an absolutely algebraic field.
Theorem 4.8.
Let $T$ be a ring which is integral over its center $C:=C(T)$. If any of the following condition holds, then $T$ has infinitely many maximal subrings.
  1. (1)
    $C$ is an Artinian ring which is either uncountable or of zero characteristic and or is not integral over its prime subring.
  2. (2)
    $C$ is a Noetherian ring with $|C|>2^{\aleph_{0}}$.
Proof.
$(1)$ First note that by [8, Propositions 1.4 and 2.4] and the previous remark, $C$ has a maximal ideal $M$ such that $C/M$ is a field which is either uncountable or of zero characteristic and or is not absolutely algebraic field, respectively. Hence in any cases, $K:=C/M$ is not an absolutely algebraic field. By [11, Proposition 1.2], $T$ has a prime ideal $Q$ such that $Q\cap C=M$. Clearly, $K=C/M\subseteq C(T/Q)$ and $T/Q$ is integral over $K$ and therefore $T/Q$ is integral over its center. Thus by $(2)$ of Theorem 4.3, we conclude that $T/Q$ has infinitely many maximal subrings. Hence $T$ has infinitely many maximal subrings.
$(2)$ First assume that $C$ is an integral domain. By the proof of [9, Corollary 2.7], for each maximal ideal $M$ of $C$, there exists a natural number $n$, such that $C/M^{n}$ is an uncountable ring. Since $C/M^{n}$ is a zero dimensional noetherian local ring with only maximal ideal $M/M^{n}$, by the first part of the proof of $(1)$, we deduce that $C/M$ is an uncountable field and similar to the proof of $(1)$, we conclude that $T$ has infinitely many maximal subrings. If $C$ is not an integral domain, then by the proof of [9, Theorem 2.9], $C$ has a minimal prime ideal $P$ such that $|C/P|=|C|$. By [11, Proposition 1.2], $T$ has a prime ideal $Q$ such that $Q\cap C=P$. Clearly $T/Q$ is integral over $C/P\subseteq C(T/Q)$ (and therefore $T/Q$ is integral over its center). Similar to the first part of the proof of $(2)$, for each maximal ideal $M/P$ of $C/P$, the field $(C/P)/(M/P)$ is uncountable. Since $T/Q$ is integral over $C/P$, by [11, Proposition 1.2], $T/Q$ has a prime ideal $N/Q$ such that $(N/Q)\cap(C/P)=M/P$ and clearly $(T/Q)/(N/Q)$ is integral over $K:=(C/P)/(M/P)\subseteq C((T/Q)/(N/Q))$ (and therefore $(T/Q)/(N/Q)$ is integral over its center). Now since $(T/Q)/(N/Q)$ contains an uncountable field $K=(C/P)/(M/P)$, we immediately conclude that $(T/Q)/(N/Q)$ has infinitely many maximal subrings by $(2)$ of Theorem 4.3. Thus $T$ has infinitely many maximal subrings too. ∎
Theorem 4.9.
Let $T$ be a left Artinian algebraic $K$-algebra over an infinite field $K$. Then either $T$ has infinitely many maximal subrings or $T$ is countable and is integral over its prime subring.
Proof.
Assume that $T$ has only finitely many maximal subrings, then by $(4)$ of Theorem 4.3, we infer that $T/J(T)$ is a von Neumann regular commutative ring and $T$ is integral over $\mathbb{Z}_{p}$, where $p=Char(K)$ is a prime number. Now note that there exists a natural number $m$ such that $J(T)^{m}=0$, for $T$ is a left Artinian ring. Since $T/J(T)$ is an Artinian commutative ring with only finitely many maximal subrings, by $(2)$ of [10, Theorem 2.5], we deduce that $T/J(T)$ is a countable ring. Hence by Lemma 2.8, $T$ is a countable ring. ∎
Corollary 4.10.
Let $T$ be a left Noetherian algebraic $K$-algebra over an infinite field $K$. Then either $T$ has infinitely many maximal subrings or $T$ is a countable left Artinian ring which is integral over its prime subring.
Proof.
First note that since $T$ is an algebraic $K$-algebra over a field $K$, we infer that $J(T)$ is a nil ideal of $T$, by [21, Theorem 4.20]. Therefore by [21, Theorem 10.30], $J(T)$ is a nilpotent ideal of $T$, for $T$ is a left Noetherian ring. Hence we conclude that $T$ is a left Artinian ring by [21, Theorem 4.15]. Thus by Theorem 4.9, we deduce the assertion. ∎
Finally we conclude this section by the following result for infinite direct product.
Theorem 4.11.
Let $\{T_{i}\}_{i\in I}$ be an infinite family of rings and $T=\prod_{i\in I}T_{i}$. If each of the following conditions holds, then $T$ has infinitely many maximal subrings:
  1. (1)
    If each $T_{i}$ is integral over its center.
  2. (2)
    If $T$ is integral over its center.
Proof.
$(1)$ Let $J$ be the set of all $i\in I$ such that $T_{i}$ is not a quasi duo ring. Now we have two cases. First assume that $J$ is infinite. For each $j\in J$, let $M_{i}$ be a maximal one-sided ideal of $T_{j}$ which is not an ideal of $T_{j}$, and $N_{j}=\prod_{i\in I}A_{i}$, where $A_{i}=T_{i}$ for $i\neq j$ and $A_{j}=M_{j}$. Clearly, each $N_{j}$ is a maximal one-sided ideal of $T$ which is not an ideal of $T$. Hence $T$ has infinitely many maximal one-sided ideal which are not ideals of $T$. Thus $T$ has infinitely many maximal subrings by $(1)$ of Proposition 3.10. Hence assume that $J$ is finite. Without lose of the generality we may assume that $J=\emptyset$, i.e., each $T_{i}$ is a quasi duo ring and each $T_{i}$ has only finitely many maximal subrings. Therefore by Proposition 4.1, for each $i\in I$, the ring $T_{i}/J(T_{i})$ is a commutative ring. Take $J:=\prod_{i\in I}J(T_{i})$, then clearly $T/J\cong\prod_{i\in I}T/J(T_{i})$. Thus $T/J$ is a product of an infinite family of commutative rings, and therefore by [9, Remark 3.18] (or see [10, Theorem 3.9]), $T/J$ has infinitely many maximal subrings and so does $T$. Hence $(1)$ holds. For $(2)$, it suffices to show that each $T_{i}$ is integral over its center and use $(1)$. Clearly, $C_{T}(T)=\prod_{i\in I}C_{T_{i}}(T_{i})$. Let $k\in I$ and $t\in T_{k}$, and put $x=(x_{i})_{i\in I}$ where $x_{i}=0$ for $i\neq k$ and $x_{k}=t$. By our assumption, $x$ is integral over the center of $T$. Hence there exist $n\in\mathbb{N}$ and $c_{0},c_{1},\ldots,c_{n-1}$ in $C_{T}(T)$ such that $x^{n}+c_{n-1}x^{n-1}+\cdots+c_{1}x+c_{0}=0$. Since $C_{T}(T)=\prod_{i\in I}C_{T_{i}}(T_{i})$, we have $c_{j}=(c_{ji})_{i\in I}$, for $0\leq j\leq n-1$, where $c_{ji}\in C_{T_{i}}(T_{i})$. Therefore by calculating the $k$-th component from $x^{n}+c_{n-1}x^{n-1}+\cdots+c_{1}x+c_{0}=0$, we immediately conclude that $t^{n}+c_{n-1k}t^{n-1}+\cdots+c_{1k}t+c_{0k}=0$. Hence $t$ is integral over $C_{T_{i}}(T_{i})$ and we are done by $(1)$. ∎
5. Direct product and matrix rings
In this section we study the number of maximal subrings in matrix rings and (certain) direct products of rings. In [10, Theorem 3.9], the authors exactly determined when a direct product of a family of commutative rings has only finitely many maximal subrings. In particular, if $T$ is an infinite direct product of a family of commutative rings $\{R_{i}\}_{i\in I}$, then $|RgMax(T)|\geq 2^{|I|}$. As we see in the final result of the previous section, if $T$ is a direct product of an infinite family $\{T_{i}\}_{i\in I}$, and either each $T_{i}$ is integral over its center, or $T$ is integral over its center, then $T$ has infinitely many maximal subrings. This immediately implies that whenever for each $i$, $T_{i}=\mathbb{M}_{n_{i}}(E_{i})$, where $E_{i}$ is a field and $n_{i}\in\mathbb{N}$, then $T$ has infinitely many maximal subrings (note, for each $i$, $\mathbb{M}_{n_{i}}(E_{i})$ is finite dimensional over its center and therefore is integral over its center). In the next proposition we exactly determine when $T$ has only finitely many maximal subrings.
Proposition 5.1.
Let $T=\prod_{i\in I}\mathbb{M}_{n_{i}}(E_{i})$, where each $E_{i}$ is a field and $n_{i}\in\mathbb{N}$. Then $T$ has only finitely many maximal subrings if and only if the following hold:
  1. (1)
    II is finite.
  2. (2)
    For each $i\in I$, if $n_{i}>1$, then $E_{i}$ is finite.
  3. (3)
    For each $i,j\in I$, where $n_{i}=n_{j}=1$, and $E_{i}$ is infinite, then $E_{i}\ncong E_{j}$.
  4. (4)
    For each $i\in I$ such that $n_{i}=1$, $E_{i}$ has only finitely many maximal subrings.
Therefore, if $T$ has only finitely many maximal subrings, then $T=E_{1}\times\cdots\times E_{r}\times\mathbb{M}_{n_{1}}(E_{r+1})\times\cdots\mathbb{M}_{n_{s}}(E_{r+s})$, where $n_{i}>1$ and $r,s\geq 0$. In particular, $T$ is a semisimple ring. Moreover, if $A=E_{1}\times\cdots\times E_{r}$ and $B=\mathbb{M}_{n_{1}}(E_{r+1})\times\cdots\mathbb{M}_{n_{s}}(E_{r+s})$, then each maximal subring of $T$ is of the form $A\times B_{1}$, where $B_{1}$ is a maximal subring of $B$, or $A_{1}\times B$, where $A_{1}$ is a maximal subring of $A$.
Proof.
First assume that $T$ has only finitely many maximal subrings. This immediately implies that each $\mathbb{M}_{n_{i}}(E_{i})$ has finitely many maximal subrings. Hence if $n_{i}=1$, $E_{i}$ has only finitely many maximal subrings and if $n_{i}>1$, then by Theorem 3.5, $\mathbb{M}_{n_{i}}(E_{i})$ is finite and therefore $E_{i}$ is finite too. Therefore $(2)$ and $(4)$ hold. Although the proof of $(1)$ is obvious by the comment preceding the proposition, but we give another proof too. To see this, let $J:=\{i\in I\ |\ n_{i}=1\}$, $J^{\prime}=I\setminus J$, $A=\prod_{i\in J}E_{i}$ and $B=\prod_{i\in J^{\prime}}\mathbb{M}_{n_{i}}(E_{i})$. Then clearly $T=A\times B$. Hence by our assumption $A$ and $B$ has only finitely many maximal subrings. Thus by [10, Theorem 3.9], we immediately conclude that $J$ is finite. It is clear that by $(5)$ of Proposition 3.10, $J^{\prime}$ is finite and therefore $I$ is finite. Finally note that $(3)$ is an immediate by Proposition 3.6. Conversely, assume that (1)(4)(1)-(4) hold and let $A$ and $B$ are defined as above. We prove that $T$ has only finitely many maximal subrings. To see this first note that $A$ and $B$ are homomorpically non-isomorphic rings, hence by Lemma 2.9, each maximal subring of $T$ is of the form $A\times B_{1}$, where $B_{1}$ is a maximal subring of $B$, or $A_{1}\times B$, where $A_{1}$ is a maximal subring of $A$. This immediately shows that $T$ has only finitely many maximal subrings, for $B$ is finite and by [10, Corollary 3.12 and $(1)$ of Proposition 3.15], $A$ has only finitely many maximal subrings. Thus we are done. ∎
Proposition 5.2.
Let $T=\prod_{i\in I}\mathbb{M}_{n_{i}}(D_{i})$, where each $D_{i}$ is a division ring, $1<n_{i}\in\mathbb{N}$. Then $T$ has only finitely many maximal subrings if and only if $T$ is a finite ring.
Proof.
Assume that $T$ has only finitely many maximal subrings. Then similar to the proof of the previous theorem we deduce that $I$ is finite and for each $i\in I$, $\mathbb{M}_{n_{i}}(D_{i})$ has only finitely many maximal subrings. Hence by Theorem 3.5, $D_{i}$ is finite and thus $T$ is finite. The converse is evident. ∎
Corollary 5.3.
Let $T=\prod_{i\in I}\mathbb{M}_{n_{i}}(D_{i})$, where each $D_{i}$ is a division ring which is algebraic over its center (in particular, if each $D_{i}$ is finite dimensional over its center). Then $T$ has only finitely many maximal subrings if and only if the following hold:
  1. (1)
    II is finite.
  2. (2)
    for each $i\in I$ whenever $n_{i}>1$, then $D_{i}$ is a finite field.
  3. (3)
    for each $i\in I$ whenever $n_{i}=1$, then $D_{i}$ is a field with only finitely many maximal subrings.
  4. (4)
    for each $i\neq j$ in $I$, whenever $n_{i}=n_{j}=1$ and $D_{i}$ is infinite, then $D_{i}\ncong D_{j}$.
Proof.
First assume that $T$ has only finitely many maximal subrings. Thus for each $i\in I$, we conclude that $\mathbb{M}_{n_{i}}(D_{i})$ has only finitely many maximal subrings. Now if $n_{i}>1$, then by Theorem 3.5 we deduce that $D_{i}$ is a finite field. Hence assume that $n_{i}=1$, we claim that $D_{i}$ is a field. For otherwise by Proposition 3.1, $D_{i}$ has infinitely many maximal subrings, which is impossible. Thus for each $i\in I$, we conclude that $D_{i}$ is a field and hence we are done by Proposition 5.1. ∎
In [10, Corollary 3.10], the authors proved that if $R$ is a commutative ring, then $R\times R$ has only finitely many maximal subrings if and only if $R$ is a semilocal ring (i.e., has only finitely many maximal ideals) with only finitely many maximal subrings (in particular, $R$ is integral over $\mathbb{Z}_{n}$ for some $n>1$) and for each maximal ideal $M$ of $R$, the field $R/M$ is finite. For non-commutative rings we have the following result.
Proposition 5.4.
Let $R$ be a ring and $T=R\times R$. If $T$ has only finitely many maximal subrings, then the following hold:
  1. (1)
    $Max(R)$, $Max_{l}(R)$ and $Max_{r}(R)$ are finite sets. In particular, $Max(T)$, $Max_{l}(T)$ and $Max_{r}(T)$ are finite sets.
  2. (2)
    For each $M\in Max(R)$, $R/M$ is finite. In particular, for each maximal ideal $Q$ of $T$, the ring $T/Q$ is finite.
  3. (3)
    $R/J(R)$ is a finite semisimple ring. In particular, $T/J(T)$ is a finite semisimple ring.
  4. (4)
    For each maximal one-sided ideal $M$ of $R$, $R/M$ is finite (i.e., each simple left/right $R$-module is finite). In particular, for each maximal one-sided ideal $N$ of $T$, $T/M$ is finite (i.e., each simple left/right $T$-module is finite).
Proof.
First note that for each maximal ideal $M$ of $R$, $T/(M\times M)\cong R/M\times R/M$ has only finitely many maximal subrings. Thus by Proposition 3.6, $R/M$ is finite. Hence the first part of $(2)$ holds. Also note that for each $M\in Max(R)$, $T$ has a maximal subring $R_{M}$ which contains $M\times M$ (by the comments preceding Theorem 2.7) and clearly whenever $M\neq N\in Max(R)$, then $R_{M}\neq R_{N}$, for $(M\times M)+(N\times N)=T$. Hence we conclude that $Max(R)$ is finite. Since $T$ has only finitely many maximal subrings, we infer that $R$ has only finitely many maximal subring. This immediately implies that $Max_{l}(R)\setminus Max(R)$ is finite, by $(1)$ of Proposition 3.10. Therefore $Max_{l}(R)$ is finite, for $Max(R)$ is finite by the first part of the proof. Similarly $Max_{r}(R)$ is finite. For the second part of $(1)$, note that maximal (left/right) ideals of $T$ are of the forms $M\times R$ or $R\times M$, where $M$ is a maximal (left/right) ideal of $R$. Hence the conclusions are immediate by the first part of $(1)$. By the latter comment and the first part of $(2)$, we deduce the second part of $(2)$. For $(3)$, note that since $R$ has only finitely many maximal left ideals, we conclude that $R/J(R)$ is a semisimple ring. Hence by $(2)$, we immediately conclude that $R/J(R)$ is finite. The second part of $(3)$ is evident, for $J(T)=J(R)\times J(R)$. $(4)$ is an immediate consequences of $(3)$, for $R/J(R)$ (resp. $T/J(T)$) is a finite ring. ∎
By the previous result we immediately conclude that if $R$ is a ring and $T=R\times R$ has only finitely many maximal subrings, then there exists a finite number of non-isomorphic finite simple left $R$-module $S_{1},\ldots,S_{n}$ such that each simple left $R$-module is isomorphic to one of the $S_{i}$.
Proposition 5.5.
Let $R$ be a ring and $T=\mathbb{M}_{n}(R)$ where $n>1$. If $T$ has only finitely many maximal subring, then the following hold:
  1. (1)
    $R$ has only finitely many maximal subrings.
  2. (2)
    for each maximal ideal $M$ of $R$, the ring $R/M$ is a finite ring. In particular, for each maximal ideal $Q$ of $T$, the ring $T/Q$ is finite.
  3. (3)
    $Max(R)$ and therefore $Max(T)$ are finite. Moreover, $Max_{l}(R)$ and $Max_{r}(R)$ are finite.
  4. (4)
    $R/J(R)$ and therefore $T/J(T)$ are finite semisimple ring. In particular, $Max_{l}(T)$ and $Max_{r}(T)$ are finite.
  5. (5)
    For each maximal one-sided ideal $M$ of $R$, $R/M$ is finite (i.e., each simple left/right $R$-module is finite). In particular, for each maximal one-sided ideal $N$ of $T$, $T/M$ is finite (i.e., each simple left/right $T$-module is finite).
Proof.
Clearly for each maximal subring $S$ of $R$, $\mathbb{M}_{n}(S)$ is a maximal subring of $T$, thus we deduce that $R$ has only finitely many maximal subrings. Hence $(1)$ holds. For $(2)$, note that for each maximal ideal $M$ of $R$, $A_{M}=\mathbb{M}_{n}(M)$ is a maximal ideal of $T$ and $T/A_{M}\cong\mathbb{M}_{n}(R/M)$. By the assumption that $T$ has only finitely many maximal subrings, we deduce that $T/A_{M}$ has only finitely many maximal subrings too. Hence by Theorem 3.5, we deduce that $R/M$ is finite (note, $n>1$ and therefore $T/A_{M}$ is not a division ring). The second part of $(2)$ is obvious by the first part of $(2)$, for maximal ideals of $T$ is of the form $A_{M}$, for some maximal ideal $M$ of $R$. For $(3)$, by the notation of $(2)$, since $n>1$, by $(4)$ of Theorem 2.7, we conclude that $\mathbb{M}_{n}(R/M)$ has a maximal subring. Hence, $T$ has a maximal subring which contains $A_{M}$, say $T_{M}$. In particular, if $M\neq N\in Max(R)$, then $T_{M}\neq T_{N}$, for $A_{M}+A_{N}=T$. Thus we deduce that $Max(R)$ is finite. Therefore $Max(T)$ is finite too. For the second part of $(3)$, note that by $(1)$, $R$ has only finitely many maximal subrings, hence by $(1)$ of Proposition 3.10, we conclude that $Max_{l}(R)\setminus Max(R)$ and $Max_{r}(R)\setminus Max(R)$ are finite. Therefore $Max_{l}(R)$ and $Max_{r}(R)$ are finite, for $Max(R)$ is finite by the first part of $(3)$. For $(4)$, note that since $Max_{l}(R)$ is finite we deduce that $R/J(R)$ is a semisimple ring. Since for each maximal ideal $M$ of $R$, $R/M$ is a finite ring, we immediately conclude that $R/J(R)$ is a finite ring. Since $J(T)=\mathbb{M}_{n}(J(R))$, we deduce that $T/J(T)$ is finite semisimple too. Therefore $Max_{l}(T)$ and $Max_{r}(T)$ are finite. $(5)$ is obvious by $(4)$. ∎
Now we have the following corollaries.
Corollary 5.6.
Let $R$ be an infinite ring with $J(R)$ is nilpotent (in particular, if $R$ is an infinite left/right Artinian ring). Then $\mathbb{M}_{n}(R)$, where $n>1$, and $R\times R$ have infinitely many maximal subrings.
Proof.
Assume that $\mathbb{M}_{n}(R)$ or $R\times R$ has only finitely many maximal subrings. Then by Propositions 5.4 or 5.5, we deduce that $R/J(R)$ is a finite ring. Therefore by Lemma 2.8, we conclude that $R$ is a finite ring, which is impossible. Hence we are done. ∎
Corollary 5.7.
Let $R$ be a zero-dimensional ring with $Char(R)=0$. Then $\mathbb{M}_{n}(R)$, where $n>1$, and $R\times R$ has infinitely many maximal subrings.
Proof.
Assume that $\mathbb{M}_{n}(R)$ or $R\times R$ has only finitely many maximal subrings. Then by Propositions 5.4 or 5.5, we deduce that for each maximal ideal $M$ of $R$, $R/M$ is a finite ring and therefore $Char(R/M)\neq 0$. Now note that there exists a prime ideal $P$ of $R$ such that $P\cap\mathbb{Z}=0$. Since $R$ is a zero-dimensional ring we deduce that $P$ is a maximal ideal of $R$ and therefore $R/P$ is finite. But from $\mathbb{Z}\cap P=0$, we deduce that $R/P$ contains a copy of $\mathbb{Z}$, which is absurd. ∎
Corollary 5.8.
Let $R$ be a $K$-algebra over an infinite field $K$. Then $\mathbb{M}_{n}(R)$, where $n>1$, and $R\times R$ has infinitely many maximal subrings.
Proof.
Assume that $\mathbb{M}_{n}(R)$ or $R\times R$ has only finitely many maximal subrings. Then by Propositions 5.4 or 5.5, we deduce that $R/J(R)$ is a finite ring. But clearly $R/J(R)$ contains a copy of $K$, which is impossible for $K$ is infinite. Hence we are done. ∎
Proposition 5.9.
Let $T$ be a semilocal ring (in particular, if $T$ is a left/right Artinian ring) with only finitely many maximal subrings. Then $T/J(T)\cong D_{1}\times\cdots\times D_{n}\times\mathbb{M}_{n_{1}}(E_{1})\times\cdots\times\mathbb{M}_{n_{k}}(E_{k})$, where each $D_{i}$ is a division ring, $1\leq i\leq n$, $E_{j}$ is a finite field, $1\leq j\leq k$ and $n_{j}>1$, for some $n,k\geq 0$. Moreover, whenever $D_{i}$ is infinite, then $D_{i}\ncong D_{j}$ for $j\neq i$.
Proof.
Since $T/J(T)$ is a semisimple ring, we deduce that $T/J(T)\cong\prod_{i\in I}\mathbb{M}_{n_{i}}(D_{i})$, where $I$ is finite, $D_{i}$ are division rings and $n_{i}\in\mathbb{N}$. Clearly $T$ has only finitely many maximal ideals. Also note that since $T$ has only finitely many maximal subrings, then by $(1)$ of Proposition 3.10, we conclude that $Max_{l}(T)\setminus Max(T)$ is finite. Therefore $Max_{l}(T)$ is finite. Similarly $Max_{r}(R)$ is finite. Hence by Corollary 2.4, we deduce that whenever $i\in I$ and $n_{i}>1$, then $D_{i}$ is finite, i.e., $D_{i}$ is a finite field. Hence we infer that $T/J(T)\cong D_{1}\times\cdots\times D_{n}\times\mathbb{M}_{n_{1}}(E_{1})\times\cdots\times\mathbb{M}_{n_{k}}(E_{k})$, where each $D_{i}$ is a division ring, $1\leq i\leq n$ and $E_{j}$ is a finite field, $1\leq j\leq k$ and $n_{j}>1$ for some $n,k\geq 0$. Finally if for some $1\leq i\neq j\leq n$, $D_{i}\cong D_{j}$, then $T$ has an ideal $I$ such that $T/I\cong D_{i}\times D_{j}\cong D_{i}\times D_{i}$. Since $T$ has only finitely many maximal subrings, we deduce that $T/I$ has only finitely many maximal subrings. Hence by Proposition 3.6, $D_{i}$ is finite. Hence we are done. ∎
We conclude this paper by the following result.
Corollary 5.10.
Let $T$ be a semilocal ring. If $T$ has only finitely many maximal subrings, then $T$ has only finitely many maximal left/right ideals.
Acknowledgement
The first author is grateful to the Research Council of Shahid Chamran University of Ahvaz (Ahvaz-Iran) for financial support (Grant Number: SCU.MM1404.721)

### v2

@@
-Non-commutative rings with infinitely many maximal subrings
+Noncommutative rings with infinitely many maximal subrings
 Abstract.
-We study rings with infinitely (only finitely) many maximal subrings.
-We prove that if $M$ is a maximal left/right ideal of a ring $T$ which is not an ideal of $T$, and $R$ is the idealizer of $M$, then $T$ has at least $|R/M|+1$ maximal left/right ideals which are not an ideal of $T$; in particular $T$ has at least $|R/M|+1$ distinct maximal subrings.
-Moreover, if $T$ is a $K$-algebra over an infinite field $K$, then either $T$ has infinitely many maximal subrings or $T$ is a quasi duo ring with certain algebraic properties similar to commutative rings.
+We investigate rings with infinitely many or only finitely many maximal subrings.
+We prove that if $M$ is a maximal left/right ideal of a ring $T$ that is not an ideal of $T$, and $R$ is the idealizer of $M$, then $T$ has at least $|R/M|+1$ maximal left/right ideals that are not two-ideals of $T$; in particular $T$ has at least $|R/M|+1$ distinct maximal subrings.
+Moreover, if $T$ is a $K$-algebra over an infinite field $K$, then either $T$ has infinitely many maximal subrings or $T$ is a quasi-duo ring with certain algebraic properties similar to those commutative rings that have not infinitely many maximal subrings.
 We prove that for a simple ring $R$, the ring $R\times R$ has only finitely many maximal subrings if and only if $R$ is finite.
-Also we study rings which are integral over their centers and have only finitely many maximal subrings.
+We also study rings which are integral over their centers and have only finitely many maximal subrings.
 We prove that if $T$ is integral over its center and $T$ has more than $2^{\aleph_{0}}$ maximal (left/right) ideals, then $T$ has infinitely many maximal subrings.
-In particular, we see that if a $J$-semisimple ring $T$ is integral over its center and has only finitely many maximal subrings, then $T$ embeds in $S\times\prod_{i\in I}E_{i}$, where each $E_{i}$ is an absolutely algebraic field and $S$ is a finite semisimple ring.
-We see that if $T$ is a left Noetherian algebraic $K$-algebra over an infinite field $K$ and $T$ has only finitely many maximal subrings, then $T$ is a countable left Artinian ring which is integral over $\mathbb{Z}_{p}$, where $p=Char(K)$.
-We exactly determine when $T=\prod_{i\in I}\mathbb{M}_{n_{i}}(E_{i})$, where each $E_{i}$ is a field and $n_{i}\in\mathbb{N}$, has only finitely many maximal subrings.
-We see that if $R$ is an infinite Artinian ring, then $\mathbb{M}_{n}(R)$, $n>1$, and $R\times R$ have infinitely many maximal subrings.
+In particular, we show that if a $J$-semisimple ring $T$ is integral over its center and has only finitely many maximal subrings, then $T$ embeds into $S\times\prod_{i\in I}E_{i}$, where each $E_{i}$ is an absolutely algebraic field and $S$ is a finite semisimple ring.
+Furthermore, if $T$ is a left Noetherian algebraic $K$-algebra over an infinite field $K$ and $T$ has only finitely many maximal subrings, then $T$ is a countable left Artinian ring which is integral over $\mathbb{Z}_{p}$, where $p=\operatorname{Char}(K)$.
+We precisely determine when $T=\prod_{i\in I}\mathbb{M}_{n_{i}}(E_{i})$, where each $E_{i}$ is a field and $n_{i}\in\mathbb{N}$, has only finitely many maximal subrings.
+Finally, we see that if $R$ is an infinite Artinian ring, then $\mathbb{M}_{n}(R)$, for $n>1$, and $R\times R$ have infinitely many maximal subrings.
 Key words and phrases:
-maximal subring, non-commutative ring, idealizer, left maximal ideal, integrality, direct product
+maximal subring, noncommutative ring, idealizer, maximal left ideal, integrality, direct product
 2010 Mathematics Subject Classification:
@@
 Shahid Chamran University of Ahvaz, Ahvaz-Iran
-a-azarang@scu.ac.ir
+azarang@scu.ac.ir
 ORCID ID: orcid.org/0000-0001-9598-2411
 1. Introduction
-Let $T$ be an associative ring with identity $1\neq 0$ and $RgMax(T)$ be the set of all maximal subrings (which contain $1_{T}$) of $T$, then we study when $RgMax(T)$ is finite/infinite.
-In [2], the author characterized fields and affine rings which have only finitely many maximal subrings; and in [10], the authors studied commutative rings with infinitely/only finitely many maximal subrings.
-We present some facts from them (which are needed in this paper) as follows:
+Let $T$ be an associative ring with identity $1\neq 0$ and let $\operatorname{RgMax}(T)$ denote the set of all maximal subrings (which contain $1_{T}$) of $T$.
+In this paper we study conditions under which $\operatorname{RgMax}(T)$ is finite or infinite.
+In [2], the author characterized fields and affine rings that have only finitely many maximal subrings; in [10], the authors studied commutative rings with either infinitely many or only finitely many maximal subrings.
+We recall some relevant facts from those works (which will be needed later) as follows:
   1. (1)
-    A field $E$ has only finitely many maximal subrings if and only if $E$ has a subfield $F$ without maximal subrings with $[E:F]$ is finite; if and only if every descending chain $\cdots\subset R_{2}\subset R_{1}\subset R_{0}=E$, where each $R_{i}$ is a maximal subring of $R_{i-1}$, $i\geq 1$, is finite; moreover $F$ is unique, all such chains have a same length, say $m$, and $R_{m}=F$, see [2, Theorems 3.2 and 3.6].
+    A field $E$ has only finitely many maximal subrings if and only if it contains a subfield $F$ without maximal subrings with $[E:F]$ is finite; equivalently, every descending chain $\cdots\subset R_{2}\subset R_{1}\subset R_{0}=E$, where each $R_{i}$ is a maximal subring of $R_{i-1}$, $i\geq 1$, is finite.
+Moreover, such a subfield $F$ is unique, all such chains have the same length, say $m$, and $R_{m}=F$; see [2, Theorems 3.2 and 3.6].
   2. (2)
     Let $R=F[\alpha_{1},\ldots,\alpha_{n}]$ be an affine integral domain over a field $F$.
-Then $R$ has only finitely many maximal subrings if and only if $F$ has only finitely many maximal subrings and each $\alpha_{i}$ is algebraic over $F$, see [2, Theorem 4.1].
+Then $R$ has only finitely many maximal subrings if and only if $F$ has only finitely many maximal subrings and each $\alpha_{i}$ is algebraic over $F$; see [2, Theorem 4.1].
   3. (3)
-    For each ring $R$, either $R$ is a Hilbert ring or $RgMax(R)$ is infinite, see [10, Corollary 1.9].
+    For any ring $R$, either $R$ is a Hilbert ring or $\operatorname{RgMax}(R)$ is infinite, see [10, Corollary 1.9].
   4. (4)
-    If $R$ is an Artinian ring which is either uncountable or of zero characteristic and or is not integral over its prime subring, then $RgMax(R)$ is infinite, see [10, Theorem2.5].
+    If $R$ is an Artinian ring that is either uncountable, or of zero characteristic, or is not integral over its prime subring, then $\operatorname{RgMax}(R)$ is infinite; see [10, Theorem2.5].
   5. (5)
-    If $R$ is a Noetherian ring with $|R|>2^{\aleph_{0}}$, then $|RgMax(R)|\geq 2^{\aleph_{0}}$, see [10, Theorem 2.7].
+    If $R$ is a Noetherian ring with $|R|>2^{\aleph_{0}}$, then $|\operatorname{RgMax}(R)|\geq 2^{\aleph_{0}}$; see [10, Theorem 2.7].
   6. (6)
-    If $\{R_{i}\ |\ i\in I\}$ is a family of rings and $R=\prod_{i\in I}R_{i}$, then in [10, Theorem 3.9], the authors exactly determined when $R$ has only finitely many maximal subrings.
-Moreover, if $I$ is infinite then $|RgMax(R)|\geq 2^{|I|}$, see [9, Remark 3.18].
+    Let $\{R_{i}\ |\ i\in I\}$ be a family of rings and $R=\prod_{i\in I}R_{i}$, in [10, Theorem 3.9], the authors precisely determined when $R$ has only finitely many maximal subrings.
+Moreover, if $I$ is infinite, then $|\operatorname{RgMax}(R)|\geq 2^{|I|}$; see [9, Remark 3.18].
   7. (7)
-    If $K$ is a field, then $K\times K$ has finitely many maximal subrings if and only if $K$ is a finite field, see [10, Corollary 3.5].
+    For a field $K$, the ring $K\times K$ has only finitely many maximal subrings if and only if $K$ is a finite field; see [10, Corollary 3.5].
   8. (8)
-    If $R$ is a semilocal ring, then either $R$ has infinitely many maximal subrings or $R$ is integral over $\mathbb{Z}_{n}$, for some $n>1$, see [10, Corollary 2.2].
+    If $R$ is a semilocal ring, then either $R$ has infinitely many maximal subrings or $R$ is integral over $\mathbb{Z}_{n}$, where $\operatorname{Char}(R)=n>1$; see [10, Corollary 2.2].
   9. (9)
-    If $R$ is a ring with $|Max(R)|>2^{\aleph_{0}}$ (resp. $|R/J(R)|>2^{2^{\aleph_{0}}}$), then $|RgMax(R)|\geq|Max(R)|$ (resp. $|RgMax(R)|\geq 2^{\aleph_{0}}$), see [10, Proposition2.8].
+    If $R$ is a ring with $|\operatorname{Max}(R)|>2^{\aleph_{0}}$ (resp.
+$|R/J(R)|>2^{2^{\aleph_{0}}}$), then $|\operatorname{RgMax}(R)|\geq|\operatorname{Max}(R)|$ (resp.
+$|\operatorname{RgMax}(R)|\geq 2^{\aleph_{0}}$); see [10, Proposition2.8].
   10. (10)
-    If $R$ is a ring, then $R\times R$ has only finitely many maximal subrings if and only if $R$ is a semilocal ring with only finitely many maximal subrings and $R/M$ is finite for each $M\in Max(R)$, see [10, Corollary 3.10].
-Finally in [10], the authors generalized $(1)$ for a finite direct products of fields too.
-For non-commutative rings, the existence of maximal subring first studied in [1], we will mentioned some of the result of [1] in the next section.
-It is interesting to know that each ring $R$ can be considered as a maximal subring of a larger ring $T$, see [1, Theorem 3.7].
-Recently, in [5] and [6] the existence and the structure of maximal subrings in division rings and the existence of maximal subrings in certain non-commutative rings investigated, respectively.
-In [20] and [7], the authors systematically studied maximal commutative subrings for general noncommutative rings and commutative maximal subrings, respectively.
-We refer the interested reader to [24], for classification of maximal subrings of finite semisimple rings and their applications to the calculation of the covering number of a finite semisimple ring; and also to [3, 4] for more results about the conductor ideals and certain ideals of the extension $R\subseteq T$, where $R$ is a maximal subring of $T$.
+    A ring $R$ satisfies that $R\times R$ has only finitely many maximal subrings if and only if $R$ is a semilocal ring with only finitely many maximal subrings and $R/M$ is finite for each $M\in\operatorname{Max}(R)$; see [10, Corollary 3.10].
+In [10], the authors also generalized statement $(1)$ for a finite direct products of fields.
+For noncommutative rings, the existence of maximal subring was first studied in [1]; we will mention some results from [1] in the next section.
+Interestingly, every ring $R$ can be realized as a maximal subring of some larger ring $T$; see [1, Theorem 3.7].
+Recently, the existence and the structure of maximal subrings in division rings were investigated in[5]; and the existence of maximal subrings in certain noncommutative rings was studied in [6].
+In [20] and [7], the authors systematically examined maximal commutative subrings of general noncommutative rings and commutative maximal subrings, respectively.
+We refer the interested reader to [24] for classification of maximal subrings of finite semisimple rings and their applications to calculating the covering number of a finite semisimple ring; see also [3, 4] for further results on conductor ideals and certain ideals of the extension $R\subseteq T$, where $R$ is a maximal subring of $T$.
+In this paper, we extend the study of maximal subring beyond the commutative setting, focusing on general associative rings with identity.
+Specifically, we investigate conditions under which a ring possesses only finitely many or infinitely many maximal subrings, with particular attention to noncommutative algebras over infinite fields, rings integral over their centers, certain matrix or product constructions.
+Our approach combines ideal-theoretic techniques, structural ring theory, and cardinality arguments, providing a unified framework that generalizes and complements earlier commutative results.
 A brief outline of this paper is as follows.
-In Section $2$, we give some preliminaries results that needed in this paper, specially about the number of maximal left/right ideals of a ring $T$ which are not ideals of $T$.
-In fact, we prove that if $T$ is a ring and $M$ is a maximal left/right ideal of $T$ which is not an ideal of $T$ and $R$ is the idealizer of $M$ in $T$, then $T$ has at least $|R/M|+1$ maximal left/right ideals $N$ such that $N$ is similar to $M$ (i.e., $T/N\cong T/M$ as left/right $T$-module) and therefore $T$ has at least $|R/M|+1$ maximal left/right ideals which are not ideals of $T$.
-Moreover, if $T$ is a $K$-algebra over an infinite field $K$, then either $T$ is a quasi duo ring or $T$ has infinitely many maximal left/right ideals which are not ideals of $T$ (in fact, either $|Max_{l}(T)\setminus Max(T)|\geq|K|$ or $|Max_{r}(T)\setminus Max(T)|\geq|K|$).
-Finally we present some needed results from [27] about the (number of) maximal left/right ideals of $\mathbb{M}_{n}(R)$, where $R$ is a ring and $n\in\mathbb{N}$, and also from [1] about maximal subrings of non-commutative rings.
-In Section $3$, we investigate on the number of maximal subrings in general non-commutative rings.
-First we see that if $D$ is a division ring with finitely many maximal subrings, then the center of $D$ is algebraically closed in $D$.
-For a simple ring $S$, we exactly determine all maximal subrings of $S\times S$.
-We see that if $S$ is a simple ring which is not a division ring, then $S$ has only finitely many maximal subrings if and only if $S$ is a finite ring.
-Consequently, if $S$ is a simple then $S\times S$ has only finitely many maximal subrings if and only if $S$ is a finite ring.
-Next, we see that if a ring $T$ has only finitely many maximal subrings, then $Max_{l}(T)\setminus Max(T)$, $Max_{r}(T)\setminus Max(T)$, $Max(T)\setminus Max_{l}(T)$, $Prm_{l}(T)\setminus Max(T)$ and $Prm_{l}(T)\setminus Max_{l}(T)$ (and analogously for right sets) are all finite, moreover if $M$ is a maximal left/right ideal of $T$ which is not an ideal of $T$, then $\mathbb{I}(M)/M$ is a finite field and if $N$ is a maximal ideal of $T$ which is not a maximal left ideal of $T$, then $T/N\cong\mathbb{M}_{n}(F)$, for some finite field $F$ and $n>1$.
-Also we see that if $T$ is a ring with only finitely many maximal subrings, then $J(T)=\bigcap_{M\in Max(T)}M$.
-In particular, if we assume that $T$ is a $K$-algebra over an infinite field $K$, then either $T$ has infinitely many maximal subrings or $T$ is a quasi duo ring.
-Finally in Section $3$, we prove that if $T$ is an infinite ring with nonzero characteristic, then either $T$ has infinitely many maximal subrings or $\{M\in Max(T)\ |\ |T/M|<\aleph_{0}\}$ is a finite set.
-In Section $4$, we studied rings which are integral over their centers and have only finitely many maximal subrings.
-In particular, algebraic $K$-algebra over an infinite field $K$.
-We prove that if $T$ is a ring which is integral over its center and $T$ has more than $2^{\aleph_{0}}$ maximal (left/right) ideals, then $T$ has infinitely many maximal subrings.
-We observe that, if $T$ is an algebraic $K$-algebra over an infinite field $K$, then either $T$ has infinitely many maximal subring or $T/J(T)$ is a commutative von Neumann regular ring which is integral over $\mathbb{Z}_{p}$, where $p=Char(K)$.
-We see that if $T$ is integral over its center, then either $T$ has infinitely many maximal subrings or $Prm_{l}(T)=Max(T)=Prm_{r}(T)$, and for each maximal ideal $M$ of $T$, either $T/M$ is an absolutely algebraic field (with only finitely many maximal subrings) or $T/M$ is a finite simple ring; in particular, if in addition $J(T)=0$, then $T$ embeds in $S\times\prod_{i\in I}E_{i}$, where $S$ is a finite semisimple ring and each $E_{i}$ is an absolutely algebraic field.
-Finally in Section $5$, we exactly determine when $T=\prod_{i\in I}\mathbb{M}_{n_{i}}(E_{i})$, where each $E_{i}$ is a field and $n_{i}\in\mathbb{N}$, has only finitely many maximal subrings.
-We see that if $R$ is an infinite Artinian ring, then $\mathbb{M}_{n}(R)$, $n>1$, and $R\times R$ have infinitely many maximal subrings.
-All rings in this paper are unital with $1\neq 0$.
-All subrings, modules and homomorphisms are also unital.
-If $R\subsetneq T$ is a ring extension and there exists no other subring between $R$ and $T$, then $R$ is called a maximal subring of $T$, or the extension $R\subseteq T$ is called a minimal ring extension, see [14, 15, 16, 18, 25].
-If $T$ is a ring, $I$ is an ideal of $T$ and $M$ is a left (resp.
-right) $T$-module, then $Min_{T}(I)$, $Max_{r}(T)$, $Max_{l}(T)$, $Max(T)$ and $l.ann_{T}(M)$ (resp.
-$r.ann_{T}(M)$), denote the set of all minimal prime ideals of $I$ in $T$, the set of all maximal right ideals of $T$, the set of all maximal left ideals of $T$, the set of all maximal ideals of $T$ and the left annihilator of $M$ in $T$ (resp.
-the right annihilator of $M$ in $T$), respectively.
-We use $Min(T)$ for $Min_{T}(0)$.
+In Section $2$, we present some preliminary results needed in the sequel, particularly concerning the number of maximal left/right ideals of a ring $T$ that are not two-sided ideals.
+Specifically, we prove that if $T$ is a ring, $M$ is a maximal left/right ideal of $T$ that is not an ideal of $T$, and $R$ is the idealizer of $M$ in $T$, then $T$ has at least $|R/M|+1$ maximal left/right ideals $N$ that are similar to $M$ (i.e., $T/N\cong T/M$ as left/right $T$-modules); consequently, $T$ has at least $|R/M|+1$ maximal left/right ideals that are not ideals of $T$.
+Moreover, if $T$ is a $K$-algebra over an infinite field $K$, then either $T$ is a quasi-duo ring or $T$ has infinitely many maximal left/right ideals that are not ideals of $T$ (indeed, either $|\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)|\geq|K|$ or $|\operatorname{{Max}_{r}}(T)\setminus\operatorname{Max}(T)|\geq|K|$).
+Finally we recall several required results from [27] regarding the number of maximal left/right ideals of $\mathbb{M}_{n}(R)$ (where $R$ is a ring and $n\in\mathbb{N}$), as well as from [1] concerning maximal subrings of noncommutative rings.
+In Section $3$, we investigate the number of maximal subrings in general noncommutative rings.
+First, we show that if $D$ is a division ring with only finitely many maximal subrings, then the center of $D$ is algebraically closed in $D$.
+For a simple ring $S$, we precisely determine all maximal subrings of $S\times S$.
+We prove that if $S$ is a simple ring that is not a division ring, then $S$ has only finitely many maximal subrings if and only if $S$ is finite.
+Consequently, for a simple ring $S$, the ring $S\times S$ has only finitely many maximal subrings if and only if $S$ is a finite ring.
+Next, we demonstrate that if a ring $T$ has only finitely many maximal subrings, then $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$, $\operatorname{{Max}_{r}}(T)\setminus\operatorname{Max}(T)$, $\operatorname{Max}(T)\setminus\operatorname{{Max}_{\ell}}(T)$, $\operatorname{{Prm}_{\ell}}(T)\setminus\operatorname{Max}(T)$ and $\operatorname{{Prm}_{\ell}}(T)\setminus\operatorname{{Max}_{\ell}}(T)$ (and their right analogous) are all finite.
+Furthermore, if $M$ is a maximal left/right ideal of $T$ that is not an ideal of $T$, then $\mathbb{I}(M)/M$ is a finite field; and if $N$ is a maximal ideal of $T$ that is not a maximal left ideal of $T$, then $T/N\cong\mathbb{M}_{n}(F)$, for some finite field $F$ and some $n>1$.
+We also show that, if $T$ has only finitely many maximal subrings, then $J(T)=\bigcap_{M\in\operatorname{Max}(T)}M$.
+In particular, if $T$ is a $K$-algebra over an infinite field $K$, then either $T$ has infinitely many maximal subrings or $T$ is a quasi-duo ring.
+Finally in Section $3$, we prove that if $T$ is an infinite ring of nonzero characteristic, then either $T$ has infinitely many maximal subrings or the set $\{M\in\operatorname{Max}(T)\ |\ |T/M|<\aleph_{0}\}$ is finite.
+Section $4$, is devoted to rings that are integral over their centers and have only finitely many maximal subrings, with special attention to algebraic $K$-algebra over an infinite field $K$.
+We prove that if $T$ is integral over its center and $T$ has more than $2^{\aleph_{0}}$ maximal (left/right) ideals, then $T$ has infinitely many maximal subrings.
+Moreover, if $T$ is an algebraic $K$-algebra over an infinite field $K$, then either $T$ has infinitely many maximal subring or $T/J(T)$ is a commutative von Neumann regular ring that is integral over $\mathbb{Z}_{p}$, where $p=\operatorname{Char}(K)$.
+We also establish that if $T$ is integral over its center, then either $T$ has infinitely many maximal subrings or $\operatorname{{Prm}_{\ell}}(T)=\operatorname{Max}(T)=\operatorname{{Prm}_{r}}(T)$, and for each maximal ideal $M$ of $T$, either $T/M$ is an absolutely algebraic field (with only finitely many maximal subrings) or $T/M$ is a finite simple ring.
+In particular, if additionally $J(T)=0$, then $T$ embeds into $S\times\prod_{i\in I}E_{i}$, where $S$ is a finite semisimple ring and each $E_{i}$ is an absolutely algebraic field.
+Finally in Section $5$, we precisely determine when a ring of the form $T=\prod_{i\in I}\mathbb{M}_{n_{i}}(E_{i})$ (with each $E_{i}$ is a field and $n_{i}\in\mathbb{N}$), has only finitely many maximal subrings.
+We also prove that if $R$ is an infinite Artinian ring, then $\mathbb{M}_{n}(R)$, for $n>1$, as well as $R\times R$, have infinitely many maximal subrings.
+All rings considered are unital with $1\neq 0$.
+All subrings, modules and homomorphisms are assumed to be unital.
+If $R\subsetneq T$ is a ring extension and such that no other subring lies strictly between $R$ and $T$, then $R$ is called a maximal subring of $T$; equivalently, the extension $R\subseteq T$ is called a minimal ring extension, see [14, 15, 16, 18, 25].
+Let $T$ be a ring and $M$ be a left $T$-module, then $\operatorname{{Max}_{r}}(T)$, $\operatorname{{Max}_{\ell}}(T)$, $\operatorname{Max}(T)$ and $\operatorname{l.ann}_{T}(M)$, denote the set of all maximal right ideals of $T$, the set of all maximal left ideals of $T$, the set of all maximal ideals of $T$ and the left annihilator of $M$ in $T$, respectively.
 The set of all left (resp.
-right) primitive ideals of $T$ is denoted by $Prm_{l}(T)$ (resp.
-$Prm_{r}(T)$).
-The characteristic of a ring $T$ is denoted by $Char(T)$.
-If $T$ is a ring, then $dim(T)$ denotes the classical Krull dimension of $T$ (i.e., the supremum of the lengths of all chains of prime ideals of $T$).
-Also we use $Aut(T)$ for the set of all ring automorphism of a ring $T$.
-$v.dim_{K}(T)$ denotes the vector dimension of a $K$-algebra $T$ over a field $K$.
-$J(T)$ denotes the Jacobson radical of a ring $T$.
-For a ring $T$, $Nil_{*}(T)$, $Nil^{*}(T)$, $N(T)$ and $U(T)$ are the lower nilradical, the upper nilradical, the set of all nilpotent elements and the units group of $T$, respectively.
-If $T$ is a ring, the opposite ring of $T$ is denoted by $T^{op}$.
-A prime ideal $P$ of a ring $T$ is called strongly (resp.
-completely) prime if $T/P$ has no nonzero nil ideal (resp.
-$T/P$ is a domain).
-If $T$ is a ring, then $\mathbb{M}_{n}(T)$ denote the ring of all $n\times n$ square matrices over $T$.
-If $M$ is a left (resp.
-right) module over a ring $T$, then the ring of all $T$-module homomorphisms of $M$ is denoted by $End({}_{T}M)$ (resp.
-$End(M_{T})$).
-If $A$ is a right ideal of a ring $T$, then $\mathbb{I}(A)$ denotes the idealizer of $A$ in $T$, that is the largest subring of $T$ which $A$ is a two-sided ideal in it, i.e., $\mathbb{I}(A)=\{t\in T\ |\ tA\subseteq A\}$, the similar notation is used when $A$ is a left ideal of $T$.
-If $A$ is a right ideal of a ring $T$, then $\mathbb{E}_{T}(A):=\mathbb{I}(A)$ is called the eigenring of $A$.
-If $X$ is a subset of a ring $T$, then $C_{T}(X)$ is the centralizer of $X$ in $T$, in particular, $C(T)=C_{T}(T)$ is the center of $T$.
-For a ring $T$, the subring $\{n.1_{T}\ |\ n\in\mathbb{Z}\}$ is called the prime subring of $T$.
-It is clear that $T$ and $T^{op}$ have a same prime subring (moreover, $C(T)=C(T^{op})$) and therefore have a same characteristic.
+right) primitive ideals of $T$ is denoted by $\operatorname{{Prm}_{\ell}}(T)$ (resp.
+$\operatorname{{Prm}_{r}}(T)$).
+The characteristic of a ring $T$ written as $\operatorname{Char}(T)$.
+We use $\operatorname{Aut}(T)$ for the group of all ring automorphisms of $T$.
+The Jacobson radical of $T$ is denoted by $J(T)$.
+We let $N(T)$ and $U(T)$ be the set of all nilpotent elements and the units group of $T$, respectively.
+The opposite ring of $T$ is written by $T^{\text{op}}$.
+The ring of all $n\times n$ matrices over $T$ is denoted by $\mathbb{M}_{n}(T)$.
+The ring of all $T$-module homomorphisms of $M$ is written as $\operatorname{End}_{T}(M)$.
+For a right ideal $A$ of $T$, the idealizer of $A$ in $T$ is denoted by $\mathbb{I}(A)$, it is the largest subring of $T$ in which $A$ is a two-sided ideal, i.e., $\mathbb{I}(A)=\{t\in T\ |\ tA\subseteq A\}$.
+The same notation is used when $A$ is a left ideal of $T$.
+In any case, the ring $\mathbb{E}_{T}(A):=\mathbb{I}(A)/A$ is called the eigenring of $A$.
+If $X$ is a subset of $T$, then $C_{T}(X)$ denotes the centralizer of $X$ in $T$; in particular, $C(T)=C_{T}(T)$ is the center of $T$.
+The subring $\{n.1_{T}\ |\ n\in\mathbb{Z}\}$ is called the prime subring of $T$.
 A ring $T$ is called left (resp.
-right) quasi duo if each maximal left (resp.
-right) ideal of $T$ is a two-sided ideal of $T$, see [22].
-$T$ is called quasi duo, if $T$ is left and right quasi duo ring.
-A ring $T$ is called left (resp.
-right) duo if each left (resp.
+right) quasi-duo if any maximal left (resp.
 right) ideal of $T$ is two-sided.
-Similarly duo rings are defined.
-If $R$ is a subring of a ring $T$ and $t\in T$, then we say that $t$ is left (resp.
-right) integral over $R$, if $t$ is a root of a left (resp.
-right) monic polynomial of degree $n\geq 1$ over $R$.
-$R$ is called left (resp.
-right) integrally closed in $T$, if every left (resp.
-right) integral element of $T$ over $R$, belongs to $R$.
-$R$ is called integrally closed in $T$, whenever $R$ is left and right integrally closed in $T$.
-If $T$ is a ring and $R$ is a subring of the center of $T$, then $T$ is called an $R$-algebra.
-It is clear that, if $T$ is an $R$-algebra, then the definitions of right and left integral elements and also left or right integrally closed for the extension $R\subseteq T$ are coincided.
-If $F\subseteq E$ is a field extension, then $tr.deg(E/F)$ denotes the transcendence degree of $E$ over $F$.
-For other notations and definitions we refer the reader to [19, 21, 26].
+If $T$ is both left and right quasi-duo, it is called a quasi-duo ring, see [22].
+Let $R$ be a subring of $T$ and $t\in T$.
+We say that $t$ is left (resp.
+right) integral over $R$, if $t$ is a root of a monic polynomial of degree $n\geq 1$ with coefficients on the left (resp.
+right) side from $R$.
+The subring $R$ is left (resp.
+right) integrally closed in $T$, if every element of $T$ that is left (resp.
+right) integral element over $R$, belongs to $R$.
+We say it is integrally closed in $T$, if $R$ is both left and right integrally closed in $T$.
+When $T$ is a ring and $R$ is a subring of $C(T)$, we call $T$ is an $R$-algebra.
+In this case, the notation of right and left integral elements (as well as, left and right integrally closed) for the extension $R\subseteq T$ coincide.
+For a field extension $F\subseteq E$, we write $\operatorname{tr.deg}(E/F)$ for the transcendence degree of $E$ over $F$.
+Other notations and definitions follow [19, 21, 26].
 2. Preliminaries
 Let $T$ be a ring and $I$ and $J$ be left ideals of $T$.
-Then $I$ and $J$ are called similar left ideals if $T/I\cong T/J$ as left $T$-module.
-If $a\in T$ and define $(I:a):=\{x\in T\ |\ xa\in I\}$, then it is not hard to see that $I$ and $J$ are similar if and only if there exists $c\in T$ such that $I+Tc=T$ and $J=(I:c)$, see [13, Proposition 1.3.6].
-Since for a left ideal $I$ of a ring $T$, we have the ring isomorphism $\mathbb{I}(I)/I\cong(End({}_{T}(T/I)))^{op}$, see [23, P.
-14, 1.11], we immediately conclude that if $I$ and $J$ are similar left ideals of a ring $T$, then $\mathbb{I}(I)/I\cong\mathbb{I}(J)/J$ as rings, i.e., up to ring isomorphism, the ring $\mathbb{I}(I)/I$ is unique for similar left ideals; in particular $|\mathbb{I}(I)/I|=|\mathbb{I}(J)/J|$.
-In other words the cardinality of the ring $\mathbb{I}(I)/I$ is invariant up to similarity.
-Moreover, if $M$ is a maximal left ideal of a ring $T$, then we have the following facts too:
+We say $I$ and $J$ are similar left ideals if $T/I\cong T/J$ as left $T$-modules.
+If $a\in T$, define $(I:a):=\{x\in T\ |\ xa\in I\}$, then it is not hard to see that $I$ and $J$ are similar if and only if there exists $c\in T$ such that $I+Tc=T$ and $J=(I:c)$, see [13, Proposition 1.3.6].
+For a left ideal $I$ of $T$, we have the ring isomorphism $\mathbb{I}(I)/I\cong(\operatorname{End}_{T}(T/I))^{\text{op}}$, see [23, P.
+14, 1.11].
+Consequently, if $I$ and $J$ are similar left ideals of $T$, then $\mathbb{I}(I)/I\cong\mathbb{I}(J)/J$ as rings.
+Thus, up to ring isomorphism, the ring $\mathbb{I}(I)/I$ is uniquely determined by similarity classes; in particular $|\mathbb{I}(I)/I|=|\mathbb{I}(J)/J|$.
+In other words, the cardinality of $\mathbb{I}(I)/I$ is invariant under similarity.
+Now, let $M$ be a maximal left ideal of $T$.
+The following facts are useful:
   1. (1)
@@
   2. (2)
-    if $N$ is a (maximal) left ideal of $T$ which is similar to $M$, then $N$ is an ideal of $T$ if and only if $N=M$ if and only if $M$ is an ideal of $T$ .
-To see the above facts, note that $(1)$ is an immediate consequences of [13, Proposition 1.3.6].
-For $(2)$, let $M$ and $N$ be similar maximal left ideals of $T$ and assume that $N$ is an ideal of $T$.
-Since $T/N\cong T/M$ as left $T$-module we immediately conclude that $N=l.ann_{T}(T/N)=l.ann_{T}(T/M)\subseteq M$ and therefore $M=N$ (note that $N$ is a maximal left ideal of $T$, for $T/N\cong T/M$ as left $T$-module).
+    If $N$ is a (maximal) left ideal of $T$ similar to $M$, then $N$ is an ideal of $T$ if and only if $N=M$, which in turn holds if and only if $M$ is an ideal of $T$.
+To see the previous facts, note that $(1)$ follows directly from [13, Proposition 1.3.6], because obviously $M+Tc=T$.
+For $(2)$, assume that $N$ and $M$ are similar left ideals and that $N$ is an ideal.
+Since $T/N\cong T/M$ as left $T$-modules, we have $N=\operatorname{l.ann}_{T}(T/N)=\operatorname{l.ann}_{T}(T/M)\subseteq M$.
+Because $N$ is a maximal left ideal, the inclusion $N\subseteq M$ forces $N=M$.
 Hence $(2)$ holds.
-Consequently, if $M$ is a maximal left ideal of $T$ which is not an ideal of $T$, and $[M]$ be the set of all (maximal) left ideals of $T$ which are similar to $M$, we deduce that $M]\subseteq Max_{l}(T)\setminus Max(T)$, by $(2)$, i.e., each (maximal) left ideals of $T$ which is similar to $M$ is not an ideal of $T$.
-Clearly, the relation similarity is an equivalence relation on the set of all left ideals of $T$ and in particular on $Max_{l}(T)$.
-By $(1)$ and $(2)$, it is obvious that if $M$ is a maximal left ideal of $T$, then $|[M]|=1$ (i.e., $M]=\{M\}$) if and only if $M$ is an ideal of $T$ (thus $M\in Max(T)\cap Max_{r}(T)$).
-It is clear that whenever $I$ is a left ideal of a ring $T$ and $c\in T$, then $(I:c)=T$ if and only if $c\in I$.
-Moreover, $I\subseteq(I:c)$ if and only if $c\in\mathbb{I}(I)$.
-In particular, if $M$ is a maximal left ideal of a ring $T$, from these facts we immediately conclude that $(M:c)=M$ if and only if $c\in\mathbb{I}(M)\setminus M$.
-Hence $M]=\{M\}\cup\{(M:c)\ |\ c\in T\setminus\mathbb{I}(M)\}$.
-In the following main result we find a lower bound for $|[M]|$.
-Before present it, let us remind that from the previous observation for each $N\in[M]$, we have $|\mathbb{I}(M)/M|=|\mathbb{I}(N)/N|$.
-Now the following is in order.
+Consequently, if $M$ is a maximal left ideal of $T$ that is not an ideal of $T$, and if $[M]$ denotes the set of all (maximal) left ideals of $T$ similar to $M$, then by $(2)$ we have $M]\subseteq\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$, i.e., every (maximal) left ideal in $[M]$ fail to be a two-ideal.
+Clearly, similarity is an equivalence relation on the set of all left ideals of $T$, and in particular on $\operatorname{{Max}_{\ell}}(T)$.
+From $(1)$ and $(2)$, it is evident that for a maximal left ideal $M$, we have $|[M]|=1$ (i.e., $M]=\{M\}$) if and only if $M$ is an ideal of $T$ (and therefore in this case $M\in\operatorname{Max}(T)\cap\operatorname{{Max}_{r}}(T)$).
+It is clear that for any left ideal $I$ of a ring $T$ and any $c\in T$, we have $(I:c)=T$ if and only if $c\in I$.
+Moreover, $I\subseteq(I:c)$ holds if and only if $c\in\mathbb{I}(I)$.
+In particular, if $M$ is a maximal left ideal of $T$, these observations immediately imply that $(M:c)=M$ if and only if $c\in\mathbb{I}(M)\setminus M$.
+Consequently, $M]=\{M\}\cup\{(M:c)\ |\ c\in T\setminus\mathbb{I}(M)\}$.
+Before starting our main result (which present a lower bound for $|[M]|$), we recall that, as noted earlier, for any $N\in[M]$, we have $|\mathbb{I}(M)/M|=|\mathbb{I}(N)/N|$.
+We now present the following key theorem.
 Theorem 2.1.
-Let $T$ be a ring and $M\in Max_{l}(T)$ is not an ideal of $T$ (i.e., $M$ is not a maximal ideal of $T$).
+Assume that $T$ is a ring and $M\in\operatorname{{Max}_{\ell}}(T)$ is not an ideal of $T$ (i.e., $M$ is not a maximal ideal of $T$).
 Then $|[M]|\geq|\mathbb{I}(M)/M|+1$.
-In particular, $|Max_{l}(T)\setminus Max(T)|\geq|\mathbb{I}(M)/M|+1$.
+In particular, $|\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)|\geq|\mathbb{I}(M)/M|+1$.
 Proof.
 Let $R=\mathbb{I}(M)$ and $D=R/M$.
-Since $M$ is not an ideal of $T$ we conclude that $R$ is a proper subring of $T$.
+Since $M$ is not an ideal of $T$ we infer that $R$ is a proper subring of $T$.
 Take a fixed element $x\in T\setminus R$.
@@
 Let $t\in(M:x+a)=(M:x+b)$, then we conclude that $tx+ta,tx+tb\in M$.
-Therefore $t(a-b)\in M$, i.e., $(M:x+a)\subseteq(M:a-b)$.
-Now if $a-b\notin M$, then we conclude that $a-b\in R\setminus M$ and therefore (M:ab)=M(M:a-b)=M.
-Hence $(M:x+a)\subseteq M$ and therefore by maximality of $(M:x+a)$ we obtain that $(M:x+a)=M$ which is absurd, for $x+a\in T\setminus R$.
-Thus $a-b\in M$ and therefore we are done.
+It follows that $t(a-b)\in M$, i.e., $(M:x+a)\subseteq(M:a-b)$.
+Now if $a-b\notin M$, then we obtain that $a-b\in R\setminus M$, a fortiori (M:ab)=M(M:a-b)=M.
+This implies that $(M:x+a)\subseteq M$ and consequently by maximality of $(M:x+a)$ we deduce that $(M:x+a)=M$ which is absurd, because $x+a\in T\setminus R$.
+Thus $a-b\in M$, as desired.
 ∎
-As an immediate consequence of the previous theorem is that if $T$ is a ring and $Max_{l}(T)\setminus Max(T)\neq\emptyset$, then $|Max_{l}(T)\setminus Max(T)|\geq 3$, i.e., if a ring $T$ has a maximal left ideal which is not an ideal of $T$, then $T$ has at least three maximal left ideals which are not ideals of $T$.
+An immediate consequence of the previous theorem is that if $T$ is a ring and $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)\neq\emptyset$, then $|\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)|\geq 3$.
+In other words, if a ring $T$ possesses a maximal left ideal that is not a two-ideal, then $T$ has at least three distinct maximal left ideals that are not two-ideals.
 Corollary 2.2.
 Let $T$ be a $K$-algebra over a field $K$.
-Then either $T$ is a left quasi duo ring or $T$ has at least $|K|+1$ maximal left ideals which are not ideals of $T$ (i.e., $|Max_{l}(T)\setminus Max(T)|\geq|K|+1$).
+Then either $T$ is a left quasi-duo ring or $T$ has at least $|K|+1$ maximal left ideals which are not ideals of $T$ (i.e., $|\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)|\geq|K|+1$).
 Proof.
-Assume that $T$ is not a left quasi duo ring and $M$ be a maximal left ideal of $T$ which is not an ideal of $T$.
+Assume that $T$ is not a left quasi-duo ring and $M$ is a maximal left ideal of $T$ which is not an ideal of $T$.
 Since $K$ is a field which is contained in the center of $T$, we conclude that $K\subseteq R:=\mathbb{I}(M)$ and clearly $K\cap M=0$.
-Thus $K$ embeds in $R/M$ and hence we are done by Theorem 2.1.
+Thus $K$ embeds in $R/M$ and we are done by Theorem 2.1.
 ∎
@@
 Corollary 2.3.
-Let $T$ be an algebra over an infinite field $K$. Then either $T$ is a left quasi duo ring or $Max_{l}(T)\setminus Max(T)$ is infinite (in fact, $|Max_{l}(T)\setminus Max(T)|\geq|K|$).
-Thus, if $T$ is a ring and $Max_{l}(T)\setminus Max(T)$ is finite, then for each $M\in Max_{l}(T)\setminus Max(T)$, we deduce that $\mathbb{I}(M)/M$ is a finite field.
-Moreover, if in addition $T$ is a $K$-algebra over a field $K$ and $Max_{l}(T)\setminus Max(T)$ is finite, then $K$ is a finite field (note that $\mathbb{I}(M)/M$ contains a copy of $K$, for each maximal left ideal $M$ of $T$).
-By Theorem 2.1, note that if $Max_{l}(T)\setminus Max(T)=\bigcup_{i\in I}[M_{i}]$ be the partition of $Max_{l}(T)\setminus Max(T)$ in to similarity classes, then $|Max_{l}(T)\setminus Max(T)|\geq|I|+\sum_{i\in I}|\mathbb{I}(M)/M|$.
-It is clear that if $K$ is a field and $n>1$, then by the above results we have $|Max_{l}(\mathbb{M}_{n}(K))|\geq|K|+1$ and $|Max_{r}(\mathbb{M}_{n}(K))|\geq|K|+1$ (note, $\mathbb{M}_{n}(K)$ is a simple $K$-algebra).
-In particular, if $K$ is an infinite field then $\mathbb{M}_{n}(K)$ has at least $|K|$-many maximal left/right ideals (also see Remark 2.5).
-Next we want to use the notation and results of [27], to show that if $R$ is an infinite division ring and $n>1$, then $T=\mathbb{M}_{n}(R)$ has infinitely many maximal left/right ideals.
-To see this, note that if $M$ is a maximal left ideal of a ring $R$ and $T=\mathbb{M}_{n}(R)$, then for each $u\in R^{n}\setminus M^{n}$, the set $D(M,u):=\{X\in T\ |\ Xu\in M^{n}\}$ is a maximal left ideal of $T$ (in fact, $T/D(M,u)\cong(R/M)^{n}$ as left $T$-module).
-Moreover, $Max_{l}(T)=\{D(M,u)\ |\ M\in Max_{l}(R),\ u\in R^{n}\setminus M^{n}\}$, see [27, Theorem 1.2].
-In particular, if $R$ is a division ring, then we deduce that $Max_{l}(\mathbb{M}_{n}(R))=\{D(0,u)\ |\ 0\neq u\in R^{n}\}$.
-Hence to show that for an infinite division ring $R$, $Max_{l}(\mathbb{M}_{n}(R))$ is infinite we must prove that there exist infinitely many $0\neq u\in R^{n}$ such that $D(0,u)$ are distinct maximal left ideals.
-By [27, Proposition 2.3], if $M$ is a maximal left ideal of a ring $R$, and $S=\mathbb{I}(M)$ be the idealizer of $M$ in $R$, and $u,v\in S^{n}$, then $D(M,u)=D(M,v)$ (in the ring $T=\mathbb{M}_{n}(R)$) if and only if there exists $c\in S\setminus M$ such that $v\equiv uc\ mod(M)$ (i.e.
-$v-uc\in M^{n}$).
-In particular, if $R$ is a division ring, since the only possible for $M$ is the zero ideal and in this case $S=R$, we deduce that for each $u,v\in R^{n}$, $D(0,u)=D(0,v)$ if and only if there exists $0\neq c\in R$ such that $v=uc$, (in other words $v$ and $u$ are right parallel vectors in $R^{n}$ as right $R$-module).
-Clearly, if $D$ is an infinite division ring and $n>1$, then $D^{n}$ has infinitely many right non-parallel vectors, thus we deduce the following fact.
+Let $T$ be a $K$-algebra over an infinite field $K$.
+Then either $T$ is a left quasi-duo ring or $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$ is infinite (in fact, $|\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)|\geq|K|$).
+Thus, if $T$ is a ring and $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$ is finite, then for each $M\in\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$, we deduce that $\mathbb{I}(M)/M$ is a finite field.
+Moreover, if additionally, $T$ is a $K$-algebra over a field $K$ and $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$ is finite, then $K$ must be a finite field (note that for any maximal left ideal $M$ of $T$, the field $\mathbb{I}(M)/M$ contains a copy of $K$).
+Let $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)=\bigcup_{i\in I}[M_{i}]$ be the partition of $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$ into similarity classes.
+By Theorem 2.1, we obtain that $|\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)|\geq|I|+\sum_{i\in I}|\mathbb{I}(M)/M|$.
+It is clear from the above results that if $K$ is a field and $n>1$, then $|\operatorname{{Max}_{\ell}}(\mathbb{M}_{n}(K))|\geq|K|+1$ and $|\operatorname{{Max}_{r}}(\mathbb{M}_{n}(K))|\geq|K|+1$ (note, $\mathbb{M}_{n}(K)$ is a simple $K$-algebra and therefore each maximal left/right ideal of $\mathbb{M}_{n}(K)$ is not an ideal of $\mathbb{M}_{n}(K)$).
+In particular, if $K$ is an infinite field, then $\mathbb{M}_{n}(K)$ has at least $|K|$-many maximal left/right ideals (also see Remark 2.5).
+Next we employ the notation and results of [27] to show that if $R$ is an infinite division ring and $n>1$, then $T=\mathbb{M}_{n}(R)$ possesses infinitely many maximal left/right ideals.
+Recall that if $M$ is a maximal left ideal of a ring $R$ and $T=\mathbb{M}_{n}(R)$, then for each $u\in R^{n}\setminus M^{n}$, the set $D(M,u):=\{X\in T\ |\ Xu\in M^{n}\}$ is a maximal left ideal of $T$; indeed, $T/D(M,u)\cong(R/M)^{n}$ as left $T$-modules.
+Moreover, $\operatorname{{Max}_{\ell}}(T)=\{D(M,u)\ |\ M\in\operatorname{{Max}_{\ell}}(R),\ u\in R^{n}\setminus M^{n}\}$, see [27, Theorem 1.2].
+In particular, when $R$ is a division ring, we obtain that $\operatorname{{Max}_{\ell}}(\mathbb{M}_{n}(R))=\{D(0,u)\ |\ 0\neq u\in R^{n}\}$.
+Hence to show that $\operatorname{{Max}_{\ell}}(\mathbb{M}_{n}(R))$ is infinite for an infinite division ring $R$, it suffices to produce infinitely many nonzero vectors $u\in R^{n}$ for which $D(0,u)$ are distinct.
+By [27, Proposition 2.3], if $M$ is a maximal left ideal of a ring $R$, $S=\mathbb{I}(M)$ its idealizer of $M$ in $R$, and $u,v\in S^{n}$, then $D(M,u)=D(M,v)$ (in $T=\mathbb{M}_{n}(R)$) if and only if there exists $c\in S\setminus M$ such that $v\equiv uc\ \mod(M)$, i.e., $v-uc\in M^{n}$.
+In particular, for a division ring $R$ the only possible maximal left ideals is $M=0$, so that $S=R$; consequently, for $u,v\in R^{n}$, we have $D(0,u)=D(0,v)$ if and only if there exists $0\neq c\in R$ such that $v=uc$.
+In other words $v$ and $u$ are right-parallel vectors in the right $R$-module $R^{n}$.
+Clearly, if $R$ is an infinite division ring and $n>1$, then as right $R$-module $R^{n}$ contains infinitely many right non-parallel vectors.
+This observations yield the following quick corollary.
 Corollary 2.4.
-Let $R$ be an infinite division ring and $n>1$, then $T=\mathbb{M}_{n}(R)$ has infinitely many maximal left (resp. right) ideals.
+Let $R$ be an infinite division ring and $n>1$. Then $T=\mathbb{M}_{n}(R)$ has infinitely many maximal left (resp. right) ideals.
 Remark 2.5.
-Let $R$ be a ring, $n>1$ and $T=\mathbb{M}_{n}(R)$.
-Then it is obvious that $Max_{l}(T)\cap Max(T)=\emptyset=Max_{r}(T)\cap Max(T)$.
-Moreover, the function $M\longmapsto M^{t}$ (where $M^{t}$ is the transpose of $M$, i.e., $M^{t}:=\{m^{t}\ |\ m\in M\}$), is a one-one corresponding from $Max_{l}(T)$ onto $Max_{r}(T)$.
-In particular, $|Max_{l}(T)|=|Max_{r}(T)|$.
+Let $R$ be a ring, $n>1$, and $T=\mathbb{M}_{n}(R)$.
+It is obvious that $\operatorname{{Max}_{\ell}}(T)\cap\operatorname{Max}(T)=\emptyset=\operatorname{{Max}_{r}}(T)\cap\operatorname{Max}(T)$.
+Moreover, the map $M\longmapsto M^{t}$ (where $M^{t}:=\{m^{t}\ |\ m\in M\}$ is the transpose of $M$), is a one-to-one correspondence from $\operatorname{{Max}_{\ell}}(T)$ onto $\operatorname{{Max}_{r}}(T)$.
+In particular, $|\operatorname{{Max}_{\ell}}(T)|=|\operatorname{{Max}_{r}}(T)|$.
 Remark 2.6.
-Let $T$ be a ring, then we have the following immediate facts:
+For any ring $T$, we have the following immediate facts:
   1. (1)
-    $Max_{l}(T)\cap Max(T)=Max_{r}(T)\cap Max(T)=Prm_{l}(T)\cap Max_{l}(T)=Prm_{r}(T)\cap Max_{r}(T)=Prm_{l}(T)\cap Max_{r}(T)=Prm_{r}(T)\cap Max_{l}(T)$.
+    $\operatorname{{Max}_{\ell}}(T)\cap\operatorname{Max}(T)=\operatorname{{Max}_{r}}(T)\cap\operatorname{Max}(T)=\operatorname{{Prm}_{\ell}}(T)\cap\operatorname{{Max}_{\ell}}(T)=\operatorname{{Prm}_{r}}(T)\cap\operatorname{{Max}_{r}}(T)=\operatorname{{Prm}_{\ell}}(T)\cap\operatorname{{Max}_{r}}(T)=\operatorname{{Prm}_{r}}(T)\cap\operatorname{{Max}_{\ell}}(T)$.
   2. (2)
-    $Max(T)\setminus Max_{l}(T)=Max(T)\setminus Max_{r}(T)$.
-Next let us remind the reader some immediate fact from maximal subrings and also some needed result from [1].
-It is easy to see that a proper subring $S$ of a ring $T$ is maximal if and only if for each $x\in T\setminus S$, we have $S[x]=T$.
-Moreover, if $R$ is a proper subring of $T$ and there exists $\alpha\in T$ such that $R[\alpha]=T$, then by a natural use of Zorn’s Lemma, $T$ has a maximal subring $S$ with $R\subseteq S$ and $\alpha\notin S$.
-In particular, if $R\subsetneq T$ is a ring extension which is finite as ring (module) over $R$, then $T$ has a maximal subring which contains $R$.
+    $\operatorname{Max}(T)\setminus\operatorname{{Max}_{\ell}}(T)=\operatorname{Max}(T)\setminus\operatorname{{Max}_{r}}(T)$.
+Next, let we recall some basic facts from maximal subrings, as well as several needed results from [1].
+A proper subring $S$ of a ring $T$ is maximal if and only if for every $x\in T\setminus S$, one has $S[x]=T$.
+Moreover, if $R$ is a proper subring of $T$ and there exists $\alpha\in T$ such that $R[\alpha]=T$, then by a standard application of Zorn’s Lemma, $T$ possesses a maximal subring $S$ with $R\subseteq S$ and $\alpha\notin S$.
+In particular, if $R\subsetneq T$ is a ring extension which is finite as ring (or as module) over $R$, then $T$ has a maximal subring containing $R$.
 Now let $R$ be a ring and $\Delta:=\{(r,r)\ |\ r\in R\}$, be the diagonal subring of $T:=R\times R$.
-Then it is not hard too see that $\Delta[(1,0)]=T$.
-Therefore $T$ has a maximal subring which contains $\Delta$.
-In fact, subrings of $T$ which contains $\Delta$ has the form $\Delta(I):=\Delta+I\times I$, where $I$ is an ideal of $R$.
-In particular, $\Delta(I)$ is a maximal subring of $T$ if and only if $I$ is a maximal ideal of $R$, see [1, Theorem 2.7].
-Therefore a ring $R$ is simple if and only if $\Delta$ is a maximal subring of $R\times R$.
-In the following we give some results from [1].
+It is easy to see that $\Delta[(1,0)]=T$.
+Consequently, $T$ has a maximal subring containing $\Delta$.
+In fact, every subring $S$ of $T$ that contains $\Delta$ has the form $\Delta(I):=\Delta+I\times I$, where $I$ is an ideal of $R$.
+Moreover, $\Delta(I)$ is a maximal subring of $T$ if and only if $I$ is a maximal ideal of $R$, see [1, Theorem 2.7].
+Therefore a ring $R$ is simple exactly when $\Delta$ is a maximal subring of $R\times R$.
+The following results from [1] will be useful.
 Theorem 2.7.
@@
     [1, Theorem 4.1].
-Let $T$ be a ring and $A$ be a maximal right/left ideal of $T$ which is not an ideal of $T$.
+Let $T$ be a ring and $A$ be a maximal right/left ideal of $T$ that is not an ideal of $T$.
 Then the idealizer of $A$ is a maximal subring of $T$.
-In particular, either a ring has a maximal subring or is a quasi duo ring.
+In particular, every ring either has a maximal subring or is a quasi-duo ring.
   2. (2)
-    [1, Proposition 4.2]. Let $R$ be a maximal subring of a ring $T$ which contains a maximal one-sided ideal $A$ of $T$ which is not an ideal of $T$. Then $R$ is the idealizer of $A$ in $T$.
+    [1, Proposition 4.2].
+Let $R$ be a maximal subring of a ring $T$ that contains a maximal left/right ideal $A$ of $T$ that is not two-sided ideal of $T$.
+Then $R$ is precisely the idealizer of $A$ in $T$.
   3. (3)
-    [1, Theorem 4.4]. Let $T$ be a ring which is not a division ring. If $T$ is left/right primitive ring (in particular, if $T$ is a simple ring), then $T$ has a maximal subring.
+    [1, Theorem 4.4]. Let $T$ be a ring that is not a division ring. If $T$ is left/right primitive ring (in particular, if $T$ is simple), then $T$ has a maximal subring.
   4. (4)
-    [1, Theorem 3.1]. For each ring $R$ and $n>1$, the matrix ring $\mathbb{M}_{n}(R)$ has a maximal subring.
-Finally, we prove some needed facts as follows.
-We remind the reader that if $T$ is a ring with $|T|<\mathfrak{a}$, where $\mathfrak{a}$ is an infinite cardinal number, then for each finitely generated left/right $T$-module $M$, we have $|M|<\mathfrak{a}$.
-To see this, note that since $M$ is finitely generated as $T$-module, we infer that there exists a natural number $n$ such that $M$ is a homomorphic image of $T^{n}$ (as left/right $T$-module).
-Hence $|M|\leq|T|^{n}<\mathfrak{a}^{n}=\mathfrak{a}$.
-Now the following is in order.
+    [1, Theorem 3.1]. For every ring $R$ and every $n>1$, the matrix ring $\mathbb{M}_{n}(R)$ has a maximal subring.
+Finally, we prove several auxiliary facts.
+Recall that if $T$ is a ring with $|T|<\mathfrak{a}$, where $\mathfrak{a}$ is an infinite cardinal number, then for each finitely generated left/right $T$-module $M$, we have $|M|<\mathfrak{a}$.
+Indeed, if $M$ is finitely generated, there exists a natural number $n$ such that $M$ is a homomorphic image of $T^{n}$ (as left/right $T$-module); hence $|M|\leq|T|^{n}<\mathfrak{a}^{n}=\mathfrak{a}$.
 Lemma 2.8.
-Let $T$ be a left Noetherian ring and $I$ be a nilpotent ideal of $T$. If $\mathfrak{a}$ be an infinite cardinal number, then $|T|<\mathfrak{a}$ if and only if $|T/I|<\mathfrak{a}$.
+Let $T$ be a left Noetherian ring and $I$ a nilpotent ideal of $T$. If $\mathfrak{a}$ is an infinite cardinal number, then $|T|<\mathfrak{a}$ if and only if $|T/I|<\mathfrak{a}$.
 Proof.
-It is clear that, if $|T|<\mathfrak{a}$, then $|T/I|<\mathfrak{a}$ for each ideal $I$ of $T$.
-Conversely, assume that $I$ be a nilpotent ideal of $T$ and $|T/I|<\mathfrak{a}$.
-Since $I$ is a nilpotent ideal of $T$, we infer that there exists a natural number $n$ such that $I^{m}=0$.
-Now consider the chain $0=I^{m}\subseteq I^{m-1}\subseteq\cdots\subseteq I\subseteq T$.
-Since $T$ is a left Noetherian ring, we conclude that for each $i$, $I^{i}$ is a finitely generated left ideal.
-Hence for each $i$, the left $T$-module $I^{i}/I^{i+1}$, $0\leq i\leq m-1$ is finitely generated.
-Thus $|I^{i}/I^{i+1}|<\mathfrak{a}$.
+The implication $|T|<\mathfrak{a}\ \Longrightarrow|T/I|<\mathfrak{a}$ is obvious for every ideal $I$ of $T$.
+Conversely, assume $I$ is a nilpotent ideal of $T$ and $|T/I|<\mathfrak{a}$.
+Since $I$ is nilpotent, there exists a natural number $n$ with $I^{m}=0$.
+Consider the chain $0=I^{m}\subseteq I^{m-1}\subseteq\cdots\subseteq I\subseteq T$.
+Because $T$ is left Noetherian, each $I^{i}$ is a finitely generated left ideal.
+Consequently, for each $i$, $0\leq i\leq m-1$, the left $T$-module $I^{i}/I^{i+1}$ is finitely generated, and therefore $|I^{i}/I^{i+1}|<\mathfrak{a}$.
 In particular, $|I^{m-1}|<\mathfrak{a}$.
-Therefore for each $i$ we deduce that $|I^{i}|<\mathfrak{a}$ and thus $|T|<\mathfrak{a}$.
+By descending induction we obtain $|I^{i}|<\mathfrak{a}$, for all $i$, and finally $|T|<\mathfrak{a}$.
 ∎
-Note that in the proof of the previous lemma we only use the fact that $I$ is a nilpotent ideal of $T$ such that $I^{i}$ is a finitely generated left ideal of $T$, for each $i$.
-It is clear that by the condition of the previous lemma we conclude that $T$ is finite (resp.
-countable) if and only if $T/I$ is finite (resp.
+Note that in the proof of the previous lemma we only used that $I$ is a nilpotent ideal whose powers $I^{i}$ are finitely generated left ideals of $T$.
+Under the hypotheses of the lemma, we immediately conclude that $T$ is finite (resp.
+countable) precisely when $T/I$ is finite (resp.
 countable).
-We conclude this section by a result about the structure of maximal subrings of a certain direct product of rings.
-If $A$ and $B$ be rings, then we say that $A$ and $B$ are homomorphically non-isomorphic rings, whenever for each proper ideals $I$ of $A$ and $J$ of $B$, the rings $A/I$ and $B/J$ are not isomorphic as ring.
-For example if $A=\mathbb{M}_{n_{1}}(R_{1})\times\cdots\times\mathbb{M}_{n_{k}}(R_{k})$, where $k\geq 1$, each $n_{i}\geq 2$ and $R_{i}$ are arbitrary rings and $B$ is any commutative ring, then $A$ and $B$ are homomorphically non-isomorphic rings (note that clearly, each homomorphic image of $A$ is a non-commutative ring).
-Now the following is in order.
+We close this section with a result concerning the structure of maximal subrings of certain direct products of rings.
+Two rings $A$ and $B$ are called homomorphically non-isomorphic, if for every proper ideal $I$ of $A$ and every proper ideal $J$ of $B$, the rings $A/I$ and $B/J$ are not isomorphic.
+For example, if $A=\mathbb{M}_{n_{1}}(R_{1})\times\cdots\times\mathbb{M}_{n_{k}}(R_{k})$, where $k\geq 1$, each $n_{i}\geq 2$ and $R_{i}$ are arbitrary rings and $B$ is any commutative ring, then $A$ and $B$ are homomorphically non-isomorphic (since every homomorphic image of $A$ is noncommutative).
+With this preparation we state the following.
 Lemma 2.9.
@@
 Proof.
-It is clear that if $R$ is of the forms that is given in the statement of the lemma, then $R$ is a maximal subring of $T$.
+It is clear that if any subring of the forms described above is maximal in $T$.
 Conversely, let $R$ be a maximal subrings of $T$.
 If $0\times B\subseteq R$, then it is not hard to see that $R=A_{1}\times B$, where $A_{1}$ is a maximal subring of $A$.
-Similarly if $A\times 0\subseteq R$, then $R=A\times B_{1}$, where $B_{1}$ is a maximal subring of $B$.
-Hence assume that $R$ is a maximal subring of $T$ which does not contain $A\times 0$ and $0\times B$.
-Let $M=(A\times 0)\cap R=\{(a,0)\ |\ a\in A_{1}\}=A_{1}\times 0$, where $A_{1}$ is a proper subset of $A$ and $N=(0\times B)\cap R=\{(0,b)\ |\ b\in B_{1}\}=0\times B_{1}$, where $B_{1}$ is a proper subset of $B$.
+Similarly, if $A\times 0\subseteq R$, then $R=A\times B_{1}$, where $B_{1}$ is a maximal subring of $B$.
+Hence assume that $R$ contains neither $A\times 0$ nor $0\times B$.
+Let $M:=(A\times 0)\cap R=\{(a,0)\ |\ a\in A_{1}\}=A_{1}\times 0$, where $A_{1}$ is a proper subset of $A$ and $N:=(0\times B)\cap R=\{(0,b)\ |\ b\in B_{1}\}=0\times B_{1}$, where $B_{1}$ is a proper subset of $B$.
 Clearly $M$ and $N$ are ideals of $R$ and $M+N=A_{1}\times B_{1}$.
-Also note that since $A\times 0\nsubseteq R$ and $R$ is a maximal subring of $T$, we have $R+(A\times 0)=T$ and therefore $R/M\cong B$ and similarly $R/N\cong A$.
-Now we have two cases (note $(1,1)\in R$): $(i)$ $(1,0)\in R$ which equivalent to $(0,1)\in R$.
-Therefore in this case $M+N=R$ and thus $R=A_{1}\times B_{1}$.
+Also note that since $A\times 0\nsubseteq R$ and $R$ is a maximal subring of $T$, we have $R+(A\times 0)=T$; Consequently, $R/M\cong B$.
+Similarly, $R/N\cong A$.
+Now we distinguish two cases (recall that $(1,1)\in R$): $(\bf i)$ $(1,0)\in R$ (equivalently $(0,1)\in R$).
+Then $M+N=R$, so $R=A_{1}\times B_{1}$.
 Hence $A_{1}$ and $B_{1}$ are proper subrings of $A$ and $B$, respectively.
-Clearly $R=A_{1}\times B_{1}\subsetneq A\times B_{1}\subsetneq T$, which contradicts to the maximality of $R$.
-$(ii)$ hence assume that $(1,0)$ and $(0,1)$ are not in $R$.
-Thus $(1,1)\notin M+N$ and therefore $M+N$ is a proper ideal of $R$ which contains $M$ and $N$.
-Hence, by $R/M\cong B$ and $R/N\cong A$, we deduce that $A$ and $B$ have epimorphic images that are isomorphic as ring which is impossible.
-Thus we are done.
+But then $R=A_{1}\times B_{1}\subsetneq A\times B_{1}\subsetneq T$, contradicting maximality of $R$.
+$({\bf ii})$ $(1,0)\notin R$ and $(0,1)\notin R$.
+Thus $(1,1)\notin M+N$, so $M+N$ is a proper ideal of $R$ contains $M$ and $N$.
+From $R/M\cong B$ and $R/N\cong A$, we obtain that $A$ and $B$ have isomorphic homomorphic images, which is impossible, because $A$ and $B$ are homomorphically non-isomorphic.
+Since both cases lead to a contradiction, our assumption that $R$ contains neither $A\times 0$ nor $0\times B$ cannot hold.
+Thus $R$ must be one of the two forms stated in the lemma.
 ∎
 3. Rings with Infinitely Many Maximal Subrings
-In this section we study the number of maximal subrings for general non-commutative rings, especially for division rings, simple rings and also $K$-algebras over an infinite field $K$.
-If $E$ is a field with prime subfield $F$, then by [10, Corollary 1.5], whenever $E$ has only finitely many maximal subrings, then $F\cong\mathbb{Z}_{p}$ for some prime number $p$ and $E$ is algebraic over $F$ (for characterization of such fields see [2]).
-Moreover, if $E$ is not an absolutely algebraic field, then $|RgMax(E)|\geq max\{tr.deg(E/F),|E|\}$, by [10, Corollary 1.5].
-For division rings we have the next result.
+In this section we study the number of maximal subrings for general noncommutative rings, with particular emphasises on division rings, simple rings and $K$-algebras over an infinite field $K$.
+Let $E$ be a field with prime subfield $F$.
+By [10, Corollary 1.5], if $E$ has only finitely many maximal subrings, then $F\cong\mathbb{Z}_{p}$ for some prime number $p$ and $E$ is algebraic over $F$ (for a full characterization of such fields see [2]).
+Moreover, if $E$ is not an absolutely algebraic field, then $|\operatorname{RgMax}(E)|\geq\max\{\operatorname{tr.deg}(E/F),|E|\}$, by [10, Corollary 1.5].
+For division rings we obtain the following result.
 Proposition 3.1.
-Let $D$ be a non-commutative division ring.
+Let $D$ be a noncommutative division ring.
 If $D$ has a maximal subring $R$ which is a division ring, then $D$ has infinitely many maximal subrings.
@@
 Proof.
-Assume that $R$ be a maximal subring of $D$ which is a division ring.
+Assume that $R$ is a maximal subring of $D$ which is a division ring.
 Then by [17, Theorem 3], there exist infinitely many subrings of the form $xRx^{-1}$, where $x\in D\setminus\{0\}$.
@@
 For the next part, assume that $D$ has only finitely many maximal subrings, then by the first part we conclude that each maximal subring of $D$ is not a division ring.
-Now if $x$ is a non-centeral element of $D$ which is algebraic over the center of $D$, then by [5, Corollary 2.10], $D$ has a maximal subring which is a division ring which is impossible.
+Now, if $x$ is a non-centeral element of $D$ which is algebraic over the center of $D$, then by [5, Corollary 2.10], $D$ has a maximal subring which is a division ring which is impossible.
 Hence the center of $D$ is algebraically closed in $D$.
-For the final part note that if $R$ is a maximal subring of $D$ and $\alpha\in C_{D}(R)$ is not a central element of $D$, then clearly $R=C_{D}(\alpha)$, i.e., $R$ is a division ring which is impossible too.
+For the final part note that if $R$ is a maximal subring of $D$ and $\alpha\in C_{D}(R)$ is not a central element of $D$, then clearly $R=C_{D}(\alpha)$, i.e., $R$ is a division ring which is absurd.
 Thus $C_{D}(R)=C_{D}(D)$.
@@
 Let $x\in D\setminus F$ and $[F(x):F]<|F|$.
-Then similar to the proof of [21, Theorem 4.20], we deduce that $F(x)$ is algebraic over $F$.
+Then analogous to the proof of [21, Theorem 4.20], we deduce that $F(x)$ is algebraic over $F$.
 Hence $x$ is algebraic over $F$ and thus we are done by the previous result.
@@
 Corollary 3.3.
-Let $D$ be an existentially complete non-commutative division ring over a field $K$. Then $D$ has infinitely many maximal subrings.
+Let $D$ be an existentially complete noncommutative division ring over a field $K$. Then $D$ has infinitely many maximal subrings.
 Proof.
 By [5, $(1)$ of Proposition 3.5], $D$ has a maximal subring of the form $R=C_{D}(x)$ for some $x\in D$. It is clear that $R$ is a division ring and hence we are done by Proposition 3.1. ∎
-For more results first we need to characterize the maximal subrings of $R\times R$ for a simple ring $R$, as follow.
+To obtain further results, we first need to characterize the maximal subrings of $R\times R$ where $R$ is a simple ring. This done as follows.
 Theorem 3.4.
-Let $R$ be a simple ring. Then $S$ is a maximal subring of $T:=R\times R$ if and only if $S$ satisfies exactly one of the following conditions:
+Let $R$ be a simple ring. Then $S$ is a maximal subring of $T:=R\times R$ if and only if $S$ satisfies exactly in one of the following conditions:
   1. (1)
@@
   2. (2)
-    There exist $\sigma_{1},\sigma_{2}\in Aut(R)$ such that $S=\{(\sigma_{1}(x),\sigma_{2}(x))\ |\ x\in R\}$.
+    There exist $\sigma_{1},\sigma_{2}\in\operatorname{Aut}(R)$ such that $S=\{(\sigma_{1}(x),\sigma_{2}(x))\ |\ x\in R\}$.
 Moreover in case $(2)$, we have $S\cong R$.
@@
 To show $S$ is a maximal subring of $T$, we first prove that $S[(1,0)]=T$.
-Clearly $(0,1)\in S[(1,0)]$, for $(1,1)\in S$.
-Now let $(a,b)\in T=R\times R$, thus there exist $x,y\in R$ such that $\sigma_{1}(x)=a$ and $\sigma_{2}(y)=b$, for $\sigma_{1},\sigma_{2}\in Aut(R)$.
+Clearly $(0,1)\in S[(1,0)]$, because $(1,1)\in S$.
+Now let $(a,b)\in T=R\times R$, thus there exist $x,y\in R$ such that $\sigma_{1}(x)=a$ and $\sigma_{2}(y)=b$, because $\sigma_{1},\sigma_{2}\in\operatorname{Aut}(R)$.
 Thus by definition of $S$ we conclude that $(\sigma_{1}(x),\sigma_{2}(x))=(a,\sigma_{2}(x))\in S$ and $(\sigma_{1}(y),\sigma_{2}(y))=(\sigma_{1}(y),b)\in S$.
@@
 Now we prove that for each $(x,y)\in T\setminus S$, we have $S[(x,y)]=T$, which shows that $S$ is a maximal subring of $T$.
-Since $\sigma_{2}\in Aut(R)$, there exists $c\in R$ such that $\sigma_{2}(c)=y$.
-Therefore $(\sigma_{1}(c),y)\in S\subseteq S[(x,y)]$ and therefore $(x-\sigma_{1}(c),0)\in S[(x,y)]$.
+Since $\sigma_{2}\in\operatorname{Aut}(R)$, there exists $c\in R$ such that $\sigma_{2}(c)=y$.
+This implies that $(\sigma_{1}(c),y)\in S\subseteq S[(x,y)]$ and a fortiori $(x-\sigma_{1}(c),0)\in S[(x,y)]$.
 Since $(x,y)\notin S$, we infer that $x_{0}:=x-\sigma_{1}(c)\neq 0$.
-Thus $Rx_{0}R=R$, for $R$ is a simple ring.
+Thus $Rx_{0}R=R$, because $R$ is a simple ring.
 Hence there exist $n\in\mathbb{N}$ and $r_{i},s_{i}\in R$, $1\leq i\leq n$, such that $\sum_{i=1}^{n}r_{i}x_{0}s_{i}=1$.
-From $\sigma_{1}\in Aut(R)$, we deduce that there exist $a_{i},b_{i}\in R$, $1\leq i\leq n$, such that $\sigma_{1}(a_{i})=r_{i}$ and $\sigma_{1}(b_{i})=s_{i}$.
-Therefore
+From $\sigma_{1}\in\operatorname{Aut}(R)$, we deduce that there exist $a_{i},b_{i}\in R$, $1\leq i\leq n$, such that $\sigma_{1}(a_{i})=r_{i}$ and $\sigma_{1}(b_{i})=s_{i}$.
+It follows that
 $(1,0)=\sum_{i=1}^{n}(\sigma_{1}(a_{i}),\sigma_{2}(a_{i}))(x_{0},0)(\sigma_{1}(b_{i}),\sigma_{2}(b_{i}))\in S[(x,y)]$
-Thus $T=S[(1,0)]\subseteq S[(x,y)]$ and therefore $S[(x,y)]=T$. Hence $S$ is a maximal subring of $T$.
-Conversely, assume that $S$ be a maximal subring of $T$.
+This implies that $T=S[(1,0)]\subseteq S[(x,y)]$ and then $S[(x,y)]=T$. Consequently, $S$ is a maximal subring of $T$.
+Conversely, assume that $S$ is a maximal subring of $T$.
 Suppose $I=R\times 0$ and $J=0\times R$.
@@
 Also note that since $S$ is a maximal subring of $T$ which does not contain $I$ and $J$, we deduce that $S+I=T=S+J$ and thus $S/I_{1}\cong R\cong S/J_{1}$.
-Thus $I_{1}$ and $J_{1}$ are maximal ideals of $S$.
-Now we have two cases: $(a)$ if $(1,0)\in S$ (which is equivalent to $(0,1)\in S$), then $(1,1)=(1,0)+(0,1)\in I_{1}+J_{1}$, which means $I_{1}+J_{1}=S$ and therefore $S=S_{1}\times S_{2}$.
+Consequently, $I_{1}$ and $J_{1}$ are maximal ideals of $S$.
+Now we have two cases: $(\bf a)$ If $(1,0)\in S$ (which is equivalent to $(0,1)\in S$), then $(1,1)=(1,0)+(0,1)\in I_{1}+J_{1}$, which means $I_{1}+J_{1}=S$ and therefore $S=S_{1}\times S_{2}$.
 Thus $S_{1}$ and $S_{2}$ are proper subring of $R$.
-But clearly $S\subsetneq R\times S_{2}\neq T$, which contradicts the maximality of $S$.
-$(b)$ Hence $(1,0),(1,0)\notin S$.
-Thus $I_{1}+J_{1}=S_{1}\times S_{2}$ is a proper ideal of $S$ which contains $I_{1}$ and $J_{1}$.
+But obviously $S\subsetneq R\times S_{2}\neq T$, which contradicts to the maximality of $S$.
+$(\bf b)$ Hence we may suppose that $(1,0),(1,0)\notin S$.
+It implies that $I_{1}+J_{1}=S_{1}\times S_{2}$ is a proper ideal of $S$ which contains $I_{1}$ and $J_{1}$.
 But as we see earlier, $I_{1}$ and $J_{1}$ are maximal ideals of $S$.
-Therefore we immediately conclude that $I_{1}=J_{1}$.
-Thus $I_{1}=J_{1}=0$ and therefore $S\cong R$.
+Therefore we obtain that $I_{1}=J_{1}$.
+So $I_{1}=J_{1}=0$ and then $S\cong R$.
 Let $\sigma:R\longrightarrow S$ be a ring isomorphism, and for each $x\in R$, let $\sigma(x)=(\sigma_{1}(x),\sigma_{2}(x))$.
 Then clearly $\sigma_{1}$ and $\sigma_{2}$ are ring endomorphisms of $R$.
-Since $R$ is a simple ring we conclude that $\sigma_{1}$ and $\sigma_{2}$ are one-one.
-Finally note that since $S=\{(\sigma_{1}(x),\sigma_{2}(x))\ |\ x\in R\}$ and $(1,0)\notin S$, we deduce that $S[(1,0)]=T$.
-This immediately implies that $T=S+S(1,0)$ (note, $(1,0)$ is a central idempotent of $T$).
+Since $R$ is a simple ring we deduce that $\sigma_{1}$ and $\sigma_{2}$ are one-one.
+Finally note that from $S=\{(\sigma_{1}(x),\sigma_{2}(x))\ |\ x\in R\}$ and $(1,0)\notin S$, we have $S[(1,0)]=T$.
+It follows that $T=S+S(1,0)$ (note, $(1,0)$ is a central idempotent of $T$).
 To show that $\sigma_{1}$ and $\sigma_{2}$ are onto, let $x\in R$, then $(x,0)\in S+S(1,0)$.
-Therefore there exist $a,b\in R$ such that $(x,0)=(\sigma_{1}(a),\sigma_{2}(a))+(\sigma_{1}(b),\sigma_{2}(b))(1,0)$, which immediately shows that $\sigma_{1}(b)=x$ and therefore $\sigma_{1}$ is onto.
-Similarly, $\sigma_{2}$ is onto and hence we are done.
+Consequently there exist $a,b\in R$ such that $(x,0)=(\sigma_{1}(a),\sigma_{2}(a))+(\sigma_{1}(b),\sigma_{2}(b))(1,0)$, which shows that $\sigma_{1}(b)=x$, i.e., $\sigma_{1}$ is onto.
+Analogously, $\sigma_{2}$ is onto, and hence we are done.
 ∎
-Next we have the following main results in this paper for simple rings.
+We now present a main result of this paper concerning simple rings.
 Theorem 3.5.
 Let $T$ be a simple ring which is not a division ring.
-Then $T$ has only finitely many maximal subring if and only if $R$ is finite (i.e., $T\cong\mathbb{M}_{n}(F)$, where $F$ is a finite field and $n\geq 2$)
+Then $T$ has only finitely many maximal subring if and only if $T$ is finite (i.e., $T\cong\mathbb{M}_{n}(F)$, where $F$ is a finite field and $n\geq 2$)
 Proof.
-First we show that $Max_{l}(T)$ is finite.
-For otherwise let $M_{1},M_{2},\ldots$ be distinct maximal left ideals of $T$.
-Then for each $i$, by $(1)$ of Theorem 2.7, $A_{i}=\mathbb{I}_{T}(M_{i})$ is a maximal subring of $T$ which contains $M_{i}$ (note, for each $i$, $M_{i}$ is not an ideal of $T$).
-Since for $i\neq j$, $M_{i}+M_{j}=T$, we immediately conclude that $A_{i}\neq A_{j}$.
+We first show that $\operatorname{{Max}_{\ell}}(T)$ is finite.
+Otherwise, let $M_{1},M_{2},\ldots$ be distinct maximal left ideals of $T$.
+Then for each $i$, by Theorem 2.7(1)(1), $A_{i}:=\mathbb{I}_{T}(M_{i})$ is a maximal subring of $T$ which contains $M_{i}$ (note, for each $i$, $M_{i}$ is not an ideal of $T$, because $T$ is simple ring which is not a division ring).
+Since whenever $i\neq j$, we have $M_{i}+M_{j}=T$, then we have $A_{i}\neq A_{j}$.
 Thus $T$ has infinitely many maximal subrings which is absurd.
-Thus $Max_{l}(T)=\{M_{1},\ldots,M_{n}\}$.
-Now since $J(T)=0$, we deduce that $M_{1}\cap\cdots\cap M_{n}=0$.
-Thus $T$ can be embedded in $T/M_{1}\times\cdots\times T/M_{n}$.
-Hence $T$ is a left Artinian ring.
-Therefore $T\cong\mathbb{M}_{n}(D)$ for some division ring $D$ and $n\geq 2$ (note, $T$ is not a division ring).
+Thus $\operatorname{{Max}_{\ell}}(T)=\{M_{1},\ldots,M_{n}\}$.
+Now from the fact that $J(T)=0$, we deduce that $M_{1}\cap\cdots\cap M_{n}=0$.
+Consequently, $T$ embeds in $T/M_{1}\times\cdots\times T/M_{n}$.
+A fortiori, $T$ is a left Artinian ring.
+It follows that $T\cong\mathbb{M}_{n}(D)$, for some division ring $D$ and $n\geq 2$ (note, $T$ is not a division ring).
 If $D$ is an infinite division ring, then $\mathbb{M}_{n}(D)$, where $n\geq 2$, has infinitely many maximal left (right) ideals, by Corollary 2.4.
-Thus $D$ is finite and therefore $D$ is a finite field.
-Hence $T$ is a finite ring.
+Thus $D$ is finite and then $D$ is a finite field.
+This implies that $T$ is a finite ring, which complete the proof.
 ∎
-We remind the reader that if $K$ is a field, then $K\times K$ has only finitely many maximal subrings if and only if $K$ is finite, see [10, Corollary 3.5].
-Now the following generalization is in order.
+We recall that for a field $K$, the ring $K\times K$ has only finitely many maximal subrings if and only if $K$ is finite, see [10, Corollary 3.5]. The following result generalizes this fact.
 Proposition 3.6.
-Let $R$ be a simple ring. Then $R\times R$ has finitely many maximal subrings if and only if $R$ is finite.
+Let $R$ be a simple ring. Then $R\times R$ has only finitely many maximal subrings if and only if $R$ is finite.
 Proof.
-It is clear that if $R$ is finite then $R\times R$ has only finitely many maximal subrings.
+It is trivial that if $R$ is finite then $R\times R$ has only finitely many maximal subrings.
 Conversely, assume that $R\times R$ has only finitely many maximal subrings.
-Thus $R$ has only finitely many maximal subrings.
+Theorem 3.4 implies that $R$ has only finitely many maximal subrings.
 If $R$ is not a division ring, then we are done by the previous theorem.
-Hence assume that $R$ is a division ring and $R\times R$ has only finitely many maximal subrings.
-By the preceding comment we may assume that $R$ is not a field and therefore $C_{R}(R)\subsetneq R$.
-Now suppose that $R$ is infinite.
-Then we have two cases: $(a)$ If there exists a non-central element $x$ of $R$ which is algebraic over the center of $R$, then by Proposition 3.1, $R$ has infinitely many maximal subrings.
-Thus $R\times R$ has infinitely many maximal subring too, which is impossible.
-$(b)$ Hence assume that each non-central element $x$ of $R$ is not algebraic over the center of $R$.
-For each $n\geq 1$ and $a\in D$, let $\sigma_{n}(a)=x^{n}ax^{-n}$ be the inner automorphism of $R$ induced from $x^{n}$.
-If $\sigma_{n}=\sigma_{m}$ for $n>m$, then $\sigma_{1}^{n-m}=1$ and therefore $x^{n-m}$ is a central element of $R$.
-This immediately implies that $x$ is integral over the center of $R$, which contradict to the assumption in this case.
+Thus suppose that $R$ is a division ring and $R\times R$ has only finitely many maximal subrings.
+By the preceding comment, we may assume that $R$ is not a field, i.e., $C_{R}(R)\subsetneq R$.
+Now, let $R$ be an infinite ring.
+Then we have two cases: $(\bf a)$ If there exists a non-central element $x$ of $R$ which is algebraic over the center of $R$, then by Proposition 3.1, $R$ has infinitely many maximal subrings, which is impossible.
+$(\bf b)$ Hence assume that each non-central element $x$ of $R$ is not algebraic over the center of $R$.
+Let $x\in R$ be a non-central element and for each $n\geq 1$ and $a\in D$, let $\sigma_{n}(a)=x^{n}ax^{-n}$ be the inner automorphism of $R$ induced from $x^{n}$.
+If $\sigma_{n}=\sigma_{m}$ for some $n>m$, then $\sigma_{1}^{n-m}=1$ and therefore $x^{n-m}$ is a central element of $R$.
+This implies that $x$ is integral over the center of $R$, which contradicts to the assumption that $x$ is not algebraic over the center.
 Thus $\sigma_{n}\neq\sigma_{m}$ for each $n\neq m$.
-Finally note that by Theorem 3.4, for each $n\geq 1$, the set $S_{n}=\{(x,\sigma_{n}(x))\ |\ x\in R\}$ is a maximal subring of $R\times R$ and clearly for $n\neq m$, we have $S_{n}\neq S_{m}$.
-Therefore $R\times R$ has infinitely many maximal subrings, which is absurd.
-Thus $R$ is a finite and hence we are done.
+Consequently, by Theorem 3.4(2)(2), for each $n\geq 1$, the set $S_{n}=\{(x,\sigma_{n}(x))\ |\ x\in R\}$ is a maximal subring of $R\times R$.
+Obviously, whenever $n\neq m$, we have $S_{n}\neq S_{m}$.
+It follows that $R\times R$ has infinitely many maximal subrings, which is absurd.
+Thus $R$ is a finite, as desired.
 ∎
@@
 By [12, Theorem 6.5.11], $D$ has at least $2^{\aleph_{0}}$ automorphisms.
-Hence by Theorem 3.4 for each $\sigma\in Aut(D)$, $S_{\sigma}:=\{(x,\sigma(x))\ |\ x\in D\}$ is a maximal subring of $D\times D$.
+Hence by Theorem 3.4 for each $\sigma\in\operatorname{Aut}(D)$, $S_{\sigma}:=\{(x,\sigma(x))\ |\ x\in D\}$ is a maximal subring of $D\times D$.
 It is clear that $S_{\sigma}\neq S_{\sigma^{\prime}}$, whenever $\sigma$ and $\sigma^{\prime}$ are two distinct automorphisms of $D$.
-Hence $D\times D$ has at least $2^{\aleph_{0}}$ maximal subrings and hence we are done.
+Hence $D\times D$ has at least $2^{\aleph_{0}}$ maximal subrings.
 ∎
@@
 Conversely, assume that $T$ has infinitely many maximal subrings, we prove that $(1)$ or $(2)$ holds.
-Suppose that $(2)$ does not hold, i.e., $R_{1}$ and $R_{2}$ are not isomorphic rings.
+Suppose that $(2)$ does not hold, i.e., $R_{1}$ and $R_{2}$ are not isomorphic as rings.
 Thus $R_{1}$ and $R_{2}$ are homomorphically non-isomorphic rings.
-Therefore by Lemma 2.9, we infer that each maximal subrings of $T$ is of the form $S_{1}\times R_{2}$ or $R_{1}\times S_{2}$, where $S_{i}$ is a maximal subring of $R_{i}$.
-Since $T$ has infinitely many maximal subrings, we immediately conclude that either $R_{1}$ or $R_{2}$ has infinitely many maximal subrings and hence we are done.
+Consequently, by Lemma 2.9, we infer that each maximal subrings of $T$ is of the form $S_{1}\times R_{2}$ or $R_{1}\times S_{2}$, where $S_{i}$ is a maximal subring of $R_{i}$.
+Since $T$ has infinitely many maximal subrings, we obtain that either $R_{1}$ or $R_{2}$ has infinitely many maximal subrings and hence we are done.
 ∎
-For a ring $T$, let us denote the sets $Max_{l}(T)\setminus Max(T)$ by $Max_{l}^{\prime}(T)$, similarly $Max_{r}^{\prime}(T):=Max_{r}(T)\setminus Max(T)$ and $Max^{\prime}(T):=Max(T)\setminus Max_{l}(T)=Max(T)\setminus Max_{r}(T)$.
-Note that if $P$ is a prime ideal of a ring $T$ and $A$ be a left (resp.
-right) ideal and $J$ be an ideal of $T$ such that $P=I\cap J$, then either $P=I$ or $P=J$.
-To see this note that clearly $JI\subseteq I\cap J\subseteq P$ (resp.
+For a ring $T$, let us denote the sets $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$ by $\operatorname{{Max}_{\ell}}^{\prime}(T)$, similarly $\operatorname{{Max}_{r}}^{\prime}(T):=\operatorname{{Max}_{r}}(T)\setminus\operatorname{Max}(T)$ and $\operatorname{Max}^{\prime}(T):=\operatorname{Max}(T)\setminus\operatorname{{Max}_{\ell}}(T)=\operatorname{Max}(T)\setminus\operatorname{{Max}_{r}}(T)$.
+Note that if $P$ is a prime ideal of a ring $T$ and $I$ be a left (resp.
+right) ideal and $J$ is an ideal of $T$ such that $P=I\cap J$, then either $P=I$ or $P=J$.
+To see this, note that $JI\subseteq I\cap J\subseteq P$ (resp.
 $IJ\subseteq I\cap J\subseteq P$), therefore $I\subseteq P$ or $J\subseteq P$.
@@
   1. (1)
-    $Max_{l}^{\prime}(T)$ and $Max_{r}^{\prime}(T)$ are finite sets and in fact $|Max_{l}^{\prime}(T)|\leq|RgMax(T)|$ and $|Max_{r}^{\prime}(T)|\leq|RgMax(T)|$.
+    $\operatorname{{Max}_{\ell}}^{\prime}(T)$ and $\operatorname{{Max}_{r}}^{\prime}(T)$ are finite sets and in fact $|\operatorname{{Max}_{\ell}}^{\prime}(T)|\leq|\operatorname{RgMax}(T)|$ and $|\operatorname{{Max}_{r}}^{\prime}(T)|\leq|\operatorname{RgMax}(T)|$.
   2. (2)
-    for each $M\in Max_{l}^{\prime}(T)\cup Max_{r}^{\prime}(T)$, the ring $\mathbb{I}(M)/M$ is a finite field. In particular, $C(T)/(C(T)\cap M)$ is a finite field.
+    For each $M\in\operatorname{{Max}_{\ell}}^{\prime}(T)\cup\operatorname{{Max}_{r}}^{\prime}(T)$, the ring $\mathbb{I}(M)/M$ is a finite field. In particular, $C(T)/(C(T)\cap M)$ is a finite field.
   3. (3)
-    for each $M\in Max_{l}(T)\cap Max(T)=Max_{r}(T)\cap Max(T)$, the center of $T/M$ is algebraically closed in $T/M$.
+    For each $M\in\operatorname{{Max}_{\ell}}(T)\cap\operatorname{Max}(T)=\operatorname{{Max}_{r}}(T)\cap\operatorname{Max}(T)$, the center of $T/M$ is algebraically closed in $T/M$.
   4. (4)
-    for each $M\in Max^{\prime}(T)$, there exist a natural number $n>1$ and a finite field $F$ such that $T/M\cong\mathbb{M}_{n}(F)$.
+    For each $M\in\operatorname{Max}^{\prime}(T)$, there exist a natural number $n>1$ and a finite field $F$ such that $T/M\cong\mathbb{M}_{n}(F)$.
   5. (5)
-    $Max^{\prime}(T)$ is finite and in fact $|Max^{\prime}(T)|\leq|RgMax(T)|$.
+    $\operatorname{Max}^{\prime}(T)$ is finite and in fact $|\operatorname{Max}^{\prime}(T)|\leq|\operatorname{RgMax}(T)|$.
   6. (6)
-    $Prm_{l}(T)\setminus Max_{l}(T)$, $Prm_{l}(T)\setminus Max(T)$, $Prm_{r}(T)\setminus Max_{r}(T)$ and $Prm_{r}(T)\setminus Max(T)$ are finite sets.
+    $\operatorname{{Prm}_{\ell}}(T)\setminus\operatorname{{Max}_{\ell}}(T)$, $\operatorname{{Prm}_{\ell}}(T)\setminus\operatorname{Max}(T)$, $\operatorname{{Prm}_{r}}(T)\setminus\operatorname{{Max}_{r}}(T)$ and $\operatorname{{Prm}_{r}}(T)\setminus\operatorname{Max}(T)$ are finite sets.
   7. (7)
     If $P$ a left (resp.
-right) primitive ideal of $T$, then either $P$ is a maximal ideal of $T$ or $P=\bigcap_{i\in I}M_{i}$, where $M_{i}\in Max_{l}(T)\cap Max(T)$, moreover in this case $I$ is infinite and $T/P$ is a domain (i.e., $P$ is a completely prime ideal of $T$).
+right) primitive ideal of $T$, then either $P$ is a maximal ideal of $T$ or $P=\bigcap_{i\in I}M_{i}$, where $M_{i}\in\operatorname{{Max}_{\ell}}(T)\cap\operatorname{Max}(T)$, moreover in this case $I$ is infinite and $T/P$ is a domain (i.e., $P$ is a completely prime ideal of $T$).
 In particular, $T/P$ is embeddable in a division ring.
   8. (8)
-    $J(T)=\bigcap_{M\in Max(T)}M$.
+    $J(T)=\bigcap_{M\in\operatorname{Max}(T)}M$.
 Proof.
-$(1)$ for each maximal left ideal $A$ of $T$ which is not an ideal of $T$, by Theorem 2.7, $\mathbb{I}(A)$ is a maximal subring of $T$ which contains $A$.
-It is clear that whenever $A\neq B$ are in $Max_{l}^{\prime}(T)$, then $A+B=T$.
-Hence we conclude that $\mathbb{I}(A)\neq\mathbb{I}(B)$.
-Therefore the function $A\longmapsto\mathbb{I}(A)$, from $Max_{l}^{\prime}(T)$ in to $RgMax(T)$ is one-one and therefore $(1)$ holds.
-For $Max_{r}^{\prime}(T)$ the proof is similar.
-$(2)$ is an immediate consequence of Theorem 2.1 and $(1)$ (note that if $M\in Max_{l}^{\prime}(T)$, then $\mathbb{I}(M)/M\cong(End({}_{T}(T/M)))^{op}$ is a division ring).
-For the final part of $(2)$ note that $C(T)+M$ is a subring of $T$ which is contained in $\mathbb{I}(M)$ and therefore $(C(T)+M)/M$ is a subring of $\mathbb{I}(M)/M$.
-Now by the first part of $(2)$, we immediately deduce that $(C(T)+M)/M$ is a finite field and thus $C(T)/(C(T)\cap M)$ is a finite field too.
-$(3)$ if $T/M$ is a field (in particular, if $T/M$ is finite), then the conclusion is obvious.
-Hence assume that $T/M$ is an infinite division ring, then by Theorem 3.1, we deduce that the center of $T/M$ is algebraically closed in $T/M$ (for otherwise, $T/M$ and therefore $T$ have infinitely many maximal subrings, which is absurd).
-$(4)$ this is an immediate conclusion of Theorem 3.5 and the fact that $T/M$ has only finitely many maximal subrings (for $T$ has only finitely many maximal subrings).
-$(5)$ assume that $Max^{\prime}(T)=\{M_{i}\ |\ i\in I\}$, where $M_{i}\neq M_{j}$ for $i\neq j\in I$.
+$(\bf 1)$ For each maximal left ideal $A$ of $T$ which is not an ideal of $T$, by Theorem 2.7, $\mathbb{I}(A)$ is a maximal subring of $T$ which contains $A$.
+It is clear that whenever $A\neq B$ are in $\operatorname{{Max}_{\ell}}^{\prime}(T)$, then $A+B=T$ and consequently $\mathbb{I}(A)\neq\mathbb{I}(B)$.
+Therefore the function $A\longmapsto\mathbb{I}(A)$, from $\operatorname{{Max}_{\ell}}^{\prime}(T)$ into $\operatorname{RgMax}(T)$ is one-one, which complete the proof of $(1)$ for $\operatorname{{Max}_{\ell}}(T)$.
+The proof for $\operatorname{{Max}_{r}}^{\prime}(T)$ is analogous.
+$(\bf 2)$ The first part is an immediate consequence of Theorem 2.1 and $(1)$ (note that if $M\in\operatorname{{Max}_{\ell}}^{\prime}(T)$, then $\mathbb{I}(M)/M\cong(\operatorname{End}({}_{T}(T/M)))^{op}$ is a division ring).
+For the last part of $(2)$, obviously $C(T)+M$ is a subring of $T$ which is contained in $\mathbb{I}(M)$.
+It follows that $(C(T)+M)/M$ is a subring of $\mathbb{I}(M)/M$.
+Now by the first part of $(2)$, we deduce that $(C(T)+M)/M$ is a finite field.
+A fortiori, $C(T)/(C(T)\cap M)$ is a finite field.
+$(\bf 3)$ If $T/M$ is a field (in particular, if $T/M$ is finite), then the conclusion is obvious.
+Hence assume that $T/M$ is an infinite division ring.
+By Theorem 3.1, we deduce that the center of $T/M$ is algebraically closed in $T/M$, because otherwise, $T/M$ and therefore $T$ have infinitely many maximal subrings, which is absurd.
+$(\bf 4)$ This is an immediate consequence of Theorem 3.5 and the fact that $T/M$ has only finitely many maximal subrings (note that $T$ has only finitely many maximal subrings and thus $T/M$ has only finitely many maximal subrings).
+$(\bf 5)$ Suppose that $\operatorname{Max}^{\prime}(T)=\{M_{i}\ |\ i\in I\}$, where $M_{i}\neq M_{j}$ for $i\neq j\in I$.
 Then by $(4)$, for each $i\in I$ there exist a natural number $n_{i}>1$ and a finite field $F_{i}$ such that $T/M_{i}\cong\mathbb{M}_{n_{i}}(F_{i})$.
-Therefore by $(4)$ of Theorem 2.7, we infer that for each $i\in I$, $M_{i}$ is contained in a maximal subring $S_{i}$ of $T$.
+Therefore by Theorem 2.7(4)(4), we infer that for each $i\in I$, $M_{i}$ is contained in a maximal subring $S_{i}$ of $T$.
 Since for $i\neq j$, we have $M_{i}+M_{j}=T$, we deduce that $S_{i}\neq S_{j}$.
-Therefore $|I|\leq|RgMax(T)|$ and hence we are done.
-$(6)$ let $\{P_{i}\ |\ i\in I\}$ be a family of distinct left primitive ideals of $T$ such that for each $i\in I$, $P_{i}$ is not a maximal (left) ideal of $T$.
-Since for each $i\in I$, $P_{i}$ is a left primitive ideal of $T$, we deduce that there exists a maximal left ideal $M_{i}$ of $T$ such that $P_{i}=l.ann_{T}(T/M_{i})$.
+Therefore $|I|\leq|\operatorname{RgMax}(T)|$ and hence we are done.
+$(\bf 6)$ Let $\{P_{i}\ |\ i\in I\}$ be a family of distinct left primitive ideals of $T$ such that for any $i\in I$, $P_{i}$ is not a maximal (left) ideal of $T$.
+Since for each $i\in I$, $P_{i}$ is a left primitive ideal of $T$, we deduce that there exists a maximal left ideal $M_{i}$ of $T$ such that $P_{i}=\operatorname{l.ann}_{T}(T/M_{i})$.
 It is clear that whenever $i\neq j$, then $M_{i}\neq M_{j}$, for $P_{i}\neq P_{j}$.
 Also note that if $M_{i}$ is an ideal of $T$, then $P_{i}=M_{i}$ is a maximal (left) ideal of $T$ which is absurd.
-Thus for each $i\in I$ we deduce that $M_{i}\in Max_{l}^{\prime}(T)$ and therefore by $(1)$ we conclude that $I$ is finite.
-Thus the first and the second parts of $(6)$ hold.
-The proof for the next two parts are similar.
-$(7)$ let $P$ be a left primitive ideal of $T$.
+Thus for each $i\in I$ we obtain that $M_{i}\in\operatorname{{Max}_{\ell}}^{\prime}(T)$ and therefore by $(1)$ we conclude that $I$ is finite.
+It follows that $\operatorname{{Prm}_{\ell}}(T)\setminus\operatorname{{Max}_{\ell}}(T)$, $\operatorname{{Prm}_{\ell}}(T)\setminus\operatorname{Max}(T)$ are finite.
+The proof for $\operatorname{{Prm}_{r}}(T)\setminus\operatorname{{Max}_{r}}(T)$ and $\operatorname{{Prm}_{r}}(T)\setminus\operatorname{Max}(T)$ are analogous.
+$(\bf 7)$ Let $P$ be a left primitive ideal of $T$.
 Hence $P$ is an intersection of a family of maximal left ideals of $T$.
-Since by $(1)$, $Max_{l}^{\prime}(T)$ is finite, we conclude that there exist $N_{1},\ldots,N_{k}\in Max_{l}^{\prime}(T)$, $k\geq 0$, and a family $\{M_{i}\ |\ i\in I\}\subseteq Max_{l}(T)\cap Max(T)$ such that $P=N_{1}\cap\cdots N_{k}\cap\bigcap_{i\in I}M_{i}$.
+Since by $(1)$, $\operatorname{{Max}_{\ell}}^{\prime}(T)$ is finite, we conclude that there exist $N_{1},\ldots,N_{k}\in\operatorname{{Max}_{\ell}}^{\prime}(T)$, $k\geq 0$, and a family $\{M_{i}\ |\ i\in I\}\subseteq\operatorname{{Max}_{\ell}}(T)\cap\operatorname{Max}(T)$ such that $P=N_{1}\cap\cdots N_{k}\cap N$, where $N=\bigcap_{i\in I}M_{i}$.
 By the comments preceding the proposition we infer that either $P=N_{1}\cap\cdots N_{k}$ or $P=\bigcap_{i\in I}M_{i}$.
-If $P=N_{1}\cap\cdots N_{k}$, then the ring $T/P$ embeds in the semisimple left $T$-module $T/N_{1}\times\cdots\times T/N_{k}$, as left $T$-module and therefore $T/P$ is a prime semisimple ring.
-Thus $T/P$ is a simple ring, i.e., $P$ is a maximal ideal of $T$.
-Hence assume that $P$ is not a maximal ideal of $T$ and $P=\bigcap_{i\in I}M_{i}$.
-Since $P$ is not maximal ideal of $T$, but is a prime ideal of $T$, we immediately conclude that $I$ is infinite.
+If $P=N_{1}\cap\cdots N_{k}$, then the ring $T/P$ embeds in the semisimple left $T$-module $T/N_{1}\times\cdots\times T/N_{k}$, as left $T$-modules.
+This implies that $T/P$ is a prime semisimple ring.
+Consequently $T/P$ is a simple ring, i.e., $P$ is a maximal ideal of $T$.
+Hence we may suppose that $P$ is not a maximal ideal of $T$ and $P=N=\bigcap_{i\in I}M_{i}$.
+Since $P$ is not maximal ideal of $T$, but is a prime ideal of $T$, we conclude that $I$ is infinite.
 Finally note that for each $i\in I$, the ring $T/M_{i}$ is a division ring.
-Therefore $\prod_{i\in I}T/M_{i}$ is a reduced ring.
+A fortiori, $\prod_{i\in I}T/M_{i}$ is a reduced ring.
 Since $T/P$ embeds in $\prod_{i\in I}T/M_{i}$, we deduce that $T/P$ is a reduced ring.
-Thus $T/P$ is a reduced prime ring and hence is a domain.
-Therefore by [12, Theorem 1.2.4], $T/P$ is embeddable in a division ring.
-$(8)$ This is trivial by $(7)$ and the fact that $J(T)$ is the intersection of all left (resp. right) primitive ideals of $T$. ∎
+Thus $T/P$ is a reduced prime ring and consequently $T/P$ is a domain.
+It follows by [12, Theorem 1.2.4] that $T/P$ is embeddable in a division ring.
+$(\bf 8)$ This is trivial by $(7)$ and the fact that $J(T)$ is the intersection of all left (resp. right) primitive ideals of $T$. ∎
+The previous observations yield the following quick corollaries.
 Corollary 3.11.
@@
 Hence assume that $T$ is not a simple ring.
-If $T$ has only finitely many maximal subrings, then by $(7)$ of Proposition 3.10, we conclude that $T$ is embeddable in a division ring which is absurd.
+If $T$ has only finitely many maximal subrings, then by Proposition 3.10(7)(7), we conclude that $T$ is embeddable in a division ring which is absurd.
 Thus $T$ has infinitely many maximal subrings.
@@
 Proof.
-Assume that $T$ has only finitely many maximal subrings, then by $(8)$ and $(5)$ of Proposition 3.10, we conclude that $0=J(T)=Q_{1}\cap\cdots\cap Q_{n}\cap N$, where $Q_{i}$ are maximal ideals of $T$ which are not maximal left ideals, $n\geq 0$ and $N=\bigcap_{i\in I}M_{i}$, where $Max(T)\cap Max_{l}(T)=\{M_{i}\ |\ i\in I\}$.
+Assume that $T$ has only finitely many maximal subrings, then by $(8)$ and $(5)$ of Proposition 3.10, we conclude that $0=J(T)=Q_{1}\cap\cdots\cap Q_{n}\cap N$, where $Q_{i}$ are maximal ideals of $T$ which are not maximal left ideals, $n\geq 0$ and $N=\bigcap_{i\in I}M_{i}$, where $\operatorname{Max}(T)\cap\operatorname{{Max}_{\ell}}(T)=\{M_{i}\ |\ i\in I\}$.
 Since $T$ is a prime ring, we infer that either there exists $k$ such that $Q_{k}=0$, i.e., $T$ is an infinite simple ring, or $N=0$.
@@
 Thus assume that $N=0$.
-Therefore $T$ embeds in $\prod_{i\in I}T/M_{i}$, and similar to the proof of $(7)$, we conclude that $T$ is embeddable in a division ring which is absurd.
+Therefore $T$ embeds in $\prod_{i\in I}T/M_{i}$, and similar to the proof of Proposition 3.10(7)(7), we conclude that $T$ is embeddable in a division ring which is impossible.
 Thus $T$ has infinitely many maximal subrings.
@@
 Lemma 3.13.
-Let $T$ be an infinite ring and $\{M_{i}\ |\ i\in I\}\subseteq Max_{l}(T)\cap Max(T)$ such that $\bigcap_{i\in I}M_{i}=0$.
+Let $T$ be an infinite ring and $\bigcap_{i\in I}M_{i}=0$ where $\{M_{i}\ |\ i\in I\}\subseteq\operatorname{{Max}_{\ell}}(T)\cap\operatorname{Max}(T)$.
 Then either $T$ has infinitely many maximal subring or $C(T)$ is integrally closed in $T$.
-In particular, if $T$ is a (left) quasi duo $J$-semisimple ring, then either $T$ has infinitely many maximal subring or $C(T)$ is integrally closed in $T$.
+In particular, if $T$ is a (left) quasi-duo $J$-semisimple ring, then either $T$ has infinitely many maximal subring or $C(T)$ is integrally closed in $T$.
 Proof.
 Assume that $T$ has only finitely many maximal subrings.
-Hence by Theorem 3.1, for each $i\in I$, the center of $T/M_{i}$ is integrally closed in $T/M_{i}$, for otherwise if there exists $i\in I$ such that $T/M_{i}$ has a non-central element which is integral over the center of $T/M_{i}$, then $T/M_{i}$ and therefore $T$ have infinitely many maximal subrings, which is absurd.
-Now assume that $\alpha$ be a non-central element of $T$ which is integral over the center of $T$.
+Consequently for any $i\in I$, the division ring $T/M_{i}$ has only finitely many maximal subrings.
+Hence by Theorem 3.1, for each $i\in I$, the center of $T/M_{i}$ is integrally closed in $T/M_{i}$, for otherwise if there exists $i\in I$ such that $T/M_{i}$ has a non-central element which is integral over the center of $T/M_{i}$, then $T/M_{i}$ has infinitely many maximal subrings, which is absurd.
+Now assume that $\alpha$ is a non-central element of $T$ which is integral over the center of $T$.
 Thus there exists $\beta\in T$ such that $\alpha\beta-\beta\alpha\neq 0$.
@@
 Therefore $\alpha+M$ is a non-central element of the division ring $T/M_{i}$.
-Clearly, $\alpha+M_{i}$ is integral over the center of $T/M_{i}$, which is not possible.
-Thus we conclude that the center of $T$ is integrally closed in $T$.
-The final part is evident.
-∎
-Remark 3.14.
-Let $T$ be a ring with only finitely many maximal subrings.
-Then $Max_{l}(T)$ is finite if and only if $T/J(T)$ is finite if and only if $Max_{r}(T)$ is finite.
-To see this, it suffices to show that if $Max_{l}(T)$ is finite, then $T/J(T)$ is finite.
-Since $Max_{l}(T)$ is finite, we immediately conclude that $T/J(T)$ is a semisimple ring.
-By assumption, $T$ and therefore $T/J(T)$ has only finitely many maximal subrings, therefore by Proposition 3.6, we deduce that each simple part of $T/J(T)$ is finite and therefore $T/J(T)$ is a finite ring.
-For the next result, by the previous remark, if $T$ is a ring with only finitely many maximal subrings, then we may assume that $Max_{l}(T)$ is infinite and therefore by Proposition 3.10, we conclude that $Max_{l}(T)\cap Max(T)$ is infinite.
-Corollary 3.15.
-Let $T$ be an infinite ring with only finitely many maximal subrings. Then there exist finitely many distinct maximal ideals $Q_{1},\ldots,Q_{m}$, $m\geq 0$, such that:
-  1. (1)
-    for each $i$, there exist a natural number $n_{i}>1$ and a finite field $F_{i}$ such that $T/Q_{i}\cong\mathbb{M}_{n_{i}}(F_{i})$.
-  2. (2)
-    $J(T)=Q_{1}\cap\cdots\cap Q_{m}\cap N$, where $N=\bigcap_{M\in Max(T)\cap Max_{l}(T)}M$, where $N,Q_{1},\ldots,Q_{m}$ are incomparable.
-  3. (3)
-    $T/J(T)\cong R\times S$, where $R$ is a finite semisimple ring (in fact, $R=\prod_{i=1}^{n}\mathbb{M}_{n_{i}}(F_{i})$) and the center of $S$ is integrally closed in $S$.
-Proof.
-First note that by $(8)$ of Proposition 3.10, $J(T)=\bigcap_{M\in Max(T)}M$.
-Since by $(5)$ of Proposition 3.10, $Max^{\prime}(T)$ is finite we deduce that there exists finitely many distinct maximal ideals $Q_{1},\ldots,Q_{m}\in Max^{\prime}(T)$, $m\geq 0$, such that $J(T)=Q_{1}\cap\cdots\cap Q_{m}\cap N$, where $N=\bigcap_{M\in Max(T)\cap Max_{l}(T)}M$ and clearly we may assume that $N$ and $Q_{i}$ are incomparable.
-Thus $(2)$ holds and note that $(1)$ is obvious by $(4)$ of Proposition 3.10.
-Since $Q_{i}$ are distinct maximal ideals which are incomparable to $N$, and $J(T)=Q_{1}\cap\cdots\cap Q_{m}\cap N$, we deduce the first part of $(3)$ (where $S=T/N$).
-The final part of $(3)$ is a consequence of Lemma 3.13.
-Hence we are done.
+Clearly, $\alpha+M_{i}$ is integral over the center of $T/M_{i}$, which is impossible.
+Thus, the center of $T$ is integrally closed in $T$.
+The last part is evident.
 ∎
@@
 Theorem 3.14.
-Let $T$ be a $K$-algebra over an infinite field $K$. Then either $T$ has infinitely many maximal subrings (in fact, $|RgMax(T)|\geq|K|$) or the following hold:
+Let $T$ be a $K$-algebra over an infinite field $K$. Then either $T$ has infinitely many maximal subrings (in fact, $|\operatorname{RgMax}(T)|\geq|K|$) or the following hold:
   1. (1)
-    $T$ is a quasi duo ring. In particular, $Max(T)=Max_{l}(T)=Max_{r}(T)=Prm_{l}(T)=Prm_{r}(T)$.
+    $T$ is a quasi-duo ring. In particular, $\operatorname{Max}(T)=\operatorname{{Max}_{\ell}}(T)=\operatorname{{Max}_{r}}(T)=\operatorname{{Prm}_{\ell}}(T)=\operatorname{{Prm}_{r}}(T)$.
   2. (2)
-    For each $M\in Max(T)$, the center of $T/M$ is integrally closed in $T/M$.
+    For each $M\in\operatorname{Max}(T)$, the center of $T/M$ is integrally closed in $T/M$.
   3. (3)
-    $J(T)=\bigcap_{M\in Max(T)}M$ and $T/J(T)$ is a reduced ring. In particular, $N(T)\subseteq J(T)$. Moreover, the center of $T/J(T)$ is integrally closed in $T/J(T)$.
+    $J(T)=\bigcap_{M\in\operatorname{Max}(T)}M$ and $T/J(T)$ is a reduced ring. In particular, $N(T)\subseteq J(T)$. Moreover, the center of $T/J(T)$ is integrally closed in $T/J(T)$.
 Proof.
-If $T$ has a maximal left ideal $M$, which is not an ideal of $T$, then by Corollary 2.2, we deduce that $|K|\leq|Max_{l}(T)\setminus Max(T)|$.
-Thus by $(1)$ of Proposition 3.10, we conclude that $|K|\leq|RgMax(T)|$.
-Hence if $|RgMax(T)|<|K|$ we deduce that each maximal left ideal of $T$ is an ideal of $T$, i.e., $T$ is a left quasi duo ring, and by a similar proof $T$ is a right quasi duo ring.
-Hence $T$ is a quasi duo ring.
+If $T$ has a maximal left ideal $M$, which is not an ideal of $T$, then by Corollary 2.2, we deduce that $|K|\leq|\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)|$.
+Thus by Proposition 3.10(1)(1), we conclude that $|K|\leq|\operatorname{RgMax}(T)|$.
+Hence if $|\operatorname{RgMax}(T)|<|K|$ we obtain that each maximal left ideal of $T$ is an ideal of $T$, i.e., $T$ is a left quasi-duo ring, and by an analogous proof, $T$ is a right quasi-duo ring.
+Consequently $T$ is a quasi-duo ring.
 Thus $(1)$ holds.
 $(2)$ is an immediate consequence of Proposition 3.1.
-For $(3)$, by $(1)$ we infer that $J(T)$ is the intersection of all maximal ideals of $T$ and therefore $T/J(T)$ embeds in the reduce ring $\prod_{M\in Max(T)}T/M$ (note that $T$ is a quasi duo ring and therefore $T/M$ is a division ring for each $M\in Max(T)$).
+For $(3)$, by $(1)$ we infer that $J(T)$ is the intersection of all maximal ideals of $T$ and therefore $T/J(T)$ embeds in the reduce ring $\prod_{M\in\operatorname{Max}(T)}T/M$ (note that $T$ is a quasi-duo ring and therefore $T/M$ is a division ring for each $M\in\operatorname{Max}(T)$).
 Hence $T/J(T)$ is a reduced ring and thus $N(T)\subseteq J(T)$.
-The final part of $(3)$ is obvious by Lemma 3.13 (note, $T/J(T)$ is an infinite ring, for it contains a copy of the infinite field $K$).
+The last part of $(3)$ is obvious by Lemma 3.13 (note, $T/J(T)$ is an infinite ring, because it contains a copy of the infinite field $K$).
 ∎
 Corollary 3.15.
-Let $T$ be a prime $K$-algebra over an infinite field $K$ with $J(T)=0$, then either $|RgMax(T)|\geq|K|$ or $T$ embeds in a division ring (therefore $T$ is a domain).
+Let $T$ be a prime $K$-algebra over an infinite field $K$ with $J(T)=0$, then either $|\operatorname{RgMax}(T)|\geq|K|$ or $T$ embeds in a division ring; in particular, $T$ is a domain.
 Proof.
-Assume that $|RgMax(T)|<|K|$, then by the proof of the previous theorem $T$ embeds in the reduced ring $R:=\prod_{M\in Max(T)}T/M$.
+Assume that $|\operatorname{RgMax}(T)|<|K|$, then by the proof of the previous theorem $T$ embeds in the reduced ring $R:=\prod_{M\in\operatorname{Max}(T)}T/M$.
 Thus $T$ is a reduce prime ring.
@@
 Since for each maximal ideal $M$ of $T$, $T/M$ is a division ring, we infer that $T$ is a strongly von Neumann regular ring.
-Hence by [12, Theorem 1.2.4], we conclude that $T$ is embeddable in a division ring (note, in fact $T$ is a division ring, for $T$ is a strongly von Neumann regular domain and therefore $C(T)$ is integrally closed in $T$, by Proposition 3.1).
-∎
-If $R$ is a commutative ring and $M$ be a maximal ideal of $R$ such that $Char(R/M)=0$, then $R/M$ is a field with infinitely many maximal subrings, see [10, Corollary 1.5].
-Therefore $R$ has infinitely many maximal subrings.
-For non-commutative rings we have the following result.
+Hence by [12, Theorem 1.2.4], we conclude that $T$ is embeddable in a division ring, as desired.∎
+If $R$ is a commutative ring and $M$ is a maximal ideal of $R$ such that $\operatorname{Char}(R/M)=0$, then $R/M$ is a field with infinitely many maximal subrings, by [10, Corollary 1.5].
+Consequently, $R$ itself has infinitely many maximal subrings.
+For noncommutative rings we obtain the following analogue.
 Corollary 3.16.
@@
   1. (1)
-    For each $M\in Max_{l}^{\prime}(T)\cup Max_{r}^{\prime}(T)$, there exists a prime number $p$ such that $pT\subseteq M$.
+    For each $M\in\operatorname{{Max}_{\ell}}^{\prime}(T)\cup\operatorname{{Max}_{r}}^{\prime}(T)$, there exists a prime number $p$ such that $pT\subseteq M$.
   2. (2)
-    For each $M\in Max^{\prime}(T)$, there exists a prime number $p$ such that $pT\subseteq M$.
+    For each $M\in\operatorname{Max}^{\prime}(T)$, there exists a prime number $p$ such that $pT\subseteq M$.
   3. (3)
-    If $P$ is a left/right primitive ideal of $T$, then either $P$ is a maximal left and right ideal of $T$ or $Char(T/P)$ is a prime number.
+    If $P$ is a left/right primitive ideal of $T$, then either $T/P$ is a division ring or $\operatorname{Char}(T/P)$ is a prime number.
 Proof.
 $(1)$ Let $M$ be a maximal left/right ideal of $T$ which is not an ideal of $T$.
-Since $T$ has only finitely many maximal subring, by $(2)$ of Proposition 3.10, we conclude that $\mathbb{I}(M)/M$ is a finite field.
-Hence $Char(\mathbb{I}(M)/M)=p$, where $p$ is a prime number.
-Thus $p.1\in M$, for $1\in\mathbb{I}(M)$.
+Since $T$ has only finitely many maximal subring, by Proposition 3.10(2)(2), we conclude that $\mathbb{I}(M)/M$ is a finite field.
+Hence $\operatorname{Char}(\mathbb{I}(M)/M)=p$, where $p$ is a prime number.
+Thus $p.1\in M$, because $1\in\mathbb{I}(M)$.
 Therefore $pT\subseteq M$.
-$(2)$ The proof is similar to $(1)$ and using $(4)$ of Proposition 3.10.
+$(2)$ The proof is similar to $(1)$ and by the use of Proposition 3.10(4)(4).
 $(3)$ Let $P$ be a left primitive ideal of $T$, which is not a left/right maximal ideal of $T$.
-Hence there exists a maximal left ideal $M$ of $T$ such that $P=l.ann_{T}(T/M)$.
-If $M$ is an ideal of $T$, then $M=l.ann_{T}(T/M)=P$, which is absurd.
+Hence there exists a maximal left ideal $M$ of $T$ such that $P=\operatorname{l.ann}_{T}(T/M)$.
+If $M$ is an ideal of $T$, then $M=\operatorname{l.ann}_{T}(T/M)=P$, which is absurd.
 Thus $M$ is not an ideal of $T$ and therefore by $(1)$, we deduce that there exists a prime number $p$ such that $p(T/M)=0$.
-Hence $p\in P$ and thus $Char(T/P)=p$.
+Hence $p\in P$ and thus $\operatorname{Char}(T/P)=p$.
 ∎
-By the previous result, if $T$ is a ring with zero characteristic and there exists a maximal left/right ideal $M$ of $T$ which is not an ideal of $T$ and either $M\cap\mathbb{Z}=0$ or $T/M$ is a torsionfree/faithful $\mathbb{Z}$-module, then $T$ has infinitely many maximal subrings.
+By the previous result, if $T$ is a ring of characteristic zero and there exists a maximal left/right ideal $M$ of $T$ that is not an ideal and moreover either $M\cap\mathbb{Z}=0$ or $T/M$ is a torsion-free/faithful $\mathbb{Z}$-module, then $T$ has infinitely many maximal subrings.
 Proposition 3.17.
-Let $T$ be an infinite ring with nonzero characteristic. If $Max_{F}(T):=\{M\in Max(T)\ |\ T/M\ \text{is}\ \text{finite}\}$ is infinite, then $T$ has infinitely many maximal subrings.
+Let $T$ be an infinite ring with nonzero characteristic.
+If $\operatorname{Max}_{F}(T):=\{M\in\operatorname{Max}(T)\ |\ T/M\ \text{is}\ \text{finite}\}$ is infinite, then $T$ has infinitely many maximal subrings.
 Proof.
-Since $n$ has only finitely many prime divisors and for each $M\in Max_{F}(T)$, $Char(T/M)$ is a prime divisor of $n$, we conclude that the exists a prime divisor $p$ of $n$ such that $Max_{p,F}(T):=\{M\in Max_{F}(T)\ |\ Char(T/M)=p\}$ is infinite.
-Also note that for each $M\in Max_{p,F}(T)$, the ring $T/M$ is a finite simple ring, hence we deduce that there exist a natural number $m$ and a finite field $E$ such that $T/M\cong\mathbb{M}_{m}(E)$.
+Since $n$ has only finitely many prime divisors and for each $M\in\operatorname{Max}_{F}(T)$, $\operatorname{Char}(T/M)$ is a prime divisor of $n$, we conclude that the exists a prime divisor $p$ of $n$ such that $\operatorname{Max}_{p,F}(T):=\{M\in\operatorname{Max}_{F}(T)\ |\ \operatorname{Char}(T/M)=p\}$ is infinite.
+Also note that for each $M\in\operatorname{Max}_{p,F}(T)$, the ring $T/M$ is a finite simple ring, hence we deduce that there exist a natural number $m$ and a finite field $E$ such that $T/M\cong\mathbb{M}_{m}(E)$.
 Now we have two cases:
-$(1)$ there exist distinct $M_{1},M_{2},\ldots\in Max_{p,F}(T)$ such that $T/M_{i}\cong\mathbb{M}_{m_{i}}(E_{i})$, where $m_{i}>1$ and $E_{i}$ is a field.
-Thus in this case for each $i$, $T$ has a maximal subring $R_{i}$ which contains $M_{i}$, for $\mathbb{M}_{m_{i}}(E_{i})$ has a maximal subring by $(4)$ of Theorem 2.7.
+$(\bf 1)$ There exist distinct $M_{1},M_{2},\ldots\in\operatorname{Max}_{p,F}(T)$ such that $T/M_{i}\cong\mathbb{M}_{m_{i}}(E_{i})$, where $m_{i}>1$ and $E_{i}$ is a field.
+Thus in this case for each $i$, $T$ has a maximal subring $R_{i}$ which contains $M_{i}$, for $\mathbb{M}_{m_{i}}(E_{i})$ has a maximal subring by Theorem 2.7(4)(4).
 Since for $i\neq j$, $M_{i}+M_{j}=T$, we deduce that $R_{i}\neq R_{j}$.
-Hence $R_{1},R_{2},\ldots$ are distinct maximal subrings of $T$ and hence we are done.
-$(2)$ thus assume that there exist infinitely many distinct maximal ideals $M_{1},M_{2},\ldots\in Max_{p,F}(T)$ such that $T/M_{i}=E_{i}$ is a finite field.
+Hence $R_{1},R_{2},\ldots$ are distinct maximal subrings of $T$, as desired.
+$(\bf 2)$ Thus assume that there exist infinitely many distinct maximal ideals $M_{1},M_{2},\ldots\in\operatorname{Max}_{p,F}(T)$ such that $T/M_{i}=E_{i}$ is a finite field.
 Clearly by our assumption $\mathbb{Z}_{p}\subseteq E_{i}$, for each $i\geq 1$.
 Now if for infinitely many $i$, $\mathbb{Z}_{p}\subsetneq E_{i}$, then similar to proof of $(1)$ we obtain that $T$ has infinitely many maximal subrings (note that, $E_{i}$ is a finite field and $\mathbb{Z}_{p}\subsetneq E_{i}$, therefore $E_{i}$ has a maximal subring, see the comments preceding Theorem 2.7, and hence $T$ has a maximal subring $R_{i}$ which contains $M_{i}$).
-Thus without lose of the generality we may assume that for each $i\geq 1$, we have $\mathbb{Z}_{p}=E_{i}$.
+Thus without loss of generality, we may assume that for each $i\geq 1$, we have $\mathbb{Z}_{p}=E_{i}$.
 Let $P_{i}=M_{2i}$ and $Q_{i}=M_{2i-1}$, where $i\geq 1$.
@@
 ∎
-It is not hard to see that by the proof of the previous result, in fact there exists a prime divisor $p$ of $n$ such that $|Max_{p,F}(T)|=|Max_{F}(T)|\leq|RgMax(T)|$ for some prime number $p$.
+It is not hard to see that by the proof of the previous result, in fact there exists a prime divisor $p$ of $n$ such that $|\operatorname{Max}_{p,F}(T)|=|\operatorname{Max}_{F}(T)|\leq|\operatorname{RgMax}(T)|$ for some prime number $p$.
 Now the next result is in order.
 Corollary 3.18.
-Let $T$ be a ring. If $|Max_{F}(T)|>\aleph_{0}$, then $T$ has infinitely many maximal subrings. In fact, $|Max_{F}(T)|\leq|RgMax(T)|$.
+Let $T$ be a ring. If $|\operatorname{Max}_{F}(T)|>\aleph_{0}$, then $T$ has infinitely many maximal subrings. In fact, $|\operatorname{Max}_{F}(T)|\leq|\operatorname{RgMax}(T)|$.
 Proof.
-Since $|Max_{F}(T)|>\aleph_{0}$ and the number of prime numbers is countable, we conclude that there exists a prime number $p$ such that $|Max_{p,F}(T)|=|Max_{F}(T)|$.
-Now by a similar proof of Proposition 3.19, we can see that $|Max_{p,F}(T)|\leq|RgMax(T)|$ and hence we are done.
+Since $|\operatorname{Max}_{F}(T)|>\aleph_{0}$ and the number of prime numbers is countable, we conclude that there exists a prime number $p$ such that $|\operatorname{Max}_{p,F}(T)|=|\operatorname{Max}_{F}(T)|$.
+Now by a similar proof of Proposition 3.17, we can see that $|\operatorname{Max}_{p,F}(T)|\leq|\operatorname{RgMax}(T)|$ and hence we are done.
 ∎
@@
   1. (1)
-    If $E_{1}=End_{T}(S_{1})$ is a field which is not an absolutely algebraic field.
+    If $E_{1}=\operatorname{End}_{T}(S_{1})$ is a field which is not an absolutely algebraic field.
   2. (2)
-    If $E_{1}=End_{T}(S_{1})$ and $E_{2}=End_{T}(S_{2})$ are infinite isomorphic ring.
+    If $E_{1}=\operatorname{End}_{T}(S_{1})$ and $E_{2}=\operatorname{End}_{T}(S_{2})$ are infinite isomorphic ring.
 Proof.
-First note that since $S_{i}$ is a simple left $T$-module, we infer that there exists maximal left ideals $M_{i}$, $i=1,2$, such that $S_{i}\cong T/M_{i}$, as left $T$-module.
-$(1)$ Note that $E_{1}^{op}\cong\mathbb{I}(M_{1})/M_{1}$.
-Now we have two cases, either $M_{1}$ is an ideal of $T$ or not.
-If $M_{1}$ is an ideal of $T$, then $T/M_{1}\cong E_{1}^{op}$ as ring and therefore $T/M_{1}$ has infinitely many maximal subrings, by [10, Corollary 1.5].
+First note that since $S_{1}$ and $S_{2}$ are simple left $T$-modules, we infer that there exist maximal left ideals $M_{1}$ and $M_{2}$, such that $S_{i}\cong T/M_{i}$, as left $T$-modules for $i=1,2$.
+$(\bf 1)$ Note that $E_{1}^{op}\cong\mathbb{I}(M_{1})/M_{1}$.
+Now we have two cases, either $M_{1}$ is an ideal of $T$ or is not.
+If $M_{1}$ is an ideal of $T$, then $T/M_{1}\cong E_{1}^{op}$ as rings and therefore $T/M_{1}$ has infinitely many maximal subrings, by [10, Corollary 1.5].
 Hence assume that $M_{1}$ is not an ideal of $T$.
-Since $E_{1}$ is not an absolutely algebraic field, we conclude that $E_{1}$ is an infinite ring.
-Therefore $\mathbb{I}(M_{1})/M_{1}$ is an infinite ring.
-Thus by Theorem 3.1, we conclude that $M_{1}]$ and therefore $Max_{l}(T)\setminus Max(T)$ are infinite and hence we are done by $(1)$ of Proposition 3.10.
-$(2)$ We have two cases, if $M_{1}$ or $M_{2}$ is not an ideal of $T$.
-Then similar to last part of the proof of $(1)$ we infer that $T$ has infinitely many maximal subrings.
-Hence assume that $M_{1}$ and $M_{2}$ are ideals of $T$.
-In this case we have the following ring isomorphisms
+Since $E_{1}$ is not an absolutely algebraic field, we deduce that $E_{1}$ is an infinite ring.
+Consequently $\mathbb{I}(M_{1})/M_{1}$ is an infinite ring.
+Thus by Theorem 3.1, we obtain that $M_{1}]$ and a fortiori $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$ are infinite.
+It follows that from Proposition 3.10(1)(1), $T$ has infinitely many maximal subrings, as desired.
+$(\bf 2)$ We have two cases, if $M_{1}$ or $M_{2}$ is not an ideal of $T$.
+Then analogous to last part of the proof of $(1)$, we infer that $T$ has infinitely many maximal subrings.
+Thus suppose that $M_{1}$ and $M_{2}$ are ideals of $T$.
+In this case, we have the following ring isomorphisms
 $T/M_{1}=\mathbb{I}(M_{1})/M_{1}\cong E_{1}^{op}\cong E_{2}^{op}\cong\mathbb{I}(M_{2})/M_{2}=T/M_{2}.$
-Thus $T/(M_{1}\cap M_{2})\cong E_{1}^{op}\times E_{1}^{op}$, as ring (note $M_{1}\neq M_{2}$, for $S_{1}\ncong S_{2}$).
-Now since $E_{1}$ and $E_{2}$ are infinite division rings, by Theorem 3.6, we deduce that $T/(M_{1}\cap M_{2})$ has infinitely many maximal subrings.
-Therefore $T$ has infinitely many maximal subrings.
+Consequently $T/(M_{1}\cap M_{2})\cong E_{1}^{op}\times E_{1}^{op}$, as rings (note $M_{1}\neq M_{2}$, because $S_{1}\ncong S_{2}$).
+Since $E_{1}$ and $E_{2}$ are infinite division rings, we deduce that $T/(M_{1}\cap M_{2})$ has infinitely many maximal subrings, by Theorem 3.6.
+Therefore $T$ has infinitely many maximal subrings, which complete the proof.
 ∎
 4. Rings integral over their centers
-In this section we investigate about the number of maximal subrings in a ring $T$ which is integral over its center.
-First we need some observation and notation.
-Let $T$ be a ring, then it is clear that $Max_{l}(T)\cap Max(T)=Max_{r}(T)\cap Max(T)=Max_{r}(T)\cap Max_{l}(T)$, we denote this set by $Max_{lr}(T)$ and we define $J^{\prime}(T):=\bigcap_{M\in Max_{lr}(T)}M$.
-It is clear that $J^{\prime}(T)$ is an ideal of $T$ and $J(T)\subseteq J^{\prime}(T)$.
-In particular, if $T$ is either left or right quasi duo ring, then $J(T)=J^{\prime}(T)$.
-Moreover if $Max_{l}(T)\setminus Max(T)=\{M_{1},\ldots,M_{n}\}$ is finite (in particular, by $(5)$ of Proposition 3.10, where $T$ has only finitely many maximal subrings), then $J(T)=J^{\prime}(T)\cap M_{1}\cap\cdots\cap M_{n}$.
-Note that by Lemma 3.13, if $T$ is an infinite ring with $J^{\prime}(T)=0$, then either $T$ has infinitely many maximal subrings or $C(T)$ is integrally closed in $T$.
-Now the following is in order.
+In this section we investigate on the number of maximal subrings in a ring $T$ that is integral over its center.
+First we need some preliminary observations and notation.
+Let $T$ be a ring.
+Then it is clear that $\operatorname{{Max}_{\ell}}(T)\cap\operatorname{Max}(T)=\operatorname{{Max}_{r}}(T)\cap\operatorname{Max}(T)=\operatorname{{Max}_{r}}(T)\cap\operatorname{{Max}_{\ell}}(T)$, we denote this set by $\operatorname{Max}_{lr}(T)$.
+Define $J^{\prime}(T):=\bigcap_{M\in\operatorname{Max}_{lr}(T)}M$.
+Then $J^{\prime}(T)$ is an ideal of $T$ and $J(T)\subseteq J^{\prime}(T)$.
+In particular, if $T$ is either left or right quasi-duo ring, then $J(T)=J^{\prime}(T)$.
+Moreover, if $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)=\{M_{1},\ldots,M_{n}\}$ is finite (which occur, in particular, when $T$ has only finitely many maximal subrings, by Proposition 3.10(5)(5)), then $J(T)=J^{\prime}(T)\cap M_{1}\cap\cdots\cap M_{n}$.
+Note also that by Lemma 3.13, if $T$ is an infinite ring with $J^{\prime}(T)=0$, then either $T$ has infinitely many maximal subrings or $C(T)$ is integrally closed in $T$.
+We now state the following result.
 Proposition 4.1.
@@
   2. (2)
-    $|Max_{lr}(T)|\leq 2^{\aleph_{0}}$. For each $M\in Max_{lr}(T)$, $T/M$ is an absolutely algebraic field (thus $|T/M|\leq\aleph_{0}$).
+    $|\operatorname{Max}_{lr}(T)|\leq 2^{\aleph_{0}}$. For each $M\in\operatorname{Max}_{lr}(T)$, $T/M$ is an absolutely algebraic field (thus $|T/M|\leq\aleph_{0}$).
   3. (3)
-    for each $M\in Max_{l}^{\prime}(T)\cup Max_{r}^{\prime}(T)$, $J^{\prime}(T)+M=T$.
+    For each $M\in\operatorname{{Max}_{\ell}}^{\prime}(T)\cup\operatorname{{Max}_{r}}^{\prime}(T)$, $J^{\prime}(T)+M=T$.
 Proof.
-Assume that $T$ has only finitely many maximal subrings.
-Thus for each $M\in Max_{lr}(T)$, the division ring $T/M$ has only finitely many maximal subrings.
-Since $T$ is integral over its center, we conclude that for each $M\in Max_{lr}(T)$, the division ring $T/M$ is integral over its center.
+$(1)$ Assume that $T$ has only finitely many maximal subrings.
+Thus for each $M\in\operatorname{Max}_{lr}(T)$, the division ring $T/M$ has only finitely many maximal subrings.
+Since $T$ is integral over its center, we conclude that for each $M\in\operatorname{Max}_{lr}(T)$, the division ring $T/M$ is integral over its center.
 Thus by Proposition 3.1, we conclude that $T/M$ is a field.
-Clearly, $T/J^{\prime}(T)$ embeds in $\prod_{M\in Max_{lr}(T)}T/M$, thus $T/J^{\prime}(T)$ is a commutative reduced ring and therefore $N(T)\subseteq J^{\prime}(T)$.
+Clearly, $T/J^{\prime}(T)$ embeds in $\prod_{M\in\operatorname{Max}_{lr}(T)}T/M$, thus $T/J^{\prime}(T)$ is a commutative reduced ring and a fortiori $N(T)\subseteq J^{\prime}(T)$.
 Since $T/J^{\prime}(T)$ is a commutative ring with only finitely many maximal subrings, by [10, Corollary 1.9], we deduce that $T/J^{\prime}(T)$ is a Hilbert ring.
-By [10, Corollary 1.5], for each $M\in Max_{lr}(T)$, the field $T/M$ is an absolutely algebraic field and therefore $|T/M|\leq\aleph_{0}$.
-Since $T/J^{\prime}(T)$ is a commutative ring, the first part of $(2)$ is an immediate consequence of [10, (1) of Proposition 2.8].
+By [10, Corollary 1.5], for each $M\in\operatorname{Max}_{lr}(T)$, the field $T/M$ is an absolutely algebraic field and consequently $|T/M|\leq\aleph_{0}$.
 Therefore we deduce that
-$|\frac{T}{J^{\prime}(T)}|\leq|\prod_{M\in Max_{lr}(T)}T/M|\leq 2^{2^{\aleph_{0}}}.$
-Hence $(1)$ and $(2)$ hold.
-Finally for $(3)$, let $M$ be a maximal left ideal of $T$ which is not an ideal of $T$.
-Since $T/J^{\prime}(T)$ is a commutative ring, we conclude that $J^{\prime}(T)\nsubseteq M$ and therefore $M+J^{\prime}(T)=T$ by maximality of $M$.
+$|\frac{T}{J^{\prime}(T)}|\leq|\prod_{M\in\operatorname{Max}_{lr}(T)}T/M|\leq 2^{2^{\aleph_{0}}}.$
+$(2)$ Since $T/J^{\prime}(T)$ is a commutative ring, the first part of $(2)$ is an immediate consequence of [10, (1) of Proposition 2.8]. For the last part is evident by the proof of $(1)$.
+$(3)$ Let $M$ be a maximal left ideal of $T$ which is not an ideal of $T$.
+Since $T/J^{\prime}(T)$ is a commutative ring, we conclude that $J^{\prime}(T)\nsubseteq M$, at once $M+J^{\prime}(T)=T$ by maximality of $M$.
 ∎
-As we mentioned in the proof of the previous proposition, if $R$ is a commutative ring with $|Max(R)|>2^{\aleph_{0}}$, then $|RgMax(R)|\geq 2^{\aleph_{0}}$, see [10, (1) of Proposition 2.8].
-For non-commutative rings we have the following result.
+As noted in the proof of the previous proposition, if $R$ is a commutative ring with $|\operatorname{Max}(R)|>2^{\aleph_{0}}$, then $|\operatorname{RgMax}(R)|\geq 2^{\aleph_{0}}$, see [10, Proposition 2.8(1)].
+For noncommutative rings we obtain the following analogue.
 Proposition 4.2.
@@
   1. (1)
-    $|Max_{l}(T)|>2^{\aleph_{0}}$.
+    $|\operatorname{{Max}_{\ell}}(T)|>2^{\aleph_{0}}$.
   2. (2)
-    $|Max_{r}(T)|>2^{\aleph_{0}}$.
+    $|\operatorname{{Max}_{r}}(T)|>2^{\aleph_{0}}$.
   3. (3)
-    $|Max(T)|>2^{\aleph_{0}}$.
+    $|\operatorname{Max}(T)|>2^{\aleph_{0}}$.
 Proof.
-We prove $(1)$ and $(3)$, the proof of $(2)$ is similar to $(1)$.
-For $(1)$, first note that if $Max_{l}(T)\setminus Max(T)$ is infinite, then by $(1)$ of Proposition 3.10, we deduce that $T$ has infinitely many maximal subrings.
-Hence assume that $Max_{l}(T)\setminus Max(T)$ is finite.
-Thus $|Max_{l}(T)\cap Max(T)|=|Max_{l}(T)|>2^{\aleph_{0}}$.
-Now since $Max_{lr}(T)=Max_{l}(T)\cap Max(T)$, we deduce that $T$ has infinitely many maximal subrings by $(2)$ of Proposition 4.1.
-$(3)$ If $Max(T)\setminus Max_{l}(T)$ is infinite, then by $(5)$ of Proposition 3.10, we conclude that $T$ has infinitely many maximal subrings.
-Hence assume that $Max(T)\setminus Max_{l}(T)$ is finite and therefore $|Max(T)\cap Max_{l}(T)|=|Max(T)|>2^{\aleph_{0}}$.
-By a similar proof of $(1)$ we obtain that $T$ has infinitely many maximal subrings.
+We prove $(1)$ and $(3)$. The proof of $(2)$ is similar to $(1)$.
+$(1)$ First note that if $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$ is infinite, then by Proposition 3.10(1)(1), we deduce that $T$ has infinitely many maximal subrings.
+Hence assume that $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$ is finite.
+Thus $|\operatorname{{Max}_{\ell}}(T)\cap\operatorname{Max}(T)|=|\operatorname{{Max}_{\ell}}(T)|>2^{\aleph_{0}}$.
+Now since $\operatorname{Max}_{lr}(T)=\operatorname{{Max}_{\ell}}(T)\cap\operatorname{Max}(T)$, we deduce that $T$ has infinitely many maximal subrings by Proposition 4.1(2)(2).
+$(3)$ If $\operatorname{Max}(T)\setminus\operatorname{{Max}_{\ell}}(T)$ is infinite, then by Proposition 3.10(5)(5), we conclude that $T$ has infinitely many maximal subrings.
+Hence assume that $\operatorname{Max}(T)\setminus\operatorname{{Max}_{\ell}}(T)$ is finite and therefore $|\operatorname{Max}(T)\cap\operatorname{{Max}_{\ell}}(T)|=|\operatorname{Max}(T)|>2^{\aleph_{0}}$.
+By an analogous proof of $(1)$, we obtain that $T$ has infinitely many maximal subrings.
 ∎
-As the following main result shows if we assume that a ring $T$ is an algebraic $K$-algebra (and therefore is integral over its center), then either $T$ has infinitely many maximal subrings or $T$ has similar algebraic properties like commutative rings with only finitely many maximal subrings, see [10].
+As the following main result shows, if a ring $T$ is an algebraic $K$-algebra (and therefore is integral over its center), then either $T$ has infinitely many maximal subrings, or $T$ exhibits algebraic properties similar to commutative rings with only finitely many maximal subrings, see [10].
 Theorem 4.3.
@@
   1. (1)
-    $T$ is a quasi duo ring. In particular, $J(T)=\bigcap_{M\in Max(T)}M$.
+    $T$ is a quasi-duo ring. In particular, $J(T)=\bigcap_{M\in\operatorname{Max}(T)}M$.
   2. (2)
-    For each $M\in Max(T)$, the ring $T/M$ is an absolutely algebraic field which contains a copy of $K$.
-In particular, $K$ is an absolutely algebraic field and $|K|\leq|T/M|\leq\aleph_{0}$.
+    For each $M\in\operatorname{Max}(T)$, the ring $T/M$ is an absolutely algebraic field which contains a copy of $K$.
+In particular, $K$ is an absolutely algebraic field and $|K|=|T/M|=\aleph_{0}$.
 Moreover, if $D$ is a division ring which is a subring of $T$, then $D$ is an absolutely algebraic field.
   3. (3)
-    $|Max(T)|\leq 2^{\aleph_{0}}$ and $|T/J(T)|\leq 2^{2^{\aleph_{0}}}$.
+    $|\operatorname{Max}(T)|\leq 2^{\aleph_{0}}$ and $|T/J(T)|\leq 2^{2^{\aleph_{0}}}$.
   4. (4)
     $T/J(T)$ is a reduced Hilbert commutative ring.
-In fact, $T/J(T)$ is a von Neumann regular ring which is integral over its prime subring (i.e., $\mathbb{Z}_{p}$, where $p=Char(K)$ is a prime number).
+In fact, $T/J(T)$ is a von Neumann regular ring which is integral over its prime subring (i.e., $\mathbb{Z}_{p}$, where $p=\operatorname{Char}(K)$ is a prime number).
 Moreover, $T$ is integral over $\mathbb{Z}_{p}$.
@@
 Thus by Theorem 3.14, we deduce that $(1)$ holds.
-The first part of $(2)$ is clear by $(2)$ of Proposition 4.1.
+The first part of $(2)$ is clear by Proposition 4.1(2)(2).
 For the second part of $(2)$, let $M$ be any maximal ideal of $T$.
 Clearly, $T/M$ contains a copy of $K$ (and similarly $D$), thus $K$ and $D$ are absolutely algebraic fields.
-$(3)$ and the first part of $(4)$ are clear by Proposition 4.1.
+Consequently, $\aleph_{0}\leq|K|\leq|T/M|\leq\aleph_{0}$ and thus the equalities hold.
+$(3)$ and the first part of $(4)$ are evident by Proposition 4.1.
 For the second part of $(4)$, note that $T/J(T)$ is a commutative ring which is integral over the field $K$.
-Hence we deduce that $T/J(T)$ is a zero-dimensional (reduced) ring and therefore is a von Neumann ring.
+Hence we deduce that $T/J(T)$ is a zero-dimensional (reduced) ring and a fortiori, it is a von Neumann ring.
 Since $T$ has only finitely many maximal subrings, we infer that $T/J(T)$ has only finitely many maximal subring.
@@
 Since $T$ is algebraic over $K$, by [21, Theorem 4.20], $J(T)$ is a nil ideal of $T$.
-Hence if $x\in T$, then there there exist $n\in\mathbb{N}$ and $a_{0},\ldots,a_{n-1}\in\mathbb{Z}_{p}$ such that $(x+J(T))^{n}+a_{n-1}(x+J(T))^{n-1}+\cdots+a_{1}(x+J(T))+a_{0}=0$ in the ring $T/J(T)$.
+Now let $x\in T$, then there exist $n\in\mathbb{N}$ and $a_{0},\ldots,a_{n-1}\in\mathbb{Z}_{p}$ such that $(x+J(T))^{n}+a_{n-1}(x+J(T))^{n-1}+\cdots+a_{1}(x+J(T))+a_{0}=0$ in the ring $T/J(T)$.
 Thus $x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0}\in J(T)$.
-Now since $J(T)$ is nil, we deduce that there exists a natural number $m$ such that $(x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0})^{m}=0$, i.e., $x$ is integral over $\mathbb{Z}_{p}$.
+From the fact that $J(T)$ is nil, we deduce that there exists a natural number $m$ such that $(x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0})^{m}=0$, i.e., $x$ is integral over $\mathbb{Z}_{p}$.
 For $(5)$, assume that $T/M\cong T/N$.
@@
 Hence by Proposition 3.6, $T/M\times T/M$ has infinitely many maximal subrings.
-Hence we conclude that $T$ has infinitely many maximal subrings, which is absurd.
-For $(6)$, since $T/I$ contains a copy of $K$ and remains algebraic over $K$ we may assume that $I=0$.
-Now note that $T/J(T)$ is a reduced ring.
-Therefore $N(T)\subseteq J(T)$.
-As we see in the proof of $(4)$, $J(T)$ is nil and thus $J(T)=N(T)$.
+It follows that $T$ has infinitely many maximal subrings, which is absurd.
+For $(6)$, nothing that $T/I$ contains a copy of $K$ and remains algebraic over $K$, thus we may assume that $I=0$.
+Because $T/J(T)$ is a reduced ring, it is obvious that $N(T)\subseteq J(T)$.
+Consequently, as we see in the proof of $(4)$, $J(T)$ is nil and then $J(T)=N(T)$.
 For $(7)$, note that $T/Q$ has no nonzero nil ideal, hence by $(6)$ we conclude that $J(T/Q)=0$, i.e., $Q$ is an intersection of a family of maximal (left) ideal of $T$.
-Thus $J(T)\subseteq Q$ and therefore by $(4)$ we immediately deduce that $T/Q$ is a field, i.e., $Q$ is a maximal (left) ideal of $T$.
+Thus $J(T)\subseteq Q$ and then by $(4)$, we deduce that $T/Q$ is a field, i.e., $Q$ is a maximal (left) ideal of $T$.
 Finally for $(8)$, let $M$ be a maximal ideal of $T$ and $R$ be a subring of $T$.
-Since $(R+M)/M$ is a subring of the absolutely algebraic field $T/M$, we deduce that $(R+M)/M$ is a field.
-Therefore $R/(R\cap M)$ is a field.
-Hence $R\cap M\in Max(R)$.
-This immediately implies that $J(R)\subseteq J(T)\cap R$ and therefore $J(R)$ is a nil ideal (note, $J(T)$ is nil by $(6)$).
+Since $(R+M)/M$ is a subring of the absolutely algebraic field $T/M$, we obtain that $(R+M)/M$ is a field.
+It follows that $R/(R\cap M)$ is a field, i.e., $R\cap M\in\operatorname{Max}(R)$.
+This implies that $J(R)\subseteq J(T)\cap R$ and at once $J(R)$ is a nil ideal (note, $J(T)$ is nil by $(6)$).
 ∎
@@
 If $P$ is a prime ideal of $T$ with $J(T/P)=0$, then either $P$ is a maximal ideal of $T$ or $T/P$ is an integral domain.
-In fact, either $T/P\cong\mathbb{M}_{n}(F)$ for some natural number $n>1$ and a finite field $F$ or $T/P$ is an integral domain.
+In fact, either $T/P\cong\mathbb{M}_{n}(F)$ for some natural number $n>1$ and finite field $F$ or $T/P$ is an integral domain.
 Proof.
-Since $T$ has only finitely many maximal subrings, then by $(1)$ of Proposition 3.10 we deduce that $Max_{l}^{\prime}(T)$ is finite.
-Hence by $J(T/P)=0$ we infer that there exist $M_{1},\ldots,M_{n}\in Max_{l}^{\prime}(T)$, $n\geq 0$ and $N=\bigcap_{i\in I}M_{i}$, where $M_{i}\in Max(R)\cap Max_{l}(T)$ such that $P=M_{1}\cap\cdots\cap M_{n}\cap N$.
-Thus by the comments preceding Proposition 3.10, either $P=M_{1}\cap\cdots\cap M_{n}$ or $P=N$, for $P$ is a prime ideal of $T$.
+Since $T$ has only finitely many maximal subrings, then by Proposition 3.10(1)(1) we deduce that $\operatorname{{Max}_{\ell}}^{\prime}(T)$ is finite.
+Hence by $J(T/P)=0$ we infer that there exist $M_{1},\ldots,M_{n}\in\operatorname{{Max}_{\ell}}^{\prime}(T)$, $n\geq 0$ and $N=\bigcap_{i\in I}M_{i}$, where $M_{i}\in\operatorname{Max}(R)\cap\operatorname{{Max}_{\ell}}(T)$ such that $P=M_{1}\cap\cdots\cap M_{n}\cap N$.
+Thus by the comments preceding Proposition 3.10, either $P=M_{1}\cap\cdots\cap M_{n}$ or $P=N$, because $P$ is a prime ideal of $T$.
 If $P=M_{1}\cap\cdots\cap M_{n}$, then $T/P$ is a left $T$-submodule of the semisimple $T$-module $T/M_{1}\times\cdots\times T/M_{n}$.
-Therefore we conclude that $T/P$ is a simple Artinian ring.
+Consequently $T/P$ is a simple Artinian ring.
 Hence $P$ is a maximal ideal of $T$.
@@
 Since $T$ is integral over its center, we infer that for each $i\in I$, the division ring $T/M_{i}$ is integral over its center.
-Hence, if for some $i\in I$, $T/M_{i}$ is not a field, then by Theorem 3.1, we obtain that $T/M_{i}$ and therefore $T$ have infinitely many maximal subrings which is impossible.
+Hence, if for some $i\in I$, $T/M_{i}$ is not a field, then by Proposition 3.1, we obtain that $T/M_{i}$ and therefore $T$ have infinitely many maximal subrings which is impossible.
 Thus for each $i\in I$, $T/M_{i}$ is a field.
 Clearly $T/P$ embeds in $\prod_{i\in I}T/M_{i}$ and therefore $T/P$ is a commutative (reduced prime) ring.
-Hence $T/P$ is an integral domain.
-For the final par, if $T/P$ is not an integral domain, then since $T/P$ is a simple Artinian ring, we conclude that $T/P\cong\mathbb{M}_{n}(D)$, where $D$ is a division ring and $n$ is a natural number.
-Now we have two cases: If $n=1$, then $T/P$ is a division ring which is integral over its center, hence by Theorem 3.1, we deduce that $T/P$ is a field, for $T/P$ has only finitely many maximal subrings.
+It follows that $T/P$ is an integral domain, because $P$ is prime.
+For the last part, if $T/P$ is not an integral domain, then since $T/P$ is a simple Artinian ring, we infer that $T/P\cong\mathbb{M}_{n}(D)$, where $D$ is a division ring and $n$ is a natural number.
+Now we have two cases: If $n=1$, then $T/P$ is a division ring which is integral over its center, hence by Proposition 3.1, we deduce that $T/P$ is a field, for $T/P$ has only finitely many maximal subrings.
 Thus $T/P$ is an integral domain which is absurd.
-Therefore $n>1$.
-We claim that $D$ is finite.
-For otherwise, by Theorem 3.5, we deduce that $T/P$ has infinitely many maximal subrings which is absurd.
-Thus $D$ is a finite field and hence we are done.
+Now suppose that $n>1$, we claim that $D$ is finite.
+Otherwise, by Theorem 3.5, we deduce that $T/P$ has infinitely many maximal subrings which is absurd.
+Thus $D$ is a finite field, as desired.
 ∎
@@
   1. (1)
-    $Prm_{l}(T)=Max(T)=Prm_{r}(T)$.
+    $\operatorname{{Prm}_{\ell}}(T)=\operatorname{Max}(T)=\operatorname{{Prm}_{r}}(T)$.
   2. (2)
@@
 Clearly, $J(T/P)=0$ and therefore by Lemma 4.4, we conclude that either $P$ is a maximal ideal of $T$ or $T/P$ is a commutative left primitive ring.
-Since a commutative primitive ring is a field, we infer that $T/P$ is a field and therefore $P$ is a maximal ideal of $T$.
+Since a commutative primitive ring is a field, we infer that $T/P$ is a field and consequently $P$ is a maximal ideal of $T$.
 Similarly each right primitive ideal of $T$ is a maximal ideal.
@@
 Conversely, let $P$ be a strongly prime ideal of $T$, i.e., $T/P$ has no nonzero nil ideal.
-It is not hard to see that $P\cap C(T)$ is a prime ideal of $C(T)$ and therefore by the assumption is a maximal ideal of $C(T)$.
-Therefore $C(T)/(P\cap C(T))\cong(C(T)+P)/P$ is a field.
-Since $T$ is integral over its center, we immediately conclude that $T/P$ is integral over the field $E:=(C(T)+P)/P$.
-Hence by [21, Theorem 4.20], $J(T/P)$ is a nil ideal and therefore $J(T/P)=0$, for $P$ is a strongly prime ideal of $T$.
-Hence by Lemma 4.4, either $P$ is a maximal ideal of $T$ or $T/P$ is an integral domain.
-In the second case note that since $T/P$ is integral over the field $E$, we immediately conclude that $P$ is maximal too.
+It is not hard to see that $P\cap C(T)$ is a prime ideal of $C(T)$.
+It follows that by the assumption $P\cap C(T)$ is a maximal ideal of $C(T)$.
+This implies that $C(T)/(P\cap C(T))\cong(C(T)+P)/P$ is a field.
+Since $T$ is integral over its center, we conclude that $T/P$ is integral over the field $E:=(C(T)+P)/P$.
+Hence by [21, Theorem 4.20], $J(T/P)$ is a nil ideal and therefore $J(T/P)=0$, because $P$ is a strongly prime ideal of $T$.
+Consequently by Lemma 4.4, either $P$ is a maximal ideal of $T$ or $T/P$ is an integral domain.
+In the second case note that since $T/P$ is integral over the field $E$, we obtain that $P$ is maximal.
 ∎
-In the next main result in this section, we see that if $T$ is a ring which is integral over its center and has only finitely many maximal subrings, then $T/J(T)$ is embeddable in $R\times S$, where $R$ is a direct product of certain absolutely algebraic fields and $S$ is a finite semisimple ring.
+The next main result in this section shows that if $T$ is a ring that is integral over its center and has only finitely many maximal subrings, then $T/J(T)$ is embeds into a ring of the form $R\times S$, where $R$ is a direct product of certain absolutely algebraic fields and $S$ is a finite semisimple ring.
 Theorem 4.6.
@@
 Then for each maximal ideal $M$ of $T$ either $T/M$ is an absolutely algebraic field with only finitely many maximal subrings or $T/M\cong\mathbb{M}_{n}(F)$ for some finite field $F$ and $n>1$.
-In particular, if $J(T)=0$, then $T$ embeds in
+In particular, if $J(T)=0$, then $T$ embeds into
 $\prod_{i\in I}E_{i}\times\prod_{j\in J}\mathbb{M}_{n_{j}}(F_{j})$
 where each $E_{i}$ is an absolutely algebraic field with only finitely many maximal subrings, $n_{j}>1$ and each $F_{j}$ is a finite field.
-Moreover $J$ is finite, $|I|\leq 2^{\aleph_{0}}$ and if $E_{i}$ is infinite for some $i\in I$, then for each $i^{\prime}\in I\setminus\{i\}$, $E_{i}\ncong E_{i^{\prime}}$.
+Moreover $J$ is finite, $|I|\leq 2^{\aleph_{0}}$ and if $E_{i}$ is infinite for some $i\in I$, then for each $i^{\prime}\in I\setminus\{i\}$, $E_{i}\ncong E_{i^{\prime}}$, as rings.
 Proof.
-Let $M$ be a maximal ideal of $T$, then by assumption $T/M$ has only finitely many maximal subrings and is integral over its center.
+Let $M$ be a maximal ideal of $T$, then by the assumption $T/M$ has only finitely many maximal subrings and is integral over its center.
 Hence by Theorem 3.1, we infer that whenever $T/M$ is a division ring, then $T/M$ is a field with only finitely many maximal subrings and therefore $T/M$ is an absolutely algebraic field, by [10, Corollary 1.5].
-Hence assume that $T/M$ is not a division ring.
+Thus assume that $T/M$ is not a division ring.
 Now by Theorem 3.5, whenever $T/M$ is an infinite ring, then $T/M$ has infinitely many maximal subrings which is impossible.
@@
 Since $T/M$ is not a division ring in this case, we deduce that $n>1$.
-Also note that since $T$ has only finitely many maximal subrings, then $T$ has only finitely many distinct maximal ideals $M$ for them the ring $T/M$ is not a field, by $(5)$ of Proposition 3.10, say $M_{1},\ldots,M_{k}$ (and assume that $T/M_{i}\cong\mathbb{M}_{n_{i}}(F_{i})$, where $F_{i}$ is a finite field and $n_{i}>1$).
-For the next part, by $(8)$ of Proposition 3.10, we have $0=J(T)=\bigcap_{M\in Max(T)}M$.
-Hence we conclude that $0=J(T)=M_{1}\cap\cdots\cap M_{k}\cap\bigcap_{i\in I}N_{i}$, where for each $i\in I$, $T/N_{i}$ is an absolutely algebraic field with only finitely many maximal subrings, by the first part of the theorem.
+Also note that since $T$ has only finitely many maximal subrings, then $T$ has only finitely many distinct maximal ideals $M$ for them the ring $T/M$ is not a field, by Proposition 3.10(5)(5), say $M_{1},\ldots,M_{k}$ (and suppose that $T/M_{i}\cong\mathbb{M}_{n_{i}}(F_{i})$, for some finite field $F_{i}$ and $n_{i}>1$).
+For the next part, by Proposition 3.10(8)(8), we have $0=J(T)=\bigcap_{M\in\operatorname{Max}(T)}M$.
+Consequently by the first part of the proof of the theorem, $0=J(T)=M_{1}\cap\cdots\cap M_{k}\cap\bigcap_{i\in I}N_{i}$, where for each $i\in I$, $E_{i}:=T/N_{i}$ is an absolutely algebraic field with only finitely many maximal subrings.
 Clearly, we may assume that for $i\neq j$ in $I$, $N_{i}\neq N_{j}$.
 Thus $T$ embeds in $T/M_{1}\times\cdots\times T/M_{k}\times\prod_{i\in I}T/N_{i}$.
-Clearly, by Proposition 4.2, we have $|Max(T)|\leq 2^{\aleph_{0}}$ and therefore $|I|\leq 2^{\aleph_{0}}$.
+It follows that from Proposition 4.2, $|\operatorname{Max}(T)|\leq 2^{\aleph_{0}}$ and then $|I|\leq 2^{\aleph_{0}}$.
 Finally we show that if $i\neq j$ in $I$ and $T/N_{i}$ is infinite, then $T/N_{i}\ncong T/N_{j}$.
@@
 ∎
-In the next section we exactly determine when the ring $T=\prod_{i\in I}\mathbb{M}_{n_{i}}(E_{i})$, where each $E_{i}$ is a field and $n_{i}\in\mathbb{N}$, has only finitely many maximal subrings.
-In the following we obtain some results for Artinian/Noetherian rings.
-First we need the following remark.
+In the next section we precisely determine when a ring of the form $T=\prod_{i\in I}\mathbb{M}_{n_{i}}(E_{i})$, where each $E_{i}$ is a field and $n_{i}\in\mathbb{N}$, has only finitely many maximal subrings.
+Below we derive some related results for Artinian/Noetherian rings.
+First we require the following remark.
 Remark 4.7.
-It is not hard to see that if $S$ is a commutative ring with nonzero characteristic, say $n$, $Max(S)=\{M_{1},\ldots,M_{k}\}$ and $J(S)$ is a nil ideal of $S$ (in particular, if $S$ is a commutative Artinian ring with nonzero characteristic $n$), then $S$ is integral over $\mathbb{Z}_{n}$ if and only if each $S/M_{i}$ is an absolutely algebraic field.
+It is not hard to see that if $S$ is a commutative ring with nonzero characteristic, say $n$, $\operatorname{Max}(S)=\{M_{1},\ldots,M_{k}\}$ and $J(S)$ is a nil ideal of $S$ (in particular, if $S$ is a commutative Artinian ring with nonzero characteristic $n$), then $S$ is integral over $\mathbb{Z}_{n}$ if and only if each $S/M_{i}$ is an absolutely algebraic field.
 Theorem 4.8.
@@
 Clearly, $K=C/M\subseteq C(T/Q)$ and $T/Q$ is integral over $K$ and therefore $T/Q$ is integral over its center.
-Thus by $(2)$ of Theorem 4.3, we conclude that $T/Q$ has infinitely many maximal subrings.
+Thus by Theorem 4.3(2)(2), we conclude that $T/Q$ has infinitely many maximal subrings.
 Hence $T$ has infinitely many maximal subrings.
@@
 Since $T/Q$ is integral over $C/P$, by [11, Proposition 1.2], $T/Q$ has a prime ideal $N/Q$ such that $(N/Q)\cap(C/P)=M/P$ and clearly $(T/Q)/(N/Q)$ is integral over $K:=(C/P)/(M/P)\subseteq C((T/Q)/(N/Q))$ (and therefore $(T/Q)/(N/Q)$ is integral over its center).
-Now since $(T/Q)/(N/Q)$ contains an uncountable field $K=(C/P)/(M/P)$, we immediately conclude that $(T/Q)/(N/Q)$ has infinitely many maximal subrings by $(2)$ of Theorem 4.3.
+Now since $(T/Q)/(N/Q)$ contains an uncountable field $K=(C/P)/(M/P)$, we immediately conclude that $(T/Q)/(N/Q)$ has infinitely many maximal subrings by Theorem 4.3(2)(2).
 Thus $T$ has infinitely many maximal subrings too.
@@
 Proof.
-Assume that $T$ has only finitely many maximal subrings, then by $(4)$ of Theorem 4.3, we infer that $T/J(T)$ is a von Neumann regular commutative ring and $T$ is integral over $\mathbb{Z}_{p}$, where $p=Char(K)$ is a prime number.
+Assume that $T$ has only finitely many maximal subrings, then by Theorem 4.3(4)(4), we infer that $T/J(T)$ is a von Neumann regular commutative ring and $T$ is integral over $\mathbb{Z}_{p}$, where $p=\operatorname{Char}(K)$ is a prime number.
 Now note that there exists a natural number $m$ such that $J(T)^{m}=0$, for $T$ is a left Artinian ring.
-Since $T/J(T)$ is an Artinian commutative ring with only finitely many maximal subrings, by $(2)$ of [10, Theorem 2.5], we deduce that $T/J(T)$ is a countable ring.
+Since $T/J(T)$ is an Artinian commutative ring with only finitely many maximal subrings, by [10, Theorem 2.5(2)(2)], we deduce that $T/J(T)$ is a countable ring.
 Hence by Lemma 2.8, $T$ is a countable ring.
@@
 First note that since $T$ is an algebraic $K$-algebra over a field $K$, we infer that $J(T)$ is a nil ideal of $T$, by [21, Theorem 4.20].
-Therefore by [21, Theorem 10.30], $J(T)$ is a nilpotent ideal of $T$, for $T$ is a left Noetherian ring.
-Hence we conclude that $T$ is a left Artinian ring by [21, Theorem 4.15].
+Therefore by [21, Theorem 10.30], $J(T)$ is a nilpotent ideal of $T$, because $T$ is a left Noetherian ring.
+Hence $T$ is a left Artinian ring, by [21, Theorem 4.15].
 Thus by Theorem 4.9, we deduce the assertion.
@@
 Proof.
-$(1)$ Let $J$ be the set of all $i\in I$ such that $T_{i}$ is not a quasi duo ring.
+$(1)$ Let $J$ be the set of all $i\in I$ such that $T_{i}$ is not a quasi-duo ring.
 Now we have two cases.
 First assume that $J$ is infinite.
-For each $j\in J$, let $M_{i}$ be a maximal one-sided ideal of $T_{j}$ which is not an ideal of $T_{j}$, and $N_{j}=\prod_{i\in I}A_{i}$, where $A_{i}=T_{i}$ for $i\neq j$ and $A_{j}=M_{j}$.
-Clearly, each $N_{j}$ is a maximal one-sided ideal of $T$ which is not an ideal of $T$.
-Hence $T$ has infinitely many maximal one-sided ideal which are not ideals of $T$.
-Thus $T$ has infinitely many maximal subrings by $(1)$ of Proposition 3.10.
-Hence assume that $J$ is finite.
-Without lose of the generality we may assume that $J=\emptyset$, i.e., each $T_{i}$ is a quasi duo ring and each $T_{i}$ has only finitely many maximal subrings.
+Without loss of generality we may suppose that for each $j\in J$, the ring $T_{j}$ is not a left quasi-duo ring.
+Consequently, for each $j\in J$, let $M_{i}$ be a maximal left ideal of $T_{j}$ which is not an ideal of $T_{j}$, and $N_{j}=\prod_{i\in I}A_{i}$, where $A_{i}=T_{i}$ for $i\neq j$ and $A_{j}=M_{j}$.
+Clearly, each $N_{j}$ is a maximal left ideal of $T$ which is not an ideal of $T$.
+Hence $T$ has infinitely many maximal left ideal which are not ideals of $T$.
+Thus $T$ has infinitely many maximal subrings by Proposition 3.10(1)(1).
+Now, assume that $J$ is finite.
+Without loss of generality we may suppose that $J=\emptyset$, i.e., each $T_{i}$ is a quasi-duo ring with only finitely many maximal subrings.
 Therefore by Proposition 4.1, for each $i\in I$, the ring $T_{i}/J(T_{i})$ is a commutative ring.
 Take $J:=\prod_{i\in I}J(T_{i})$, then clearly $T/J\cong\prod_{i\in I}T/J(T_{i})$.
-Thus $T/J$ is a product of an infinite family of commutative rings, and therefore by [9, Remark 3.18] (or see [10, Theorem 3.9]), $T/J$ has infinitely many maximal subrings and so does $T$.
+Thus $T/J$ is a product of an infinite family of commutative rings.
+Consequently, by [9, Remark 3.18] (or see [10, Theorem 3.9]), $T/J$ has infinitely many maximal subrings and so does $T$.
 Hence $(1)$ holds.
-For $(2)$, it suffices to show that each $T_{i}$ is integral over its center and use $(1)$.
+$(2)$ It suffices to show that each $T_{i}$ is integral over its center and use $(1)$.
 Clearly, $C_{T}(T)=\prod_{i\in I}C_{T_{i}}(T_{i})$.
@@
 Since $C_{T}(T)=\prod_{i\in I}C_{T_{i}}(T_{i})$, we have $c_{j}=(c_{ji})_{i\in I}$, for $0\leq j\leq n-1$, where $c_{ji}\in C_{T_{i}}(T_{i})$.
-Therefore by calculating the $k$-th component from $x^{n}+c_{n-1}x^{n-1}+\cdots+c_{1}x+c_{0}=0$, we immediately conclude that $t^{n}+c_{n-1k}t^{n-1}+\cdots+c_{1k}t+c_{0k}=0$.
-Hence $t$ is integral over $C_{T_{i}}(T_{i})$ and we are done by $(1)$.
+Now by calculating the $k$-th component of $x^{n}+c_{n-1}x^{n-1}+\cdots+c_{1}x+c_{0}=0$, we obtain that $t^{n}+c_{n-1k}t^{n-1}+\cdots+c_{1k}t+c_{0k}=0$.
+It follows that $t$ is integral over $C_{T_{i}}(T_{i})$ and we are done by $(1)$.
 ∎
 5. Direct product and matrix rings
-In this section we study the number of maximal subrings in matrix rings and (certain) direct products of rings.
-In [10, Theorem 3.9], the authors exactly determined when a direct product of a family of commutative rings has only finitely many maximal subrings.
-In particular, if $T$ is an infinite direct product of a family of commutative rings $\{R_{i}\}_{i\in I}$, then $|RgMax(T)|\geq 2^{|I|}$.
-As we see in the final result of the previous section, if $T$ is a direct product of an infinite family $\{T_{i}\}_{i\in I}$, and either each $T_{i}$ is integral over its center, or $T$ is integral over its center, then $T$ has infinitely many maximal subrings.
-This immediately implies that whenever for each $i$, $T_{i}=\mathbb{M}_{n_{i}}(E_{i})$, where $E_{i}$ is a field and $n_{i}\in\mathbb{N}$, then $T$ has infinitely many maximal subrings (note, for each $i$, $\mathbb{M}_{n_{i}}(E_{i})$ is finite dimensional over its center and therefore is integral over its center).
-In the next proposition we exactly determine when $T$ has only finitely many maximal subrings.
+In this section we study the number of maximal subrings in matrix rings and in certain direct products of rings.
+In [10, Theorem 3.9], the authors precisely determined when a direct product of a family of commutative rings has only finitely many maximal subrings.
+In particular, if $T$ is an infinite direct product of a family of commutative rings $\{R_{i}\}_{i\in I}$, then $|\operatorname{RgMax}(T)|\geq 2^{|I|}$.
+As shown in the final result of the previous section, if $T=\prod_{i\in I}T_{i}$ is a direct product of an infinite family of rings, and either each $T_{i}$ is integral over its center, or $T$ itself is integral over its center, then $T$ has infinitely many maximal subrings.
+This immediately implies that whenever each $T_{i}=\mathbb{M}_{n_{i}}(E_{i})$, with $E_{i}$ is a field and $n_{i}\in\mathbb{N}$, then $T$ has infinitely many maximal subrings (note that each $\mathbb{M}_{n_{i}}(E_{i})$ is finite-dimensional over its center and therefore integral over its center).
+In the next proposition we give precise characterization of when such a product has only finitely many maximal subrings.
 Proposition 5.1.
@@
   3. (3)
-    For each $i,j\in I$, where $n_{i}=n_{j}=1$, and $E_{i}$ is infinite, then $E_{i}\ncong E_{j}$.
+    For each $i\neq j$ in $I$, if $n_{i}=n_{j}=1$, and $E_{i}$ is infinite, then $E_{i}\ncong E_{j}$, as rings.
   4. (4)
     For each $i\in I$ such that $n_{i}=1$, $E_{i}$ has only finitely many maximal subrings.
-Therefore, if $T$ has only finitely many maximal subrings, then $T=E_{1}\times\cdots\times E_{r}\times\mathbb{M}_{n_{1}}(E_{r+1})\times\cdots\mathbb{M}_{n_{s}}(E_{r+s})$, where $n_{i}>1$ and $r,s\geq 0$.
+Consequently, if $T$ has only finitely many maximal subrings, then $T=E_{1}\times\cdots\times E_{r}\times\mathbb{M}_{n_{1}}(E_{r+1})\times\cdots\mathbb{M}_{n_{s}}(E_{r+s})$, where $n_{i}>1$ and $r,s\geq 0$.
 In particular, $T$ is a semisimple ring.
@@
 This immediately implies that each $\mathbb{M}_{n_{i}}(E_{i})$ has finitely many maximal subrings.
-Hence if $n_{i}=1$, $E_{i}$ has only finitely many maximal subrings and if $n_{i}>1$, then by Theorem 3.5, $\mathbb{M}_{n_{i}}(E_{i})$ is finite and therefore $E_{i}$ is finite too.
+Hence if $n_{i}=1$, $E_{i}$ has only finitely many maximal subrings and if $n_{i}>1$, then by Theorem 3.5, $\mathbb{M}_{n_{i}}(E_{i})$ is finite and at once $E_{i}$ is finite too.
 Therefore $(2)$ and $(4)$ hold.
-Although the proof of $(1)$ is obvious by the comment preceding the proposition, but we give another proof too.
+Although the proof of $(1)$ is obvious by the comment preceding the proposition, but we give another proof.
 To see this, let $J:=\{i\in I\ |\ n_{i}=1\}$, $J^{\prime}=I\setminus J$, $A=\prod_{i\in J}E_{i}$ and $B=\prod_{i\in J^{\prime}}\mathbb{M}_{n_{i}}(E_{i})$.
 Then clearly $T=A\times B$.
-Hence by our assumption $A$ and $B$ has only finitely many maximal subrings.
-Thus by [10, Theorem 3.9], we immediately conclude that $J$ is finite.
+Consequently by our assumption $A$ and $B$ has only finitely many maximal subrings.
+Thus by [10, Theorem 3.9], we obtain that $J$ is finite.
 It is clear that by $(5)$ of Proposition 3.10, $J^{\prime}$ is finite and therefore $I$ is finite.
-Finally note that $(3)$ is an immediate by Proposition 3.6.
-Conversely, assume that (1)(4)(1)-(4) hold and let $A$ and $B$ are defined as above.
+Finally note that $(3)$ is an immediate consequence of Proposition 3.6.
+Conversely, suppose that (1)(4)(1)-(4) hold and let $A$ and $B$ are defined as above.
 We prove that $T$ has only finitely many maximal subrings.
-To see this first note that $A$ and $B$ are homomorpically non-isomorphic rings, hence by Lemma 2.9, each maximal subring of $T$ is of the form $A\times B_{1}$, where $B_{1}$ is a maximal subring of $B$, or $A_{1}\times B$, where $A_{1}$ is a maximal subring of $A$.
-This immediately shows that $T$ has only finitely many maximal subrings, for $B$ is finite and by [10, Corollary 3.12 and $(1)$ of Proposition 3.15], $A$ has only finitely many maximal subrings.
+To see this first note that $A$ and $B$ are homomorphically non-isomorphic rings, hence by Lemma 2.9, each maximal subring of $T$ is of the form $A\times B_{1}$, where $B_{1}$ is a maximal subring of $B$, or $A_{1}\times B$, where $A_{1}$ is a maximal subring of $A$.
+This implies that $T$ has only finitely many maximal subrings, for $B$ is finite and by [10, Corollary 3.12 and of Proposition 3.15(1)(1)], AA has only finitely many maximal subrings.
 Thus we are done.
@@
 Assume that $T$ has only finitely many maximal subrings.
-Then similar to the proof of the previous theorem we deduce that $I$ is finite and for each $i\in I$, $\mathbb{M}_{n_{i}}(D_{i})$ has only finitely many maximal subrings.
+Then analogous to the proof of the previous theorem we deduce that $I$ is finite and for each $i\in I$, $\mathbb{M}_{n_{i}}(D_{i})$ has only finitely many maximal subrings.
 Hence by Theorem 3.5, $D_{i}$ is finite and thus $T$ is finite.
@@
   4. (4)
-    for each $i\neq j$ in $I$, whenever $n_{i}=n_{j}=1$ and $D_{i}$ is infinite, then $D_{i}\ncong D_{j}$.
+    For each $i\neq j$ in $I$, if $n_{i}=n_{j}=1$ and $D_{i}$ is infinite, then $D_{i}\ncong D_{j}$, as rings.
 Proof.
@@
 Hence assume that $n_{i}=1$, we claim that $D_{i}$ is a field.
-For otherwise by Proposition 3.1, $D_{i}$ has infinitely many maximal subrings, which is impossible.
+Otherwise by Proposition 3.1, $D_{i}$ has infinitely many maximal subrings, which is impossible.
 Thus for each $i\in I$, we conclude that $D_{i}$ is a field and hence we are done by Proposition 5.1.
 ∎
-In [10, Corollary 3.10], the authors proved that if $R$ is a commutative ring, then $R\times R$ has only finitely many maximal subrings if and only if $R$ is a semilocal ring (i.e., has only finitely many maximal ideals) with only finitely many maximal subrings (in particular, $R$ is integral over $\mathbb{Z}_{n}$ for some $n>1$) and for each maximal ideal $M$ of $R$, the field $R/M$ is finite.
-For non-commutative rings we have the following result.
+In [10, Corollary 3.10], the authors proved that for a commutative ring $R$, the ring $R\times R$ has only finitely many maximal subrings if and only if $R$ is a semilocal ring (i.e., has only finitely many maximal ideals) with only finitely many maximal subrings (in particular, $R$ is integral over $\mathbb{Z}_{n}$ for some $n>1$) and for ever maximal ideal $M$ of $R$, the residue field $R/M$ is finite.
+For noncommutative rings we obtain the following generalization.
 Proposition 5.4.
@@
   1. (1)
-    $Max(R)$, $Max_{l}(R)$ and $Max_{r}(R)$ are finite sets. In particular, $Max(T)$, $Max_{l}(T)$ and $Max_{r}(T)$ are finite sets.
+    $\operatorname{Max}(R)$, $\operatorname{{Max}_{\ell}}(R)$ and $\operatorname{{Max}_{r}}(R)$ are finite sets.
+In particular, $\operatorname{Max}(T)$, $\operatorname{{Max}_{\ell}}(T)$ and $\operatorname{{Max}_{r}}(T)$ are finite sets.
   2. (2)
-    For each $M\in Max(R)$, $R/M$ is finite. In particular, for each maximal ideal $Q$ of $T$, the ring $T/Q$ is finite.
+    For each $M\in\operatorname{Max}(R)$, $R/M$ is finite. In particular, for each maximal ideal $Q$ of $T$, the ring $T/Q$ is finite.
   3. (3)
@@
     For each maximal one-sided ideal $M$ of $R$, $R/M$ is finite (i.e., each simple left/right $R$-module is finite).
-In particular, for each maximal one-sided ideal $N$ of $T$, $T/M$ is finite (i.e., each simple left/right $T$-module is finite).
+In particular, for each maximal one-sided ideal $N$ of $T$, $T/N$ is finite (i.e., each simple left/right $T$-module is finite).
 Proof.
@@
 Hence the first part of $(2)$ holds.
-Also note that for each $M\in Max(R)$, $T$ has a maximal subring $R_{M}$ which contains $M\times M$ (by the comments preceding Theorem 2.7) and clearly whenever $M\neq N\in Max(R)$, then $R_{M}\neq R_{N}$, for $(M\times M)+(N\times N)=T$.
-Hence we conclude that $Max(R)$ is finite.
+Also note that for each $M\in\operatorname{Max}(R)$, $T$ has a maximal subring $R_{M}$ which contains $M\times M$ (by the comments preceding Theorem 2.7) and clearly whenever $M\neq N\in\operatorname{Max}(R)$, then $R_{M}\neq R_{N}$, because $(M\times M)+(N\times N)=T$.
+Hence we conclude that $\operatorname{Max}(R)$ is finite.
 Since $T$ has only finitely many maximal subrings, we infer that $R$ has only finitely many maximal subring.
-This immediately implies that $Max_{l}(R)\setminus Max(R)$ is finite, by $(1)$ of Proposition 3.10.
-Therefore $Max_{l}(R)$ is finite, for $Max(R)$ is finite by the first part of the proof.
-Similarly $Max_{r}(R)$ is finite.
+This implies that $\operatorname{{Max}_{\ell}}(R)\setminus\operatorname{Max}(R)$ is finite, by $(1)$ of Proposition 3.10.
+Consequently $\operatorname{{Max}_{\ell}}(R)$ is finite, because $\operatorname{Max}(R)$ is finite by the first part of the proof.
+Analogously $\operatorname{{Max}_{r}}(R)$ is finite.
 For the second part of $(1)$, note that maximal (left/right) ideals of $T$ are of the forms $M\times R$ or $R\times M$, where $M$ is a maximal (left/right) ideal of $R$.
@@
 For $(3)$, note that since $R$ has only finitely many maximal left ideals, we conclude that $R/J(R)$ is a semisimple ring.
-Hence by $(2)$, we immediately conclude that $R/J(R)$ is finite.
-The second part of $(3)$ is evident, for $J(T)=J(R)\times J(R)$.
-$(4)$ is an immediate consequences of $(3)$, for $R/J(R)$ (resp.
-$T/J(T)$) is a finite ring.
+Hence by $(2)$, we have $R/J(R)$ is finite.
+The second part of $(3)$ is evident, because $J(T)=J(R)\times J(R)$.
+$(4)$ is an immediate consequences of $(3)$, because $R/J(R)$ and $T/J(T)$ are finite rings.
 ∎
-By the previous result we immediately conclude that if $R$ is a ring and $T=R\times R$ has only finitely many maximal subrings, then there exists a finite number of non-isomorphic finite simple left $R$-module $S_{1},\ldots,S_{n}$ such that each simple left $R$-module is isomorphic to one of the $S_{i}$.
+From the previous result we immediately conclude that if $R$ is a ring for which $T=R\times R$ has only finitely many maximal subrings, then there exist finitely many pairwise non-isomorphic finite simple left $R$-modules $S_{1},\ldots,S_{n}$ such that every simple left $R$-module is isomorphic to one of the $S_{i}$.
 Proposition 5.5.
@@
   3. (3)
-    $Max(R)$ and therefore $Max(T)$ are finite. Moreover, $Max_{l}(R)$ and $Max_{r}(R)$ are finite.
+    $\operatorname{Max}(R)$ and therefore $\operatorname{Max}(T)$ are finite. Moreover, $\operatorname{{Max}_{\ell}}(R)$ and $\operatorname{{Max}_{r}}(R)$ are finite.
   4. (4)
-    $R/J(R)$ and therefore $T/J(T)$ are finite semisimple ring. In particular, $Max_{l}(T)$ and $Max_{r}(T)$ are finite.
+    $R/J(R)$ and therefore $T/J(T)$ are finite semisimple rings. In particular, $\operatorname{{Max}_{\ell}}(T)$ and $\operatorname{{Max}_{r}}(T)$ are finite.
   5. (5)
     For each maximal one-sided ideal $M$ of $R$, $R/M$ is finite (i.e., each simple left/right $R$-module is finite).
-In particular, for each maximal one-sided ideal $N$ of $T$, $T/M$ is finite (i.e., each simple left/right $T$-module is finite).
+In particular, for each maximal one-sided ideal $N$ of $T$, $T/N$ is finite (i.e., each simple left/right $T$-module is finite).
 Proof.
-Clearly for each maximal subring $S$ of $R$, $\mathbb{M}_{n}(S)$ is a maximal subring of $T$, thus we deduce that $R$ has only finitely many maximal subrings.
-Hence $(1)$ holds.
-For $(2)$, note that for each maximal ideal $M$ of $R$, $A_{M}=\mathbb{M}_{n}(M)$ is a maximal ideal of $T$ and $T/A_{M}\cong\mathbb{M}_{n}(R/M)$.
+(1)(1)It is obvious that for each maximal subring $S$ of $R$, $\mathbb{M}_{n}(S)$ is a maximal subring of $T$, see [1, Theorem 3.1(2)(2)]. Thus $R$ has only finitely many maximal subrings.
+$(2)$ Note that for each maximal ideal $M$ of $R$, $A_{M}=\mathbb{M}_{n}(M)$ is a maximal ideal of $T$ and $T/A_{M}\cong\mathbb{M}_{n}(R/M)$.
 By the assumption that $T$ has only finitely many maximal subrings, we deduce that $T/A_{M}$ has only finitely many maximal subrings too.
 Hence by Theorem 3.5, we deduce that $R/M$ is finite (note, $n>1$ and therefore $T/A_{M}$ is not a division ring).
-The second part of $(2)$ is obvious by the first part of $(2)$, for maximal ideals of $T$ is of the form $A_{M}$, for some maximal ideal $M$ of $R$.
-For $(3)$, by the notation of $(2)$, since $n>1$, by $(4)$ of Theorem 2.7, we conclude that $\mathbb{M}_{n}(R/M)$ has a maximal subring.
+The second part of $(2)$ is obvious by the first part of $(2)$, because each maximal ideal of $T$ is of the form $A_{M}$, for some maximal ideal $M$ of $R$.
+$(3)$ With the notation of $(2)$, since $n>1$, by Theorem 2.7(4)(4), we conclude that $\mathbb{M}_{n}(R/M)$ has a maximal subring.
 Hence, $T$ has a maximal subring which contains $A_{M}$, say $T_{M}$.
-In particular, if $M\neq N\in Max(R)$, then $T_{M}\neq T_{N}$, for $A_{M}+A_{N}=T$.
-Thus we deduce that $Max(R)$ is finite.
-Therefore $Max(T)$ is finite too.
-For the second part of $(3)$, note that by $(1)$, $R$ has only finitely many maximal subrings, hence by $(1)$ of Proposition 3.10, we conclude that $Max_{l}(R)\setminus Max(R)$ and $Max_{r}(R)\setminus Max(R)$ are finite.
-Therefore $Max_{l}(R)$ and $Max_{r}(R)$ are finite, for $Max(R)$ is finite by the first part of $(3)$.
-For $(4)$, note that since $Max_{l}(R)$ is finite we deduce that $R/J(R)$ is a semisimple ring.
-Since for each maximal ideal $M$ of $R$, $R/M$ is a finite ring, we immediately conclude that $R/J(R)$ is a finite ring.
-Since $J(T)=\mathbb{M}_{n}(J(R))$, we deduce that $T/J(T)$ is finite semisimple too.
-Therefore $Max_{l}(T)$ and $Max_{r}(T)$ are finite.
-$(5)$ is obvious by $(4)$.
-∎
-Now we have the following corollaries.
+In particular, if $M\neq N\in\operatorname{Max}(R)$, then $T_{M}\neq T_{N}$, because $A_{M}+A_{N}=T$.
+Thus $\operatorname{Max}(R)$ is finite, a fortiori $\operatorname{Max}(T)$ is finite.
+For the second part of $(3)$, note that by $(1)$, $R$ has only finitely many maximal subrings, hence Proposition 3.10(1)(1), implies that $\operatorname{{Max}_{\ell}}(R)\setminus\operatorname{Max}(R)$ and $\operatorname{{Max}_{r}}(R)\setminus\operatorname{Max}(R)$ are finite.
+Therefore $\operatorname{{Max}_{\ell}}(R)$ and $\operatorname{{Max}_{r}}(R)$ are finite, because $\operatorname{Max}(R)$ is finite by the first part of $(3)$.
+$(4)$ Since $\operatorname{{Max}_{\ell}}(R)$ is finite, we infer that $R/J(R)$ is a semisimple ring.
+Nothing that for each maximal ideal $M$ of $R$, $R/M$ is a finite ring, consequently $R/J(R)$ is a finite ring.
+Now from the fact that $J(T)=\mathbb{M}_{n}(J(R))$, we deduce that $T/J(T)$ is finite semisimple too.
+Therefore $\operatorname{{Max}_{\ell}}(T)$ and $\operatorname{{Max}_{r}}(T)$ are finite.
+$(5)$ It is obvious by $(4)$. ∎
+We now state several corollaries.
 Corollary 5.6.
@@
 Corollary 5.7.
-Let $R$ be a zero-dimensional ring with $Char(R)=0$. Then $\mathbb{M}_{n}(R)$, where $n>1$, and $R\times R$ has infinitely many maximal subrings.
+Let $R$ be a zero-dimensional ring with $\operatorname{Char}(R)=0$. Then $\mathbb{M}_{n}(R)$, where $n>1$, and $R\times R$ have infinitely many maximal subrings.
 Proof.
 Assume that $\mathbb{M}_{n}(R)$ or $R\times R$ has only finitely many maximal subrings.
-Then by Propositions 5.4 or 5.5, we deduce that for each maximal ideal $M$ of $R$, $R/M$ is a finite ring and therefore $Char(R/M)\neq 0$.
+Then by Propositions 5.4 or 5.5, we deduce that for each maximal ideal $M$ of $R$, $R/M$ is a finite ring and therefore $\operatorname{Char}(R/M)\neq 0$.
 Now note that there exists a prime ideal $P$ of $R$ such that $P\cap\mathbb{Z}=0$.
@@
 Then $T/J(T)\cong D_{1}\times\cdots\times D_{n}\times\mathbb{M}_{n_{1}}(E_{1})\times\cdots\times\mathbb{M}_{n_{k}}(E_{k})$, where each $D_{i}$ is a division ring, $1\leq i\leq n$, $E_{j}$ is a finite field, $1\leq j\leq k$ and $n_{j}>1$, for some $n,k\geq 0$.
-Moreover, whenever $D_{i}$ is infinite, then $D_{i}\ncong D_{j}$ for $j\neq i$.
+Moreover, whenever $D_{i}$ is infinite, then $D_{i}\ncong D_{j}$ as rings, for each $j\neq i$.
 Proof.
-Since $T/J(T)$ is a semisimple ring, we deduce that $T/J(T)\cong\prod_{i\in I}\mathbb{M}_{n_{i}}(D_{i})$, where $I$ is finite, $D_{i}$ are division rings and $n_{i}\in\mathbb{N}$.
+Since $T/J(T)$ is a semisimple ring, we deduce that $T/J(T)\cong\prod_{i\in I}\mathbb{M}_{n_{i}}(D_{i})$, where $I$ is finite, $D_{i}$ is a division ring and $n_{i}\in\mathbb{N}$.
 Clearly $T$ has only finitely many maximal ideals.
-Also note that since $T$ has only finitely many maximal subrings, then by $(1)$ of Proposition 3.10, we conclude that $Max_{l}(T)\setminus Max(T)$ is finite.
-Therefore $Max_{l}(T)$ is finite.
-Similarly $Max_{r}(R)$ is finite.
-Hence by Corollary 2.4, we deduce that whenever $i\in I$ and $n_{i}>1$, then $D_{i}$ is finite, i.e., $D_{i}$ is a finite field.
-Hence we infer that $T/J(T)\cong D_{1}\times\cdots\times D_{n}\times\mathbb{M}_{n_{1}}(E_{1})\times\cdots\times\mathbb{M}_{n_{k}}(E_{k})$, where each $D_{i}$ is a division ring, $1\leq i\leq n$ and $E_{j}$ is a finite field, $1\leq j\leq k$ and $n_{j}>1$ for some $n,k\geq 0$.
-Finally if for some $1\leq i\neq j\leq n$, $D_{i}\cong D_{j}$, then $T$ has an ideal $I$ such that $T/I\cong D_{i}\times D_{j}\cong D_{i}\times D_{i}$.
+Also note that since $T$ has only finitely many maximal subrings, then by Proposition 3.10(1)(1), we conclude that $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$ is finite.
+Therefore $\operatorname{{Max}_{\ell}}(T)$ is finite.
+Similarly $\operatorname{{Max}_{r}}(R)$ is finite.
+Obviously for each $i\in I$, the ring $\mathbb{M}_{n_{i}}(D_{i})$ has only finitely many maximal subrings, thus by Corollary 2.4, we deduce that whenever $i\in I$ and $n_{i}>1$, then $D_{i}$ is finite, i.e., $D_{i}$ is a finite field.
+Consequently $T/J(T)\cong D_{1}\times\cdots\times D_{n}\times\mathbb{M}_{n_{1}}(E_{1})\times\cdots\times\mathbb{M}_{n_{k}}(E_{k})$, where each $D_{i}$ is a division ring, $1\leq i\leq n$ and $E_{j}$ is a finite field, $1\leq j\leq k$ and $n_{j}>1$, for some $n,k\geq 0$.
+Finally if for some $1\leq i\neq j\leq n$, $D_{i}\cong D_{j}$ as rings, then $T$ has an ideal $I$ such that $T/I\cong D_{i}\times D_{j}\cong D_{i}\times D_{i}$.
 Since $T$ has only finitely many maximal subrings, we deduce that $T/I$ has only finitely many maximal subrings.
-Hence by Proposition 3.6, $D_{i}$ is finite.
+Thus by Proposition 3.6, $D_{i}$ is finite, which is impossible.
 Hence we are done.
 ∎
-We conclude this paper by the following result.
+We conclude this paper by the following immediate corollary.
 Corollary 5.10.
@@
 Acknowledgement
-The first author is grateful to the Research Council of Shahid Chamran University of Ahvaz (Ahvaz-Iran) for financial support (Grant Number: SCU.MM1404.721)
+The author is grateful to the Research Council of Shahid Chamran University of Ahvaz (Ahvaz-Iran) for financial support (Grant Number: SCU.MM1404.721)